### v1

# PARCO: Learning Parallel Autoregressive Policies for Efficient Multi-Agent Combinatorial Optimization

Abstract
Multi-agent combinatorial optimization problems such as routing and scheduling have great practical relevance but present challenges due to their NP-hard combinatorial nature, hard constraints on the number of possible agents, and hard-to-optimize objective functions. This paper introduces PARCO (Parallel AutoRegressive Combinatorial Optimization), a novel approach that learns fast surrogate solvers for multi-agent combinatorial problems with reinforcement learning by employing parallel autoregressive decoding. We propose a model with a Multiple Pointer Mechanism to efficiently decode multiple decisions simultaneously by different agents, enhanced by a Priority-based Conflict Handling scheme. Moreover, we design specialized Communication Layers that enable effective agent collaboration, thus enriching decision-making. We evaluate PARCO in representative multi-agent combinatorial problems in routing and scheduling and demonstrate that our learned solvers offer competitive results against both classical and neural baselines in terms of both solution quality and speed. We make our code openly available at https://github.com/ai4co/parco.
[…]
1 Introduction
Combinatorial optimization (CO) problems, such as routing and scheduling problems, involve determining an optimal sequence of actions in a combinatorial space and have applications ranging from warehouse operatons (Xie et al., 2023) to network design (Chabarek et al., 2008) and safety-critical systems (Girardey et al., 2010). CO problems are notoriously hard to solve and cannot generally be solved optimally in polynomial time, i.e., they are NP-hard (Jünger et al., 1995). Multi-agent settings gained significant interest due to their applicability in realistic scenarios, such as path finding (Reijnen et al., 2020), drone routing (Ann et al., 2015), disaster management (Bektas, 2006; Cheikhrouhou and Khoufi, 2021) and order delivery (Yakıcı and Karasakal, 2013; Archetti and Bertazzi, 2021), but present even more challenges due to additional constraints and different optimization objectives, including minimizing a global lateness objective or the makespan (Mahmoudinazlou and Kwon, 2024).
While traditional algorithmic methods have significantly contributed to solving a range of problems (Laporte and Osman, 1995; Hejazi and Saghafian, 2005), these approaches often concentrate on single-agent scenarios. With the advent of modern computational techniques, neural network-based approaches have started to yield promising results for complex CO problems in a field known as Neural Combinatorial Optimization (NCO). In particular, Reinforcement Learning (RL) has shown promise due to its ability to learn directly from interactions with environments instead of relying on (costly) labeled datasets in the shape of optimal solutions (Bello et al., 2016; Kwon et al., 2020; Kim et al., 2023a).
Among NCO methods, Autoregressive (AR) sequence generation has gained attention for its ability to manage hard constraints (Kool et al., 2018; Kwon et al., 2021). In the context of NCO, this capability is crucial for addressing complex problems with multiple construction constraints, such as the precedence constraints prevalent in scheduling (Zhang et al., 2020a) as well as pickup and delivery problems (Savelsbergh and Sol, 1995; Parragh et al., 2008). However, one issue of autoregressive sequence generation is the high latency associated with it, especially when considering large problem instances. The problem of high inference latency of AR methods is especially prevalent in the domain of large language models, as multiple stacks of transformer layers have to be computed sequentially for each individual token (Bae et al., 2023).
Figure 1: Solution construction with Autoregressive (top) and Parallel Autoregressive policies (bottom).
Motivated by recent studies on LLMs demonstrating parallel decoding can not only tackle the high generation latency problem of next-token predictors but also improve their answer quality (Qi et al., 2020; Ning et al., 2023; Gloeckle et al., 2024) this paper introduces PARCO (Parallel AutoRegressive Combinatorial Optimization), a novel approach designed to address multi-agent combinatorial problems efficiently. We propose to apply parallel decoding in the NCO domain to increase solution construction efficiency and effectiveness. In contrast to most previous NCO methods, PARCO constructs solutions in parallel for multiple agents, made feasible by a priority-based conflict handler to manage conflicting decisions of different agents. Importantly, we enable the collaborative behavior of agents through specialized inter-agent Communication Layers, essential for finding high-quality solutions to multi-agent combinatorial problems.
Contributions.
We summarize our contributions as follows:
  • We introduce PARCO, a novel approach for tackling multi-agent CO problems by effectively constructing solutions in parallel via a Multiple Pointer Mechanism.
  • We introduce an efficient Priority-based Conflict-Handling scheme to resolve conflict between agents and enhance decision-making.
  • We design a special Communication Layer to help agents communicate and collaborate to obtain a better state-space representation of partial solutions.
  • We show competitive performance against classical solvers and neural baselines in routing and scheduling problems, both in terms of solution quality and computational efficiency.
2 Related Work
Neural Combinatorial Optimization
Recent advancements in NCO have shown promising end-to-end solutions for combinatorial optimization problems, as highlighted by Bengio et al. (2021) and Yang and Whinston (2023). NCO has led to the development of aligned neural architectures (Bi et al., 2022; Jin et al., 2023; Luo et al., 2023; Kim et al., 2023c), hybrid methods with OR solvers (Li et al., 2021b; Kim et al., 2024a; Yan and Wu, 2024; Ye et al., 2024a; Kim et al., 2024b), multi-level solution pipelines (Ma et al., 2023a; Li et al., 2023; Xiao et al., 2023; Ye et al., 2023), alongside improved training algorithms (Kim et al., 2023c; Jiang et al., 2023; Drakulic et al., 2023; Sun and Yang, 2023; Gao et al., 2023; Xiao et al., 2023; Li et al., 2024b; Wang et al., 2024) to enhance the heuristic search. These innovations have expanded NCO’s application across a wide array of problems (Chen et al., 2023; Kim et al., 2023b; Zhou et al., 2023; Ma et al., 2023b; Luttmann and Xie, 2024). However, integrating appropriate inductive biases requires manual tuning of model architectures and training algorithms, presenting challenges such as computationally intensive training, the need for specialized hardware, and issues with interpretability and generalizability (Liu et al., 2023).
Multi-agent Constructive Methods for NCO
While several seminal works as Vinyals et al. (2015); Kool et al. (2018); Kwon et al. (2020) propose models that can be used in loose multi-agent settings, such methods cannot be employed directly to model heterogeneous agents with different attributes, such as different capacities or starting locations in routing problems. Zhang et al. (2020b); Falkner and Schmidt-Thieme (2020); Li et al. (2022); Liu et al. (2024c) introduce models to construct solutions of different agents simultaneously. However, the above parallel construction is, in fact, sequential in terms of the model and environment stepping itself, thus not benefitting from true model-side parallel decoding and often lacking in agent collaboration. On the other hand, Son et al. (2024); Zheng et al. (2024) propose to model multi-agent min-max routing as sequential decision-making, allowing for switching the agents after a single-agent solution is complete. Most related to PARCO is Zong et al. (2022), who propose a multi-agent model for the min-sum multi-agent PDP with different decoders. However, this approach is not only tailored to a specific CO problem but also a specific number of agents due to both different decoders for each agent and a non-flexible context embedding, which is dependent on the agent number, while PARCO is more flexible and has a more general communication representation. Moreover, while Zong et al. (2022) manages conflicts by randomly giving precedence to one agent, PARCO learns how to prioritize agents with higher probabilities of selecting a specific node.
3 Background
3.1 Markov Decision Processes
CO problems can be framed as Markov Decision Processes (MDPs). In this formulation, the problem is defined by a set of states $\mathcal{S}$, where each state $s_{t}\in\mathcal{S}$ represents the configuration of the problem at time step $t$. At each step, an agent selects an action $a_{t}$ from the action space $\mathcal{A}$ according to a policy $\pi:\mathcal{S}\rightarrow\mathcal{A}$, which maps states to actions. The system then transitions from state $s_{t}$ to state $s_{t+1}$ according to a transition function $T:\mathcal{S}\times\mathcal{A}\rightarrow\mathcal{S}$. The agent reaches the terminal state once it has generated a viable solution for the problem instance $\bm{x}$. Typically, a reward $R$ is only obtained in the terminal state and takes the form of the objective function of the respective CO problem.
3.2 Autoregressive Methods for NCO
Given the sequential nature of MDPs, autoregressive (AR) methods pose a natural choice for the policy $\pi$. AR methods construct a viable solution by sequentially generating actions based on the current state and previously selected actions. Formally, the process can be represented as:
$p_{\theta}(\bm{a}|\bm{x})$ $\triangleq\prod_{t=1}^{T}g_{\theta}(a_{t}|a_{t-1},...,a_{0},\bm{h})$ (1)
where $\bm{h}=f_{\theta}(\bm{x})$ is the encoding of problem instance $\bm{x}$ via encoder network $f$, which is used to decode actions autoregressively via decoder $g_{\theta}$. $p_{\theta}$ is a solver that maps $\bm{x}$ to a solution $\bm{a}$. $\bm{a}=(a_{1},...,a_{T})$ represents the (optimal) actions executed in $T$ construction steps, resulting in a feasible solution to CO problems.
Notably, the same AR scheme has been applied to solve multi-agent CO problems, such as vehicle routing problems (Kool et al., 2018; Kwon et al., 2020), in which a single agent tasked with visiting all nodes under specific route length constraints, distinguishing it from a real multi-agent scenario by focusing on optimizing a solitary agent’s path within set limits. To tackle such problems, AR methods sequentially construct multiple routes as if handling a single agent at a time. However, inducing collaborative behavior among agents using the AR method poses a challenge, as it inherently focuses on one agent at a time. This limitation restricts the method’s ability to directly foster interactions among the agents, which is crucial for optimizing collective outcomes in multi-agent scenarios.
3.3 Training AR Models via Reinforcement Learning
In this work, we focus on Reinforcement Learning (RL) solvers for training $p_{\theta}$ as they can be trained without relying on (often hard-to-obtain) labeled solutions.
We can thus cast the optimization problem as follows:
[…]
where $P(\bm{x})$ is problem distribution, $R(\bm{a},\bm{x})$ is reward of $\bm{a}$ given $\bm{x}$. We can solve Eq. 2 by employing various RL methods such as variations of policy gradients (REINFORCE) (Kwon et al., 2020; Kim et al., 2022).
4 Methodology
4.1 Parallel Multi-Agent Environments
We propose to step multiple actions over the environment simultaneously to enhance efficiency in practice, which can reduce the number of steps required compared to single-agent stepping. In the general multi-agent CO setting, agents $k=1,...,m$ are selecting actions $\bm{a}_{t}=(a_{t}^{1},...,a_{t}^{m})$ from a shared action space $\mathcal{A}$ in parallel at a given decoding step $t$. Given the agent actions $\bm{a}_{t}$, the state of the problem instance transitions from $s_{t}$ to $s_{t+1}$ via some transition function $T:\mathcal{S}\times\mathcal{A}_{1}\times...\times\mathcal{A}_{m}\rightarrow\mathcal{S}$, which usually follows deterministic rules in case of non-stochastic CO problems. After reaching the terminal state, the agents receive a shared reward $R(A,\bm{x})$, with $\bm{x}$ the problem instance and $A=(\bm{a}_{1},...,\bm{a}_{T})$ the sequence of agents actions, which is the objective of the respective CO problem. Fig. 1 illustrates an example of AR and Parallel AR solution construction.
In this work, we propose a model architecture, PARCO, to solve such cooperative multi-agent CO problems efficiently. Drawing on studies on multi-agent RL (MARL) which promotes parameter sharing in cooperative MARL settings with a shared action space (Yu et al., 2022), PARCO establishes a central auto-regressive policy $\pi_{\theta}:\mathcal{S}\rightarrow P(\mathcal{A}_{1}\times...\times\mathcal{A}_{m})$ to decode the next action for the $m$ agents in parallel. This not only makes PARCO agnostic to the number of agents prevalent in the respective CO problem, but also enhances both solution quality and construction speed.
4.2 Parallel Autoregressive Model
In this section, we present our general architecture to solve CO problems efficiently cast as multi-agent sequential decision-making problems. PARCO follows the general encoder-decoder architecture prevalent in autoregressive NCO (Kool et al., 2018; Bello et al., 2016; Vinyals et al., 2015).
Figure 2: Overview of the PARCO model.
Encoder
The job of the encoder in autoregressive NCO approaches is to learn a mapping from the graph representing the CO problem to a higher dimensional embedding space. This typically involves a problem depending on the initial embedding layer, projecting the graph nodes using their feature representations. This is also the case for PARCO, whose encoder first maps features of the graph instance to an embedding space to learn deep relationships between the nodes through a stack of transformer blocks, similar to Kool et al. (2018). The core of each transformer block is a multi-head attention (MHA) layer, which enables message passing between the nodes in the graph and can be defined as follows (Vaswani et al., 2017):
$\text{MHA}(Q,K,V)$ $=\left(\mathop{\big{\|}}_{i=1}^{h}\text{Attn}(QW^{Q}_{i},KW^{K}_{i}VW^{V}_{i})\right)W^{O}$
where
$\text{Attn}(Q,K,V)$ $=\text{softmax}\left(\frac{QK^{\top}}{\sqrt{d_{k}}}\right)V\,.$
[…]
PARCO additionally defines an agent projection layer that maps agents into the same embedding space as the problem nodes, whose exact definition heavily depends on the problem. In routing problems, for example, agent embeddings can be represented by the depot location they are starting from and their capacities, while an agent in scheduling problems can be represented by the machine whose schedule it is constructing. Lastly, encoded actions and agents can further be enhanced by performing cross-attention between them, depending on the structure of the problem (Kwon et al., 2021). We describe the specific encoder details for the different problems covered in this work in the Appendix.
Having a separate agent encoder contrasts other multi-agent (NCO) approaches like Zhang et al. (2020b), which use separate policies $\pi^{k}_{\theta_{m}}$ for each agent. This has a substantial drawback, as a model trained for a given number of agents cannot be used to solve problems with a different number of agents. The PARCO encoder, on the other hand, can generate representations for arbitrary agents, making it agnostic to the problem size.
Decoder
The decoder is the core of the PARCO architecture. We adopt the pointer mechanism from Kool et al. (2018), which yields unnormalized log-probabilities $\bm{z}\in\mathbb{R}^{N}$ for each action/node:
$\bm{z}=C\cdot tanh\left(\frac{\bm{u}^{\top}L}{\sqrt{d_{h}}}\right)$ (3)
where $C$ is a scale parameter, $L$ is a projection of the node embeddings $\bm{h}$ and $\bm{u}$ is the output of an MHA layer, in which the query is defined by a single context embedding $\bm{q}^{c}_{t}\in\mathbb{R}^{d_{h}}$, specifying the current state of the problem at decoding step $t$ and keys and values are the node embeddings from the encoder. Each attention head of the MHA layer can further fuse the projection of the encoder embeddings with dynamic embeddings representing current information about the particular node, i.e.
$\text{Attn}(\bm{q}^{c}_{t}W^{Q}_{i},\,\bm{h}W^{K}_{i}+W^{K}_{\delta}\delta_{t},\,\bm{h}W^{V}_{i}+W^{V}_{\delta}\delta_{t})$ (4)
with $\delta_{t}$ encapsulating the dynamic elements of the problem nodes and $W^{K}_{\delta},W^{V}_{\delta}\in\mathbb{R}^{|\delta|\times d_{m}}$ projecting them into embedding space.
For PARCO we introduce a novel Multiple Pointer Mechanism based on Eq. 3 of the following form:
$Z=C\cdot tanh\left(\frac{UL^{\top}}{\sqrt{d_{h}}}\right)$ (5)
Here, $U\in\mathbb{R}^{m\times d_{h}}$ is the output of an equivalent MHA layer than the one used in Kool et al. (2018). However, instead of a single context embedding, PARCO uses the $m$ agent embeddings of the encoder, enriched by the current state’s context and individual agent’s state embeddings, as queries $Q^{c}_{t}\in\mathbb{R}^{m\times d_{h}}$. The resulting $Z\in\mathbb{R}^{m\times N}$ log-probabilities are used to sample $m$ actions - one for each agent - in parallel.
Communication Layers
To enable effective agent coordination, we introduce a communication layer before the Multiple Pointer Mechanism, which updates the agent embeddings through messages exchanged with other agents. This component is vital for parallel decoding of multiple actions, as the quality of an action chosen by one agent is significantly influenced by how it affects the actions of other agents.
Given the $m$ agent queries at decoding step $t$, the communication layer applies a transformer block (Vaswani et al., 2017) to capture intra-agent relationships and spot potential conflicts between agents:
$H^{\prime}$ $=\text{Norm}(\text{MHA}(Q^{c}_{t},Q^{c}_{t},Q^{c}_{t})+Q^{c}_{t})$ (6)
$Q^{\prime}_{t}$ $=\text{Norm}(\text{MLP}(H^{\prime})+H^{\prime})$ (7)
where Norm denotes a normalization layer (Ioffe and Szegedy, 2015; Zhang and Sennrich, 2019) and MLP represents the multi-layer perceptron. Self-attention in the MHA layer enables message passing between agents based on their embeddings. By applying the communication layer after fusing the agent embeddings with the dynamic context embeddings, it can capture dynamic relationships that could not be resolved during encoding, thereby enabling the decoder to make more informed decisions. The updated query $Q^{\prime}_{s}$ is finally passed into the Multiple Pointer Mechanism.
4.2.1 Conflicts Handlers
When sampling from the probability distribution generated by the Multiple Pointer Mechanism, it is possible for multiple agents to select the same node simultaneously, resulting in a conflict. These conflicts must be resolved to ensure the generation of a feasible solution. To address this, we introduce a Priority-based Conflict Handler $\mathcal{H}$, which resolves such conflicts by leveraging a predefined priority scheme. The conflict handler can be defined as a function $\mathcal{H}:\mathbb{N}^{m}\times\mathbb{R}^{m}\rightarrow\mathbb{N}^{m}$ with number of agents $m$. The priorities are determined based on the current actions and states. Given an input vector of actions $\bm{a}=(a_{1},a_{2},\dots,a_{m})$ where $a_{i}\in\mathbb{N}$, the corresponding priority vector $\bm{p}=(p_{1},p_{2},\dots,p_{m})$ where $p_{i}\in\mathbb{R}$ and the fallback actions (i.e., the previous node) $\bm{r}=(r_{1},r_{2},\dots,r_{m})$ where $r_{i}\in\mathbb{N}$, the output is another vector of actions $\bm{a}^{\prime}=(a^{\prime}_{1},a^{\prime}_{2},\dots,a^{\prime}_{m})$ such that any conflicts are resolved based on the priority order as illustrated in Algorithm 1.
Algorithm 1 Priority-based Conflict Handler
0:  Actions $\bm{a}\in\mathbb{N}^{m}$, Priorities $\bm{p}\in\mathbb{R}^{m}$, Fallback actions $\bm{r}\in\mathbb{N}^{m}$
0:  Resolved Actions $\bm{a^{\prime}}\in\mathbb{N}^{m}$
1:  $\sigma\leftarrow\text{argsort}(\bm{p},\text{descending}=\text{True})$Sort indices by priority}
2:  $\hat{\bm{a}}\leftarrow\bm{a}[\sigma]$Reorder actions according to priority}
3:  Initialize conflict mask $M\leftarrow\bm{0}^{m}$
4:  for $i=2$ to $m$ do
5:     if $\hat{\bm{a}}_{i}\in\{\hat{\bm{a}}_{1},\ldots,\hat{\bm{a}}_{i-1}\}$ then
6:        $M_{i}\leftarrow 1$Identify conflicts}
7:     end if
8:  end for
9:  $\hat{\bm{a}}\leftarrow(1-M)\odot\hat{\bm{a}}+M\cdot r$Resolve conflicts by assigning fallback action $r$}
10:  $\bm{a}^{\prime}\leftarrow\hat{\bm{a}}[\sigma^{-1}]$Reorder actions back to original sequence}
11:  return  $\bm{a^{\prime}}$
4.2.2 Step Definition
We define a single parallel step of the model similarly to Eq. 1 as follows:
$p_{\theta}(\bm{a}|\bm{x})$ $=\prod_{i=1}^{m}g_{\theta}(a_{t,i}^{\prime}|a_{t-1,i}^{\prime},\dots,a_{0,i}^{\prime},\bm{h})$ (8)
Here, $a_{t,i}^{\prime}$ denotes the action executed by agent $i$ at time $t$, after passing through the conflict handler $\mathcal{H}$:
$a_{t,i}^{\prime}=\mathcal{H}(a_{t,1},a_{t,2},\dots,a_{t,m};\bm{p}_{t},\bm{r}_{t}),$
where $\bm{p}_{t}$ and $\bm{r}_{t}$ are priorities and fallback actions, respectively, at the current decoding step $t$. The conflict handler $\mathcal{H}$ ensures that multiple agents do not select the same node, thus maintaining the integrity and feasibility of the solution.
4.3 Training
PARCO is a centralized multi-agent sequential decision-making model, with a shared policy $\pi_{\theta}$ for all agents and a global reward $R$. Thus, PARCO can be trained using any of the training algorithms adopted in the single-agent NCO literature. We train PARCO via the REINFORCE gradient estimator (Williams, 1992) with a shared baseline as outlined by Kwon et al. (2020) and Kim et al. (2022):
$\nabla_{\theta}\mathcal{L}\approx\frac{1}{B\cdot L}\sum_{i=1}^{B}\sum_{j=1}^{L}G_{ij}\nabla_{\theta}\,\text{log}\,p_{\theta}(A_{ij}|\bm{x}_{i})$ (9)
where $B$ is the size of the mini-batch and $G_{ij}$ is the advantage $R(A_{ij},\bm{x}_{i})-b^{\text{shared}}(\bm{x}_{i})$ of a solution $A_{ij}$ compared to the shared baseline $b^{\text{shared}}_{i}$ of problem instance $\bm{x}_{i}$.
5 Experiments
In this section, we present the experimental results of PARCO in two routing problems, the min-max heterogenous capacitated vehicle routing problem (HCVRP) and the open multi-depot capacitated pickup and delivery problem (OMDCPDP), and a scheduling problem, namely the flexible flow shop problem (FFSP). We provide more details about the problem and experimental setups in Appendix A and Appendix B respectively.
5.1 Problem Descriptions
HCVRP
The min-max HCVRP consists of $m$ agents sequentially visiting customers to satisfy their demands, with constraints including each customer can be visited exactly once and the amount of demand satisfied by a single vehicle in a trip cannot exceed its capacity, which can be reloaded by going back to the depot. The goal is to minimize the makespan, i.e., the worst route. Baselines include SISR (Christiaens and Vanden Berghe, 2020), Genetic Algorithm (GA) (Karakatič and Podgorelec, 2015), Simulated Annealing (SA) (İlhan, 2021), the Attention Model (AM) (Kool et al., 2018), Equity Transformer (ET) (Son et al., 2024), the model from Li et al. (2022) (DRLLi), and the state-of-the-art neural baseline 2D-Ptr (Liu et al., 2024c).
OMDCPDP
The OMDCPDP problem is a practical variant of the pickup and delivery problem in which agents have a stacking limit of orders that can be carried at any given time. Pickup and delivery locations are paired, and pickups must be visited before deliveries. Multiple agents start from different depots with no need to go back (open). The goal is to minimize the sum of arrival times to delivery locations, i.e. minimizing the cumulative lateness. We include ORTools (Furnon and Perron, 2024) as a classical baseline, the Heterogeneous Attention Model (HAM) (Li et al., 2021a) for sequential decision-making and MAPDP (Zong et al., 2022) for parallel decision-making.
FFSP
In the flexible flow shop problem (FFSP), $N$ jobs must be processed across $S$ stages, each with multiple machines ($m>1$). Jobs follow a specified sequence through these stages, but within each stage, any available machine can process the job, with the key constraint that no machine can handle more than one job simultaneously. The goal is to schedule the jobs so that all jobs are finished in the shortest time possible. Notable benchmarks include the MatNet model (Kwon et al., 2021), the Random and Shortest Job First (SJF) dispatching rules, as well as the evolutionary algorithms Particle Swarm Optimization (PSO), and Genetic Algorithm (GA) (Hejazi and Saghafian, 2005).
Table 1: Benchmarks and results of our model for min-max HCVRP problems of varying sizes and agent numbers. Highlighting cost ($\downarrow$) and average gaps ($\downarrow$) from the best-known solutions of classical heuristic solvers. Inference times are shown in seconds in parentheses ()($\cdot$)( ⋅ ). (g.) refers to greedy performance while (s.) refers to sampling 1280 solutions.
[…]
$m$ 3 5 7 3 5 7 3 5 7 avg.
SISR 6.57 (271) 4.00 (274) 2.91 (276) 8.52 (425) 5.10(430) 3.69 (434) 10.29 (615) 6.17 (623) 4.45 (625) 0.00
GA 9.21 (233) 6.89 (320) 5.98 (405) 12.32 (347) 8.95 (465) 7.58 (578) 15.33 (479) 10.93 (623) 9.10 (772) 74.90
SA 7.04 (130) 4.39 (289) 3.30 (362) 9.17 (318) 5.61 (417) 4.17 (515) 11.13 (434) 6.80 (557) 5.01 (678) 10.21
AM (g.) 8.49 (0.08) 5.51 (0.08) 4.15 (0.09) 10.81 (0.10) 6.87 (0.11) 5.18 (0.10) 12.68 (0.14) 8.10 (0.13) 6.13 (0.13) 33.80
ET (g.) 7.58 (0.15) 4.76 (0.17) 3.58 (0.16) 9.76 (0.21) 6.01 (0.20) 4.43 (0.23) 11.74 (0.25) 7.25 (0.25) 5.23 (0.26) 17.67
DRLLi (g.) 7.43 (0.19) 4.71 (0.22) 3.60 (0.25) 9.64 (0.25) 5.97 (0.30) 4.52 (0.33) 11.44 (0.32) 7.06 (0.37) 5.38 (0.43) 17.08
2D-Ptr (g.) 7.20 (0.11) 4.48 (0.11) 3.31 (0.11) 9.24 (0.15) 5.65 (0.15) 4.14 (0.14) 11.12 (0.18) 6.75 (0.18) 4.92 (0.17) 10.54
PARCO (g.) 7.12 (0.12) 4.40 (0.11) 3.25 (0.11) 9.14 (0.15) 5.53 (0.15) 4.04 (0.15) 10.98 (0.19) 6.61 (0.18) 4.79 (0.18) 8.56
AM (s.) 7.62 (0.14) 4.82 (0.13) 3.63 (0.14) 9.92 (0.20) 6.19 (0.21) 4.64 (0.22) 11.82 (0.29) 7.45 (0.28) 5.58 (0.28) 20.69
ET (s.) 7.14 (0.21) 4.46 (0.22) 3.33 (0.22) 9.19 (0.30) 5.64 (0.30) 4.17 (0.31) 11.20 (0.41) 6.85 (0.38) 4.98 (0.40) 10.89
DRLLi (s.) 6.97 (0.30) 4.34 (0.36) 3.25 (0.43) 9.10 (0.45) 5.54 (0.55) 4.13 (0.65) 10.90 (0.60) 6.65 (0.76) 4.98 (0.92) 8.78
2D-Ptr (s.) 6.82 (0.13) 4.20 (0.13) 3.09 (0.14) 8.85 (0.17) 5.36 (0.18) 3.90 (0.19) 10.71 (0.22) 6.46 (0.23) 4.68 (0.24) 4.84
PARCO (s.) 6.82 (0.31) 4.17 (0.31) 3.06 (0.31) 8.79 (0.36) 5.28 (0.36) 3.83 (0.35) 10.61 (0.40) 6.36 (0.40) 4.58 (0.40) 3.65
5.2 Experimental Setup
We perform all experiments on a machine equipped with two Intel(R) Xeon(R) Gold 6338 CPU @ 2.00GHz CPUs with a total $128$ threads and $8$ NVIDIA RTX 4090 graphic cards with $24$ GB of VRAM. Training runs of PARCO take less than 24 hours each. During inference, we employ only one CPU and a single GPU. We report key metrics such as solution cost, inference times, and gaps in best-known solutions. We keep most settings of PARCO consistent across experiments, i.e., we use a Communication Layer and Priority-based Conflict Handling based on the highest probability action from the model output. We provide additional details regarding training and testing setups in the Appendix.
(a) Effect of Communication Layers.
(b) Effect of Conflict Handlers.
(c) PARCO vs AR decoding speed.
Figure 3: Ablation studies on PARCO components. PARCO scales better than AR models as ET.
Table 2: Benchmarks and results of our model for OMDCPDP problems of varying sizes and agent numbers.
[…]
$N$ 50 100 200 1000 avg.
$m$ 10 18 25 20 35 50 40 70 100 100 250 500
OR-Tools 23.73 18.64 16.92 43.34 34.80 31.89 78.35 64.93 60.16 562.11 333.01 310.55 0.00%
HAM (g.) 34.85 24.79 18.52 65.95 51.56 35.41 124.16 98.32 67.45 740.34 475.76 326.07 33.52%
MAPDP (g.) 27.95 20.13 17.54 52.44 38.23 $3$ 10.80%
PARCO (g.) 24.70 19.19 17.44 44.88 35.51 32.79 81.14 66.95 61.95 531.95 344.67 302.09 1.96%
HAM (s.) 32.84 25.45 18.12 63.59 49.12 34.90 122.77 96.32 67.01 739.77 471.09 321.50 31.02%
MAPDP (s.) 25.49 20.02 17.17 49.55 37.37 33.25 - - - - - - 7.04%
PARCO (s.) 24.27 18.71 17.00 43.36 34.82 32.19 79.23 65.83 61.12 526.52 338.78 298.71 -0.01%
Table 3: FFSP results. PARCO improves the MatNet performance and is also more than 4×4$\times$4 × faster overall.
Method FFSP20 FFSP50 FFSP100
MS Gap Time (training) MS Gap Time (training) MS Gap Time (training)
CPLEX 46.4 21.0 60s - -
CPLEX 36.6 6.4 600s - -
Random 47.8 22.9 0.1s 93.2 44 0.2s 167.2 78.0 0.3s
Shortest Job First 31.3 6.4 0.1s 57.0 7.8 0.1s 99.3 10.1 0.2s
Genetic Algorithm 30.6 5.7 25.4s 56.4 7.2 57.8s 98.7 9.5 104.5s
Particle Swarm Opt. 29.1 4.2 46.8s 55.1 5.9 92.9s 97.3 8.1 173.4s
MatNet 27.26 3.0 0.9s (2h) 51.52 0.5 2.1s (5.25h) 91.58 0.0 4.9s (28h)
PARCO (w/o CL) 27.02 2.1 0.2s (0.25h) 52.29 2.0 0.4s (1.25h) 92.15 2.0 0.9s (5.25h)
PARCO 26.47 0.0 0.2s (0.5h) 51.25 0.0 0.5s (2h) 91.55 0.0 1.1s (6.25h)
5.3 Experimental Results
Main experiments
The main experiments showcasing the performance of PARCO are in Table 1, Table 2, Table 3 for HCVRP, OMDCPDP, and FFSP, respectively. Our PARCO consistently outperforms SotA neural baselines across a variety of experiments. We additionally note another property of PARCO: unlike neural baselines in HCVRP and MAPDP in the OMDCPDP, which are trained specifically for a single size and number of agents, PARCO is a single model trained on multiple location and agent distributions at the same time; nonetheless, our single model can outperform baselines trained ad hoc on specific distributions, demonstrating our method’s flexibility.
Generalization
We additionally study the generalization performance in Table 2 on the right, which are the sizes and number of agents unseen during training in OMDCPDP. Unlike MAPDP, which is constrained to a specific number of agents, PARCO can successfully generalize to unseen sizes and $m$. Notably, PARCO can even outperform Google ORTools for larger-scale instances, demonstrating remarkable scalability.
Table 4: Effect of different ##\## of agents in FFSP.
$J\times S\times m$ Metric MatNet PARCO
$50\times 3\times 6$ Obj. 35.69 33.58
Gap 6.3% 0.0%
Duration 1.1h 10.3m
Avg. Steps 200.05 19.10
$50\times 3\times 8$ Obj. 28.59 25.62
Gap 11.6% 0.0%
Duration 1.5h 9.4m
Avg. Steps 205.95 16.80
$50\times 3\times 10$ Obj. 25.05 21.46
Gap 16.7% 0.0%
Duration 1.8h 7.5m
Avg. Steps 215.85 13.45
Effect of Communication Layers
We showcase the importance of Communication Layers in Fig. 3 (a). We benchmark 1) No Communication (only context features), 2) communication via an MLP, 3) communication via an MHA layer, and 4) Our Communication layer. Our Communication Layer consistently outperforms other methods.
Effect of Conflict Handlers
Fig. 3 (b) shows the effect of the Priority-based Conflict Handler with different priorities methods $\bm{p}$ in the OMDCPDP. We consider the following: 1) Random: the priority is chosen randomly as in Zong et al. (2022), 2) First: the priority is chosen by the agent index $k=1,\dots,m$, 3) Smallest: gives priority to the agent that has traveled the least, 4) Closest: priority is given to the closest agent and 5) High Probability: priority is given to the agent whose probability of selecting the conflicting action is the highest. High Probability can consistently outperform other methods. We note that this is also the most general and problem-agnostic since methods 2-4 are problem-specific.
Scalability
Finally, we showcase PARCO’s scalability in terms of speed in large-scale instances in Fig. 3 (c) compared to the constructive autoregressive ET (Son et al., 2024). Interestingly, PARCO’s gap grows with more agents. This is because our method can fully exploit agent parallelism and can thus be a strong candidate for large-scale real-time CO applications.
6 Conclusion
In this paper, we introduced a novel approach, PARCO - Parallel AutoRegressive Combinatorial Optimization - that utilizes parallel autoregressive (AR) policies to tackle multi-agent NCO problems. We introduced a new Communication Layer to allow for the agents to effectively coordinate their next steps during decoding, alongside a Multiple Pointer Mechanism coupled with a Priority-based Conflicts Handler that can generate feasible solutions efficiently. We validate the effectiveness of PARCO through extensive experiments with different CO problems from the routing and scheduling domain and demonstrate its competitive performance against classical heuristic and neural baselines.
While PARCO is already faster than most evaluated neural baselines and decreases the number of required decoding steps by more than 90% in FFSP50 instances with 10 agents as shown in Table 4, it still does not achieve the potential reduction from $\mathcal{O}(N)$ to $\mathcal{O}(\frac{N}{m})$ in many cases. This is mainly attributed to the fact that agents can have conflicts that we currently resolve with the priority-based conflict handler. Giving one agent precedence and forcing others to stay in their current position results in more than necessary decoding steps. In future work, we will try to address this limitation by constructing a custom loss function, which penalizes conflicts and makes agents learn to avoid them during training. Further, we plan to extend our work with learnable improvement methods to enable fast first solutions coupled with powerful local search. Given PARCO’s solution quality and efficiency, a further exciting avenue of future research is to couple our model with population-based approaches (Grinsztajn et al., 2022; Chalumeau et al., 2023; Hottung et al., 2024) that could search the latent space for diverse and better solutions.
Acknowledgements
We gratefully acknowledge OMELET for supporting us with computing. We also thank people in the AI4CO open research community who have contributed to RL4CO, which this work is based on.
Appendix
[…]
Appendix A Problem Descriptions
A.1 HCVRP
A.1.1 Problem Definition
The min-max HCVRP (Heterogeneous Capacitated Vehicle Routing Problem) consists of $m$ agents sequentially visiting customers to satisfy their demands, with constraints including each customer can be visited exactly once and the amount of demand satisfied by a single vehicle in a trip cannot exceed its capacity, which can be reloaded by going back to the depot. The goal is to minimize the makespan, i.e., the worst route.
A.1.2 Mathematical Formulation
Consider a problem with $N+1$ nodes (including $N$ customers and a depot) and $m$ vehicles. The depot is indexed as 0, and customers are indexed from 1 to $N$.
Indices
$i,j$ Node indices, where $i,j=0,\ldots,N$ (0 represents the depot)
$k$ Vehicle index, where $k=1,\ldots,m$
Parameters
$N$ Number of customer nodes (excluding depot)
$m$ Number of vehicles
$X_{i}$ Location of node $i$
$d_{i}$ Demand of node $i$ ($d_{0}=0$ for the depot)
$Q_{k}$ Capacity of vehicle $k$
$f_{k}$ Speed of vehicle $k$
$c_{ij}$ Distance between nodes $i$ and $j$
Decision Variables
$x_{ijk}$$\left\{\begin{array}[]{ll}1&\text{if vehicle }k\text{ travels directly from node }i\text{ to node }j\\ 0&\text{otherwise}\end{array}\right.$
$l_{ijk}$ Remaining load of vehicle $k$ before travelling from node $i$ to node $j$
Objective Function:
$\min\max_{k=1,\ldots,m}\left(\sum_{i=0}^{N}\sum_{j=0}^{N}\frac{c_{ij}}{f_{k}}x_{ijk}\right)$ (10)
Subject to:
$\sum_{k=1}^{m}\sum_{j=0}^{N}x_{ijk}$ $=1$ $i=1,\ldots,N$ (11)
$\sum_{i=0}^{N}x_{ijk}-\sum_{h=0}^{N}x_{jhk}$ $=0$ $j=0,\ldots,N,\;k=1,\ldots,m$ (12)
$\sum_{k=1}^{m}\sum_{i=0}^{N}l_{ijk}-\sum_{k=1}^{m}\sum_{h=0}^{N}l_{jhk}$ $=d_{j}$ $j=1,\ldots,N$ (13)
$d_{j}x_{ijk}\leq l_{ijk}$ $\leq(Q_{k}-d_{i})\cdot x_{ijk}$ $i,j=0,\ldots,N,\;k=1,\ldots,m$ (14)
$x_{ijk}$$\in\{0,1\}$ $i,j=0,\ldots,N,\;k=1,\ldots,m$ (15)
$l_{ijk}\geq 0,d_{i}$ $\geq 0$ $i,j=0,\ldots,N,\;k=1,\ldots,m$ (16)
Constraint Explanations: The formulation is subject to several constraints that define the feasible solution space. Equation (11) ensures that each customer is visited exactly once by one vehicle. The flow conservation constraint (12) guarantees that each vehicle that enters a node also leaves that node, maintaining route continuity. Demand satisfaction is enforced by constraint (13), which ensures that the difference in load before and after serving a customer equals the customer’s demand. The vehicle capacity constraint (14) ensures that the load carried by a vehicle does not exceed its capacity and is sufficient to meet the next customer’s demand.
A.2 OMDCPDP
A.2.1 Problem Definition
The OMDCPDP (Open Multi-Depot Capacitated Pickup and Delivery Problem) is a practical variant of the pickup and delivery problem in which agents have a stacking limit of orders that can be carried at any given time. Pickup and delivery locations are paired, and pickups must be visited before deliveries. Multiple agents start from different depots without returning (open). The goal is to minimize the sum of arrival times to delivery locations, i.e., minimizing the cumulative lateness.
A.2.2 Mathematical Formulation
Indices
$i,j$ Node indices, where $i,j=1,\ldots,2N$
$k$ Vehicle index, where $k=1,\ldots,m$
Sets
$P$ Set of pickup nodes, $P=\{1,\ldots,N\}$
$D$ Set of delivery nodes, $D=\{N+1,\ldots,2N\}$
Parameters
$N$ Number of pickup-delivery pairs
$m$ Number of vehicles
$c_{ij}$ Travel time between nodes $i$ and $j$
$Q_{k}$ Capacity (stacking limit) of vehicle $k$
$o_{k}$ Initial location (depot) of vehicle $k$
Decision Variables
$x_{ijk}$$\left\{\begin{array}[]{ll}1&\text{if vehicle }k\text{ travels directly from node }i\text{ to node }j\\ 0&\text{otherwise}\end{array}\right.$
$y_{ik}$$\left\{\begin{array}[]{ll}1&\text{if vehicle }k\text{ visits node }i\\ 0&\text{otherwise}\end{array}\right.$
$t_{i}$ Arrival time at node $i$
$l_{ik}$ Load of vehicle $k$ after visiting node $i$
Objective Function:
$\min\sum_{i=N+1}^{2N}t_{i}$ (17)
Subject to:
$\sum_{k=1}^{m}y_{ik}$ $=1$ $i=1,\ldots,2N$ (18)
$\sum_{j=1}^{2N}x_{o_{k},j,k}$ $=1$ $k=1,\ldots,m$ (19)
$\sum_{i=1}^{2N}x_{ijk}-\sum_{h=1}^{2N}x_{jhk}$ $=0$ $j=1,\ldots,2N,\;k=1,\ldots,m$ (20)
$y_{ik}$ $=\sum_{j=1}^{2N}x_{ijk}$ $i=1,\ldots,2N,\;k=1,\ldots,m$ (21)
$t_{i}+c_{ij}-M(1-x_{ijk})$ $\leq t_{j}$ $i,j=1,\ldots,2N,\;k=1,\ldots,m$ (22)
$t_{i}$ $\leq t_{i+N}$ $i\in P$ (23)
$l_{ik}+1-M(1-x_{ijk})$ $\leq l_{jk}$ $i\in P,j\neq i+N,k=1,\ldots,m$ (24)
$l_{ik}-1+M(1-x_{ijk})$ $\geq l_{jk}$ $i\in D,j\neq i-N,k=1,\ldots,m$ (25)
$0\leq l_{ik}$ $\leq Q_{k}$ $i=1,\ldots,2N,\;k=1,\ldots,m$ (26)
$x_{ijk},y_{ik}$$\in\{0,1\}$ $i,j=1,\ldots,2N,\;k=1,\ldots,m$ (27)
$t_{i}$ $\geq 0$ $i=1,\ldots,2N$ (28)
Constraints Explanations: Equation (18) ensures that each node is visited exactly once. Constraint (19) guarantees that each vehicle starts from its designated depot. The flow conservation constraint (20) ensures route continuity for each vehicle. Equation (21) defines the relationship between x and y variables. Time consistency is enforced by constraint (22), while (23) ensures that pickups are visited before their corresponding deliveries. Constraints (24) and (25) manage the load changes during pickup and delivery operations. Finally, the vehicle capacity constraint (26) ensures that the load never exceeds the vehicle’s stacking limit.
A.3 FFSP
A.3.1 Problem Definition
The flexible flow shop problem (FFSP) is a challenging and extensively studied optimization problem in production scheduling, involving $N$ jobs that must be processed across $i=1\dots S$ stages, each with multiple machines ($m_{i}>1$). Jobs follow a specified sequence through these stages, but within each stage, any available machine can process the job, with the key constraint that no machine can handle more than one job simultaneously. The FFSP can naturally be viewed as a multi-agent CO problem by considering each machine as an agent that constructs its own schedule. Adhering to autoregressive CO, agents construct the schedule sequentially, selecting one job (or no job) at a time. The job selected by a machine (agent) at a specific stage in the decoding process is scheduled at the earliest possible time, that is, the maximum of the time the job becomes available in the respective stage (i.e., the time the job finished on prior stages) and the machine becoming idle. The process repeats until all jobs for each stage have been scheduled, and the ultimate goal is to minimize the makespan, i.e., the total time required to complete all jobs.
A.3.2 Mathematical Formulation
We use the mathematical model outlined in Kwon et al. (2021) to define the FFSP:
Indices
$i$ Stage index
$j,l$ Job index
$k$ Machine index in each stage
Parameters
$N$ Number of jobs
$S$ Number of stages
$m_{i}$ Number of machines in stage $i$
$M$ A very large number
$p_{ijk}$ Processing time of job $j$ in stage $i$ on machine $k$
Decision variables
$C_{ij}$ Completion time of job $j$ in stage $i$
$X_{ijk}$$\left\{\begin{array}[]{ll}1&\text{if job }j\text{ is assigned to machine }k\text{ in stage }i\\ 0&\text{otherwise}\end{array}\right.$
$Y_{ilj}$$\left\{\begin{array}[]{ll}1&\text{if job }l\text{ is processed earlier than job }j\text{ in stage }i\\ 0&\text{otherwise}\end{array}\right.$
Objective:
$\min\left(\max_{j=1..n}\{C_{Sj}\}\right)$ (29)
Subject to:
$\sum_{k=1}^{m_{i}}X_{ijk}$ $=1$ $i=1,\dots,S;\;j=1,\dots,N$ (30)
$Y_{iij}$ $=0$ $i=1,\dots,S;\;j=1,\dots,N$ (31)
$\sum_{j=1}^{N}\sum_{l=1}^{N}Y_{ilj}$ $=\sum_{k=1}^{m_{i}}\max\left(\sum_{j=1}^{n}(X_{ijk}-1),0\right)$ $i=1,\dots,S$ (32)
$Y_{ilj}$ $\leq\max\left(\max_{k=1\dots m_{i}}\{X_{ijk}+X_{ilk}\}-1,0\right)$ $i=1,\dots,S;\;j,l=1,2,\dots,N$ (33)
$\sum_{l=1}^{N}Y_{ilj}$ $\leq 1$ $i=1,2,\dots,S;\;j=1,2,\dots,N$ (34)
$\sum_{j=1}^{N}Y_{ilj}$ $\leq 1$ $i=1,2,\dots,S;\;l=1,2,\dots,N$ (35)
$C_{1j}$ $\geq\sum_{k=1}^{m_{1}}p_{1jk}\cdot X_{1jk}$ $j=1,2,\dots,N$ (36)
$C_{ij}$ $\geq C_{i-1j}+\sum_{k=1}^{m_{i}}p_{ijk}\cdot X_{ijk}$ $i=2,3,\dots,S;\;j=1,2,\dots,N$ (37)
$C_{ij}+M(1-Y_{ilj})$ $\geq C_{il}+\sum_{k=1}^{m_{i}}p_{ijk}\cdot X_{ijk}$ $i=1,2,\dots,S;\;j,l=1,2,\dots,N$ (38)
Constraint Explanations: Here, the objective function Eq. 29 minimizes the makespan of the resulting schedule, that is, the completetion time of the job that finishes last. The schedule has to adhere to several constraints: First, constraint set Eq. 30 ensures that each job is assigned to exactly one machine at each stage. Constraint sets Eq. 31 through Eq. 35 define the precedence relationships between jobs within a stage. Specifically, constraint set Eq. 31 ensures that a job has no precedence relationship with itself. Constraint set Eq. 32 ensures that the total number of precedence relationships in a stage equals $N-m_{i}$ minus the number of machines with no jobs assigned. Constraint set Eq. 33 dictates that precedence relationships can only exist among jobs assigned to the same machine. Additionally, constraint sets Eq. 34 and Eq. 35 restrict a job to having at most one preceding job and one following job.
Moving on, constraint set Eq. 36 specifies that the completion time of a job in the first stage must be at least as long as its processing time in that stage. The relationship between the completion times of a job in consecutive stages is described by constraint set Eq. 37. Finally, constraint set Eq. 38 ensures that no more than one job can be processed on the same machine simultaneously.
Appendix B Experimental Details
B.1 HCVRP
B.1.1 Baselines
SISR
The Slack Induction by String Removals (SISR) approach (Christiaens and Vanden Berghe, 2020) offers a heuristic method for addressing vehicle routing problems (VRPs), focusing on simplifying the optimization process. It combines techniques for route dismantling and reconstruction, along with vehicle fleet minimization strategies. SISR is applied across various VRP scenarios, including those with specific pickup and delivery tasks. In our experiments, we adhere to the hyperparameters provided in the original paper with $\bar{c}=10,L^{\mathrm{max}}=10,\alpha=10^{-3},\beta=10^{-2},T_{0}=100,T_{f}=1,\mathrm{iter}=3\times 10^{5}\times N$.
GA
The Genetic Algorithm (GA) (Karakatič and Podgorelec, 2015) is used to address vehicle routing problems (VRPs) and other NP-hard challenges by simulating natural evolutionary processes. Particularly effective in Multi-Depot Vehicle Routing Problems (MDVRP), GA quickly generates adequate solutions with reasonable computational resources. In our experiment, we follow the same carefully tuned hyperparameters from Liu et al. (2024c) with $n=200,\mathrm{iter}=40\times N,P_{m}=0.8,P_{c}=1$.
SA
The Simulated Annealing (SA) method (İlhan, 2021) targets the capacitated vehicle routing problem (CVRP) using a population-based approach combined with crossover operators. It incorporates local search and the improved 2-opt algorithm to refine routes alongside crossover techniques to speed up convergence. In our experiment, we follow the same carefully tuned hyperparameters from Liu et al. (2024c) with $T_{0}=100,T_{f}=10^{-7},L=20\times N,\alpha=0.98$.
AM
The Attention Model (AM) (Kool et al., 2018) applies the attention mechanism to tackle combinatorial optimization problems like the Traveling Salesman and Vehicle Routing Problems. It utilizes attention layers for model improvement and trains using REINFORCE with deterministic rollouts. In our studies, we adopt adjustments from the DRLLi framework, which involves selecting vehicles sequentially and then choosing the next node for each. Additionally, vehicle-specific features are incorporated into the context vector generation to distinguish between different vehicles.
ET
The Equity-Transformer (ET) approach (Son et al., 2024) addresses large-scale min-max routing problems by employing a sequential planning approach with sequence generators like the Transformer. It focuses on equitable workload distribution among multiple agents, applying this strategy to challenges like the min-max multi-agent traveling salesman and pickup and delivery problems. In our experiments, we modify the decoder mask in ET to generate feasible solutions for HCVRP, and integrate vehicle features into both the input layer and the context encoder.
DRLLi
The DRL approach for solving HCVRP by Li et al. (2022) employs a transformer architecture similar to Kool et al. (2018) in which the vehicle and node selection happens in two steps via a two selection decoder, thus requiring two actions. We employ their original model with additional context of variable vehicle speeds, noting that in the original setting each model was trained on a single distribution of number of agents $m$, each with always the same characteristics.
2D-Ptr
The 2D Array Pointer network (2D-Ptr) (Liu et al., 2024c) addresses the heterogeneous capacitated vehicle routing problem (HCVRP) by using a dual-encoder setup to map vehicles and customer nodes effectively. This approach facilitates dynamic, real-time decision-making for route optimization. Its decoder employs a 2D array pointer for action selection, prioritizing actions over vehicles. The model is designed to adapt to changes in vehicle and customer numbers, ensuring robust performance across different scenarios. For our study, 2D-Ptr was configured following the original study’s parameters for consistent comparison with 8 heads, 128 hidden dimensions for model structure; $\mathrm{lr}=10^{-4},\lambda_{\mathrm{Adam}}=0.995,\kappa_{\mathrm{L2}}=3$ for turning; 50 epochs, 2500 batches per epoch, 512 instances per batch for training.
B.1.2 Datasets
Train data generation
Neural baselines were trained with the specific number of nodes $N$ and number of agents $m$ they were tested on. In PARCO, we select a varying size and number of customer training schemes: at each training step, we sample $N\sim\mathcal{U}(60,100)$ and $m\sim\mathcal{U}(3,7)$. As we show in the Table 1, a single PARCO model can outperform baseline models even when they were fitted on a specific distribution. The coordinates of each customer location $(x_{i},y_{i})$, where $i=1,\ldots,N$, are sampled from a uniform distribution $\mathcal{U}(0.0,1.0)$ within a two-dimensional space. The depot location is similarly sampled using the same uniform distribution. The demand $d_{i}$ for each customer $i$ is also drawn from a uniform distribution $\mathcal{U}(1,10)$, with the depot having no demand, i.e., $d_{0}=0$. Each vehicle $k$, where $k=1,\ldots,m$, is assigned a capacity $Q_{k}$ sampled from a uniform distribution $\mathcal{U}(20,41)$. The speed $f_{k}$ of each vehicle is uniformly distributed within the range $\mathcal{U}(0.5,1.0)$.
Testing
Testing is performed on the 1280 instances per test setting from Liu et al. (2024c). In Table 1, (g.) refers to the greedy performance of the model, i.e., taking a single trajectory by taking the maximum action probability; (s.) refers to sampling $1280$ solutions in the latent space and selecting the one with the lowest cost (i.e., highest reward). We report optimality gaps to ORTools and solution time in seconds in parentheses.
B.1.3 PARCO Network Hyperparameters
Encoder
Initial Embedding. This layer projects initial raw features to hidden space. For the depot, the initial embedding is the positional encoding of the depot’s location $X_{0}$. For agents, the initial embedding is the encoding for the initial location, capacity, and speed. Main Encoder. we employ $L=3$ attention layers in the encoder, with hidden dimension $d_{h}=128$, $8$ attention heads in the MHA, MLP hidden dimension set to $512$, with RMSNorm (Zhang and Sennrich, 2019) as normalization before the MHA and the MLP.
Decoder
Context Embedding. This layer projects dynamic raw features to hidden space. The context is the embedding for the depot states, current node states, current time, remaining capacities, time of backing to the depot, and number of visited nodes. These features are then employed to update multiple queries $q_{k},~{}k=1,\dots,m$ simultaneously. Main Decoder. Similarly to the encoder, we employ the same hidden dimension and number of attention heads for the Multiple Pointer Mechanism.
Communication Layer
We employ a single transformer layer with hidden dimension $d_{h}=128$, $8$ attention heads in the MHA, MLP hidden dimension set to $512$, with RMSNorm (Zhang and Sennrich, 2019) as normalization before the MHA and the MLP. Unlike the encoder layer, which acts between all $m+N$ nodes, communication layers are lighter because they communicate between $m$ agents.
Agent Handler
We use the High Probability Handler for managing conflicts: priority is given to the agent whose (log-) probability of selecting the conflicting action $a$ is the highest. In other words, priorities $\mathbf{p}_{k}=\log p_{\theta}(a_{k}|x)$ for $k=1,\dots,m$.
B.1.4 PARCO Training Hyperparameters
For each problem size, we train a single PARCO model that can effectively generalize over multiple size and agent distributions. We train PARCO with RL via SymNCO (Kim et al., 2022) with $K=10$ symmetric augmentations as shared REINFORCE baseline for $100$ epochs using the Adam optimizer (Kingma and Ba, 2014) with a total batch size $512$ (using 4 GPUs in Distributed Data Parallel configuration) and an initial learning rate of $10^{-4}$ with a step decay factor of $0.1$ after the $80$th and $95$th epochs. For each epoch, we sample $4\times 10^{5}$ randomly generated data. Training takes around 15 hours in our configuration.
B.2 OMDCPDP
B.2.1 Datasets
Train data generation
Neural baselines were trained with the specific number of nodes $2N$ and number of agents $m$ they were tested on. In PARCO, we select a varying size and number of customer training schemes: at each training step, we sample $2N\sim\mathcal{U}(50,100)$ and $m\sim\mathcal{U}(10,50)$. As we show in the Table 2, a single PARCO model can outperform baseline models even when they were fitted on a specific distribution. The coordinates of each customer location $(x_{i},y_{i})$, where $i=1,\ldots,2N$, are sampled from a uniform distribution $\mathcal{U}(0.0,1.0)$ within a two-dimensional space. Similarly, we sample $m$ initial vehicle locations from the same distribution. We set the demand $d_{i}$ for each customer to $1$ and the capacity of each vehicle to $3$. This emulates realistic settings in which a single package per customer will be picked up and delivered.
Testing
Testing is performed on 100 new instances for each setting of $N$ and $m$ in Table 2 with the distributions from the training settings. We finally test PARCO on large-scale, real-world data from complex, realistic distributions with $32$ instances for each vehicle, pickup, and delivery location in Seoul City, South Korea. This dataset is provided in our source code. Fig. 4 provides an example of such data. In Table 2, (g.) refers to the greedy performance of the model, i.e., taking a single trajectory by taking the maximum action probability; (s.) refers to sampling $1280$ solutions in the latent space and selecting the one with the lowest cost (i.e., highest reward).
Figure 4: Real-world instance for the OMDCPDP problem in Seoul City, South Korea, with $N=1000$ locations and $m=100$ agents (green) showing relations (gray) of pickups (red triangles) and their respective deliveries (black crosses).
B.2.2 Baselines
OR-Tools
The OR-Tools (Furnon and Perron, 2024) is open-source software designed to address various combinatorial optimization problems. This toolkit offers a comprehensive suite of solvers suitable for linear programming, mixed-integer programming, constraint programming, and routing and scheduling challenges. Specifically for routing problems like the OMDCPDP, OR-Tools can integrate additional constraints to enhance solution accuracy. Documentation for the Python OR-Tools library, along with example codes for diverse optimization challenges, is readily accessible on their official website. For our experiments, we maintained consistent parameters across various problem sizes and numbers of agents. We configured the global span cost coefficient to $10,000$, selected PATH_CHEAPEST_ARC as the initial solution strategy, followed by GUIDED_LOCAL_SEARCH for local optimization. The solving time was capped at 300 seconds to standardize comparison and measure efficiency.
HAM
The Heterogeneous Attention Model (HAM) (Li et al., 2021a) utilizes a neural network integrated with a heterogeneous attention mechanism that distinguishes between the roles of nodes and enforces precedence constraints, ensuring the correct sequence of pickup and delivery nodes. This approach helps the deep reinforcement learning model to make informed node selections during route planning. We used the same hyperparameter settings from the original HAM experiments.
MAPDP
The Multi-Agent Reinforcement Learning-based Framework for Cooperative Pickup and Delivery Problem (MAPDP) (Zong et al., 2022) pioneers the use of multi-agent reinforcement learning (MARL) for the cooperative PDP with multiple vehicle agents. This innovative framework introduces a centralized MARL architecture to generate cooperative decisions among agents, incorporating a paired context embedding to capture the inter-dependency of heterogeneous nodes. We adapted the MAPDP to fit our OMDCPDP task, utilizing the same encoder as PARCO to ensure a fair comparison. For the decoder and training phases, we kept the same random conflict handler and retained the same hyperparameters as detailed in the original study with 8 heads, 128 hidden dimensions for model structure; $\mathrm{lr}=10^{-3},L_{\mathrm{Adam}}=3$ for turning.
B.2.3 PARCO Network Hyperparameters
Most hyperparameters are kept similar to Section B.1.3; we report all of them nonetheless below.
Encoder
Initial Embedding. This layer projects initial raw features to hidden space. For depots, the initial embeddings encode the location $o_{k}$ and the respective vehicle’s capacity $Q_{k}$. For pickup nodes, the initial embeddings encode the location and paired delivery nodes’ location. For delivery nodes, the initial embeddings encode the location and paired pickup nodes’ location. Main Encoder. We employ $l=3$ attention layers in the encoder, with hidden dimension $d_{h}=128$, $8$ attention heads in the MHA, MLP hidden dimension set to $512$, with RMSNorm (Zhang and Sennrich, 2019) as normalization before the MHA and the MLP.
Decoder
Context Embedding. This layer projects dynamic raw features to hidden space. The context is the embedding for the depot states $o_{k}$, current node states, current length, remaining capacity, and number of visited nodes. These features are then employed to update multiple queries $q_{k},~{}k=1,\dots,m$ simultaneously. Main Decoder. Similarly to the encoder, we employ the same hidden dimension and number of attention heads for the Multiple Pointer Mechanism.
Communication Layer
We employ a single transformer layer with hidden dimension $d_{h}=128$, $8$ attention heads in the MHA, MLP hidden dimension set to $512$, with RMSNorm (Zhang and Sennrich, 2019) as normalization before the MHA and the MLP. Note that unlike the encoder layer, which acts between all $m+N$ nodes, communication layers are lighter because they communicate between $m$ agents.
Agent Handler
We use the High Probability Handler for managing conflicts: priority is given to the agent whose (log-) probability of selecting the conflicting action $a$ is the highest. In other words, priorities $\mathbf{p}_{k}=\log p_{\theta}(a_{k}|x)$ for $k=1,\dots,m$.
B.2.4 PARCO Training Hyperparameters
For each problem size, we train a single PARCO model that can effectively generalize over multiple size and agent distributions. We train PARCO with RL via SymNCO (Kim et al., 2022) with $K=8$ symmetric augmentations as shared REINFORCE baseline for $100$ epochs using the Adam optimizer (Kingma and Ba, 2014) with a total batch size $128$ on a single GPU and an initial learning rate of $10^{-4}$ with a step decay factor of $0.1$ after the $80$th and $95$th epochs. For each epoch, we sample $10^{5}$ randomly generated data. Training takes less than 4 hours in our configuration.
B.3 FFSP
B.3.1 Baselines
MatNet
The comparison of PARCO with other FFSP baselines is shown in table Table 3. Being a neural solver, we benchmark PARCO mainly against MatNet (Kwon et al., 2021), the SotA NCO architecture for the FFSP. MatNet is an encoder-decoder architecture, which is inspired by the attention model (Kool et al., 2018). It extends the encoder of the attention model with a dual graph attention layer, a horizontal stack of two transformer blocks, capable to encode nodes of different types in a bipartite graph - like machines and jobs in the FFSP. Kwon et al. (2021) train MatNet using POMO (Kwon et al., 2020).
CPLEX
Besides the aforementioned MatNet model, we implement the mathematical model described above in the exact solver CPLEX with a time budget of 60 and 600 seconds per instance. However, with both time budgets, CPLEX is only capable of generating solutions to the FFSP20 instances.
Random and Shortest Job First
Third, we use different (meta-)heuristics to benchmark PARCO on the FFSP. The Random and Shortest Job First (SJF) heuristics are simple construction strategies that build valid schedules in an iterative manner. Starting from an empty schedule, the Random construction heuristic iterates through time steps $t=0,\dots T$ and stages $i=1\dots S$ and randomly assigns jobs available at the given time to an idle machine of the respective stage until all jobs are scheduled. Likewise, the SJF proceeds by assigning job-machine pairs with the shortest processing time first.
Genetic Algorithm
The Genetic Algorithms (GA) are metaheuristics widely used by the OR community to tackle the FFSP (Kahraman et al., 2008). The GA iteratively improves multiple candidate solutions called chromosomes. Each chromosome consists of $S\times N$ real numbers, where $S$ is the number of stages and $N$ is the number of jobs. For each job at each stage, the integer part of the corresponding number indicates the assigned machine index, while the fractional part determines job priority when multiple jobs are available simultaneously. Child chromosomes are created through crossover, inheriting integer and fractional parts independently from two parents. Mutations, applied with a 30% chance, use one of four randomly selected methods: exchange, inverse, insert, or change. The implementation uses 25 chromosomes. One initial chromosome is set to the Shortest Job First (SJF) heuristic solution and the best-performing chromosome is preserved across iterations. Each instance runs for 1,000 iterations.
Particle Swarm Optimization
Finally, Particle Swarm Optimization (PSO) iteratively updates multiple candidate solutions called particles, which are updated by the weighted sum of the inertial value, the local best, and the global best at each iteration (Singh and Mahapatra, 2012). In this implementation, particles use the same representation as GA chromosomes. The algorithm employs 25 particles, with an inertial weight of 0.7 and cognitive and social constants set to 1.5. One initial particle represents the SJF heuristic solution. Like GA, PSO runs for 1,000 iterations per instance.
B.3.2 Datasets
Train data generation
We follow the instance generation scheme outlined in Kwon et al. (2021) sample processing times for job-machine pairs independently from a uniform distribution within the bounds $[2,10]$. For the main experiments on the FFSP shown in Table 3, we also use the same instance sizes as Kwon et al. (2021) with $N=20,50\text{ and }100$ jobs (denoted FFSP20, FFSP50 and FFSP100, respectively) that need to be processed in $S=3$ stages on $m_{i}=4$ machines each (i.e. $m=12$ machines in total). For the agent sensitivity analysis shown in Table 4, we fix the number of jobs to 50 but alter the number of agents. Still, we use $S=3$ for this experiment, but alter the number of machines per stage to $m_{i}=6,8\text{ and }10$, yielding a total of 18, 24 and 30 agents, respectively.
Testing
Testing is performed on 1,000 test instances generated according to the above sampling scheme. We use the same instances as Kwon et al. (2021) for the experiments of Table 3 and generate new ones for the experiments of Table 4.
B.3.3 PARCO Network Hyperparameters
Encoder
To solve the FFSP with our PARCO method, we use a similar encoder as Kwon et al. (2021). The MatNet encoder generates embeddings for all machines of all stages and the jobs they need to process, plus an additional dummy job embedding, which can be selected by any machine in each decoding step to skip to the next step. To compare PARCO with MatNet, we use similar hyperparameters for both models. We use $L=3$ encoder layers, generating embeddings of dimensionality $d_{h}=256$, which are split over $h=16$ attention heads in the MHA layers. Further, we employ Instance Normalization (Ulyanov et al., 2016) and a feed-forward network with 512 neurons in the transformer blocks of the encoder.
Decoder
The machines are regarded as the agents in our PARCO framework. As such, their embeddings are used as queries $Q_{t}^{s}$ in the Multiple Pointer Mechanism Eq. 5, while job embeddings are used as the keys and values. In each decoding step, the machine embeddings are fused with a projection of the time the respective machine becomes idle. Similarly, job embeddings are augmented with a linear transformation of the time they become available in the respective stage before entering the attention head in Eq. 4.
Communication Layer
We employ a single transformer block with hidden dimension $d_{h}=256$ and $h=16$ attention heads in the MHA, an MLP with $512$ hidden units and Instance Normalization.
Agent Handler
We use the High Probability Handler for managing conflicts: priority is given to the agent whose (log-) probability of selecting the conflicting action is the highest. Formally, priorities $\mathbf{p}_{k}=\log p_{\theta}(a_{k}|x)$ for $k=1,\dots,m$.
B.3.4 PARCO Training Hyperparameters
Regarding the training setup, each training instance $i$ is augmented by a factor of 24, and the average makespan over the augmented instances is used as a shared baseline $b^{\text{shared}}_{i}$ for the REINFORCE gradient estimator of Eq. 9. We use the ADAM optimizer (Kingma and Ba, 2014) with a learning rate of $4\times 10^{-4}$, which we alter during training using a cosine annealing scheme. We train separate models for FFSP20, FFSP50 and FFSP100 instances for 100, 150 and 200 epochs, respectively, using 1,000 randomly generated instances per epoch split into batches of size 50.111Note: to avoid OOMs, For FFSP100 instances, batches are further split into mini-batches of size 25 whose gradients are accumulated. For the learned methods, the times given in parenthesis indicate the time required to train the respective models, whereas the other time information indicates the time required to solve a single instance. For the agent sensitivity analysis in Table 4, Matnet, as well as our PARCO model, are trained for 20 epochs on 1,000 instances per epoch. Duration in this table indicates the time required for training the respective models on the hardware specified in Section B.4.
B.3.5 Diagram for MatNet Decoding vs. PARCO Decoding for the FFSP
The following figures visualize the decoding for the machines of a given stage using MatNet and PARCO. As one can see in Fig. 6, MatNet requires a decoder forward pass for each machine to schedule a job on each of them. In contrast, as detailed in Fig. 6, PARCO can schedule jobs on all machines simultaneously through its Multiple Pointer Mechanism and Agent Handler, leading to significant efficiency gains.
Figure 5: A FFSP Decoding Step with MatNet
Figure 6: A FFSP Decoding Step with PARCO
B.4 Hardware and Software
B.4.1 Hardware
Experiments were carried out on a machine equipped with two Intel(R) Xeon(R) Gold 6338 CPU @ 2.00GHz CPUs with a total $128$ threads and $8$ NVIDIA RTX 4090 graphic cards with $24$ GB of VRAM. During inference, we employ only one CPU and a single GPU.
B.4.2 Software
We used Python 3.12, PyTorch 2.4 (Paszke et al., 2019) coupled with PyTorch Lightning (Falcon and The PyTorch Lightning team, 2019) with most code based on the RL4CO library (Berto et al., 2024a). The main OS is Ubuntu 22.04.4 LTS. There were no major difficulties when trying the code on other platforms, so we expect experiments to be reproducible in other architectures.
B.5 Source Code
We value open reproducibility. We provide the source code to reproduce our experiments on Github at https://github.com/ai4co/parco.
Appendix C Additional Discussion
Following Section 6, we outline further discussion here. We additionally remark that several areas of multi-agent planning, including multi-agent pathfinding (Lin and Ma, 2023; Wang et al., 2023; Tang et al., 2024a) and social robot navigation (Li et al., 2024a; He et al., 2024) employ communication as a tool with great importance, from which we also borrow ideas such as prioritized conflict resolution (Tang et al., 2024b).
An exciting future direction going forward is the development of foundation models, in which PARCO could aid in improving the efficiency and performance in solving multi-agent CO problems. Recent examples in the literature have demonstrated the possibility of scaling up solvers with (self)supervision (Drakulic et al., 2023; Luo et al., 2023; Corsini et al., 2024; Pirnay and Grimm, 2024; Luo et al., 2024), developing multi-task solvers for VRPs (Liu et al., 2024a; Lin et al., 2024; Zhou et al., 2024; Berto et al., 2024b) as well as more general CO tasks (Drakulic et al., 2024; Jiang et al., 2024). Such approaches model CO problems as sequential autoregressive decision-making; parallelizing this process with PARCO could enrich solution speed/quality. We note that most NCO approaches in e.g., VRPs, mostly focus on solving without explicitly modeling (heterogeneous) agents as in PARCO, which is of practical relevance in several real-world tasks. Better heuristics for prioritized planning might also be found utilizing automated discovery via recent LLM approaches for CO (Liu et al., 2024b; Ye et al., 2024b). Finally, integrating PARCO in online optimization (Hottung et al., 2021; Choo et al., 2022) and end-to-end pipelines for construction and improvement, as done recently in (Berto et al., 2024a; Kong et al., 2024), would further improve the performance of PARCO given larger computational time budgets.

### v2

@@
-PARCO: Learning Parallel Autoregressive Policies for Efficient Multi-Agent Combinatorial Optimization
+Parallel AutoRegressive Models for Multi-Agent Combinatorial Optimization
 Abstract
-Multi-agent combinatorial optimization problems such as routing and scheduling have great practical relevance but present challenges due to their NP-hard combinatorial nature, hard constraints on the number of possible agents, and hard-to-optimize objective functions.
-This paper introduces PARCO (Parallel AutoRegressive Combinatorial Optimization), a novel approach that learns fast surrogate solvers for multi-agent combinatorial problems with reinforcement learning by employing parallel autoregressive decoding.
-We propose a model with a Multiple Pointer Mechanism to efficiently decode multiple decisions simultaneously by different agents, enhanced by a Priority-based Conflict Handling scheme.
-Moreover, we design specialized Communication Layers that enable effective agent collaboration, thus enriching decision-making.
-We evaluate PARCO in representative multi-agent combinatorial problems in routing and scheduling and demonstrate that our learned solvers offer competitive results against both classical and neural baselines in terms of both solution quality and speed.
-We make our code openly available at https://github.com/ai4co/parco.
-[…]
+Combinatorial optimization problems involving multiple agents are notoriously challenging due to their NP-hard nature and the necessity for effective agent coordination.
+Despite advancements in learning-based methods, existing approaches often face critical limitations, including suboptimal agent coordination, poor generalizability, and high computational latency.
+To address these issues, we propose Parallel AutoRegressive Combinatorial Optimization (PARCO), a reinforcement learning framework designed to construct high-quality solutions for multi-agent combinatorial tasks efficiently.
+To this end, PARCO integrates three key components: (1) transformer-based communication layers to enable effective agent collaboration during parallel solution construction, (2) a multiple pointer mechanism for low-latency, parallel agent decision-making, and (3) priority-based conflict handlers to resolve decision conflicts via learned priorities.
+We evaluate PARCO in multi-agent vehicle routing and scheduling problems where our approach outperforms state-of-the-art learning methods and demonstrates strong generalization ability and remarkable computational efficiency.
+Machine Learning, ICML
 1 Introduction
-Combinatorial optimization (CO) problems, such as routing and scheduling problems, involve determining an optimal sequence of actions in a combinatorial space and have applications ranging from warehouse operatons (Xie et al., 2023) to network design (Chabarek et al., 2008) and safety-critical systems (Girardey et al., 2010).
-CO problems are notoriously hard to solve and cannot generally be solved optimally in polynomial time, i.e., they are NP-hard (Jünger et al., 1995).
-Multi-agent settings gained significant interest due to their applicability in realistic scenarios, such as path finding (Reijnen et al., 2020), drone routing (Ann et al., 2015), disaster management (Bektas, 2006; Cheikhrouhou and Khoufi, 2021) and order delivery (Yakıcı and Karasakal, 2013; Archetti and Bertazzi, 2021), but present even more challenges due to additional constraints and different optimization objectives, including minimizing a global lateness objective or the makespan (Mahmoudinazlou and Kwon, 2024).
-While traditional algorithmic methods have significantly contributed to solving a range of problems (Laporte and Osman, 1995; Hejazi and Saghafian, 2005), these approaches often concentrate on single-agent scenarios.
-With the advent of modern computational techniques, neural network-based approaches have started to yield promising results for complex CO problems in a field known as Neural Combinatorial Optimization (NCO).
-In particular, Reinforcement Learning (RL) has shown promise due to its ability to learn directly from interactions with environments instead of relying on (costly) labeled datasets in the shape of optimal solutions (Bello et al., 2016; Kwon et al., 2020; Kim et al., 2023a).
-Among NCO methods, Autoregressive (AR) sequence generation has gained attention for its ability to manage hard constraints (Kool et al., 2018; Kwon et al., 2021).
-In the context of NCO, this capability is crucial for addressing complex problems with multiple construction constraints, such as the precedence constraints prevalent in scheduling (Zhang et al., 2020a) as well as pickup and delivery problems (Savelsbergh and Sol, 1995; Parragh et al., 2008).
-However, one issue of autoregressive sequence generation is the high latency associated with it, especially when considering large problem instances.
-The problem of high inference latency of AR methods is especially prevalent in the domain of large language models, as multiple stacks of transformer layers have to be computed sequentially for each individual token (Bae et al., 2023).
-Figure 1: Solution construction with Autoregressive (top) and Parallel Autoregressive policies (bottom).
-Motivated by recent studies on LLMs demonstrating parallel decoding can not only tackle the high generation latency problem of next-token predictors but also improve their answer quality (Qi et al., 2020; Ning et al., 2023; Gloeckle et al., 2024) this paper introduces PARCO (Parallel AutoRegressive Combinatorial Optimization), a novel approach designed to address multi-agent combinatorial problems efficiently.
-We propose to apply parallel decoding in the NCO domain to increase solution construction efficiency and effectiveness.
-In contrast to most previous NCO methods, PARCO constructs solutions in parallel for multiple agents, made feasible by a priority-based conflict handler to manage conflicting decisions of different agents.
-Importantly, we enable the collaborative behavior of agents through specialized inter-agent Communication Layers, essential for finding high-quality solutions to multi-agent combinatorial problems.
-Contributions.
-We summarize our contributions as follows:
-  • We introduce PARCO, a novel approach for tackling multi-agent CO problems by effectively constructing solutions in parallel via a Multiple Pointer Mechanism.
-  • We introduce an efficient Priority-based Conflict-Handling scheme to resolve conflict between agents and enhance decision-making.
-  • We design a special Communication Layer to help agents communicate and collaborate to obtain a better state-space representation of partial solutions.
-  • We show competitive performance against classical solvers and neural baselines in routing and scheduling problems, both in terms of solution quality and computational efficiency.
+Combinatorial optimization (CO) problems involve determining an optimal sequence of actions in discrete spaces with several crucial domains, including logistics and supply chain management (Weinand et al., 2022).
+Many practical CO problems require a solution to be constructed by coordinating multiple distinct entities (i.e., agents), each with unique characteristics.
+We call such problems multi-agent CO.
+This class of problems naturally arises in real-world applications such as coordinated vehicle routing for disaster management (Liberatore et al., 2014), manufacturing (Jungwattanakit et al., 2008) and last-mile delivery optimization (Archetti & Bertazzi, 2021), where heterogeneous agents must operate under complex constraints.
+CO problems are notoriously hard to solve and cannot generally be solved optimally in polynomial time due to their NP-hardness (Jünger et al., 1995; Peres & Castelli, 2021).
+While traditional methods such as exact and heuristic methods have been developed to solve a variety of problems (Laporte & Osman, 1995; Hejazi & Saghafian, 2005), these approaches often concentrate on single-agent scenarios and require long execution times.
+Moreover, multi-agent CO poses numerous additional challenges such as additional constraints and different optimization objectives including minimizing lateness or the makespan among agents (Soman & Patil, 2020; Mahmoudinazlou & Kwon, 2024).
+Recently, Neural Combinatorial Optimization (NCO) has emerged as a powerful alternative by learning efficient neural solvers (Bengio et al., 2021).
+In particular, Reinforcement Learning (RL) has shown promise due to its ability to learn directly from interactions with CO problems instead of relying on costly labeled datasets in the shape of optimal solutions, and could even outperform traditional approaches by automatically discovering better (neural) heuristics (Bello et al., 2016; Kwon et al., 2020; Kim et al., 2023a).
+Among NCO methods, Autoregressive (AR) models—constructing solutions step-by-step—have garnered attention for their ability to generate solutions for a variety of problems with hard constraints (Kool et al., 2019a; Kwon et al., 2021; Bi et al., 2024).
+This capability is crucial for addressing complex problems with multiple constraints, such as heterogeneous capacity (Baldacci et al., 2008) or precedences in pickup and delivery (Parragh et al., 2008) and machine scheduling problems (Zhang et al., 2020a).
+However, despite their promising results in single-agent settings, AR-CO methods pose two challenges that hinder their practical adoption in solving complex multi-agent CO problems.
+First, existing AR methods for multi-agent CO suffer from poor coordination, resulting in unsatisfactory solution quality and poor generalization across varying problem sizes and agent configurations (Zong et al., 2022; Son et al., 2024; Zheng et al., 2024a; Liu et al., 2024b).
+Moreover, AR sequence generation is associated with high latency due to each single action (or “token”) depending on each single previous one, akin to large language models (Leviathan et al., 2023; Gao et al., 2024).
+This issue becomes increasingly pronounced when dealing with large problem instances involving numerous agents.
+Figure 1: Example multi-agent vehicle routing problem.
+[Top] Autoregressive policies construct solutions one agent at a time, with inefficient coordination and requiring several steps.
+[Bottom] Our parallel autoregressive policies ensure not only improved decision-making thanks to inter-agent communication that enables effective coordination but also faster construction in fewer steps.
+This paper introduces PARCO (Parallel AutoRegressive Combinatorial Optimization), a novel learning framework to address multi-agent combinatorial problems effectively via parallel solution construction as illustrated in Figure 1.
+We design specialized transformer-based Communication Layers to enhance coordination among agents during decision-making, enabling collaborative behaviors.
+Our model leverages a Multiple Pointer Mechanism to efficiently generate actions for various agents simultaneously at each solution construction step.
+Priority-based Conflict Handlers ensure feasibility and resolve potential conflicts among agents’ decisions based on learned priorities.
+We summarize PARCO’s contributions as follows:
+  • We propose a general Parallel Autoregressive framework for solution construction in multi-agent CO.
+  • We introduce Communication Layers to enhance agent coordination at each parallel construction step.
+  • We design a Multiple Pointer Mechanism that reduces latency by efficiently decoding solutions in parallel.
+  • We enhance solution quality with Priority-based Conflict Handlers that tie break with learned priorities.
+  • We evaluate PARCO on multi-agent vehicle routing and scheduling, where we outperform state-of-the-art learning methods across diverse problems [^1] Code at: https://github.com/ai4co/parco.
 2 Related Work
 Neural Combinatorial Optimization
-Recent advancements in NCO have shown promising end-to-end solutions for combinatorial optimization problems, as highlighted by Bengio et al.
-(2021) and Yang and Whinston (2023).
-NCO has led to the development of aligned neural architectures (Bi et al., 2022; Jin et al., 2023; Luo et al., 2023; Kim et al., 2023c), hybrid methods with OR solvers (Li et al., 2021b; Kim et al., 2024a; Yan and Wu, 2024; Ye et al., 2024a; Kim et al., 2024b), multi-level solution pipelines (Ma et al., 2023a; Li et al., 2023; Xiao et al., 2023; Ye et al., 2023), alongside improved training algorithms (Kim et al., 2023c; Jiang et al., 2023; Drakulic et al., 2023; Sun and Yang, 2023; Gao et al., 2023; Xiao et al., 2023; Li et al., 2024b; Wang et al., 2024) to enhance the heuristic search.
-These innovations have expanded NCO’s application across a wide array of problems (Chen et al., 2023; Kim et al., 2023b; Zhou et al., 2023; Ma et al., 2023b; Luttmann and Xie, 2024).
-However, integrating appropriate inductive biases requires manual tuning of model architectures and training algorithms, presenting challenges such as computationally intensive training, the need for specialized hardware, and issues with interpretability and generalizability (Liu et al., 2023).
-Multi-agent Constructive Methods for NCO
-While several seminal works as Vinyals et al.
+Recent advancements in Neural Combinatorial Optimization (NCO) have shown promising end-to-end solutions for combinatorial optimization problems (Bengio et al., 2021; Yang & Whinston, 2023).
+NCO has led to the development of a variety of methods for diverse problems, including the incorporation of problem-specific biases (Jin et al., 2023; Luo et al., 2023; Kim et al., 2023c, b; Son et al., 2023; Gao et al., 2023), bi-level solution pipelines (Li et al., 2021b; Ye et al., 2023b; Zheng et al., 2024b), learning-guided search (Ye et al., 2023a; Kim et al., 2024; Sun & Yang, 2023; Li et al., 2024b, 2025), improvement methods (Hottung et al., 2021; Ma et al., 2021, 2024), effective training algorithms (Kim et al., 2022; Grinsztajn et al., 2023; Drakulic et al., 2023; Corsini et al., 2024; Pirnay & Grimm, 2024; Luo et al., 2024), downstream applications (Chen et al., 2023; Zhou et al., 2023; Ma et al., 2023; Luttmann & Xie, 2024; Yan & Wu, 2024), and the recent development of end-to-end foundation models (Liu et al., 2024a; Zhou et al., 2024; Drakulic et al., 2024; Berto et al., 2024b; Jiang et al., 2024; Li et al., 2024a).
+Among such methods, RL-based end-to-end autoregressive (AR) models present several advantages, including eliminating the need for labeled solutions, reducing reliance on handcrafted heuristics, and achieving high efficiency in generating high-quality solutions (Kool et al., 2019a; Kwon et al., 2020, 2021; Berto et al., 2024a; Bonnet et al., 2024).
+Multi-Agent AR Methods for CO
+While the seminal works in AR-CO methods of Vinyals et al.
 (2015); Kool et al.
-(2018); Kwon et al.
-(2020) propose models that can be used in loose multi-agent settings, such methods cannot be employed directly to model heterogeneous agents with different attributes, such as different capacities or starting locations in routing problems.
-Zhang et al.
-(2020b); Falkner and Schmidt-Thieme (2020); Li et al.
-(2022); Liu et al.
-(2024c) introduce models to construct solutions of different agents simultaneously.
-However, the above parallel construction is, in fact, sequential in terms of the model and environment stepping itself, thus not benefitting from true model-side parallel decoding and often lacking in agent collaboration.
-On the other hand, Son et al.
-(2024); Zheng et al.
-(2024) propose to model multi-agent min-max routing as sequential decision-making, allowing for switching the agents after a single-agent solution is complete.
-Most related to PARCO is Zong et al.
-(2022), who propose a multi-agent model for the min-sum multi-agent PDP with different decoders.
-However, this approach is not only tailored to a specific CO problem but also a specific number of agents due to both different decoders for each agent and a non-flexible context embedding, which is dependent on the agent number, while PARCO is more flexible and has a more general communication representation.
-Moreover, while Zong et al.
-(2022) manages conflicts by randomly giving precedence to one agent, PARCO learns how to prioritize agents with higher probabilities of selecting a specific node.
-3 Background
+(2019a); Kwon et al.
+(2020, 2021) propose models that can be used in loose multi-agent settings, such methods cannot be employed directly to model heterogeneous agents with different attributes and constraints.
+Building on Kool et al.
+(2019a), Son et al.
+(2024) and Zheng et al.
+(2024a) introduce attention-based policies for multi-agent min-max routing.
+These models adopt a sequential AR construction strategy, solving for one agent at a time and switching agents only after completing a single-agent solution.
+While this approach outperforms decentralized methods (Cao et al., 2022; Park et al., 2023) it remains inherently sequential.
+In contrast, other multi-agent AR methods determine a location for every agent in each solution construction step, but the different agents select their actions sequentially in either random (Zhang et al., 2020b) or learned order (Falkner & Schmidt-Thieme, 2020; Li et al., 2022; Liu et al., 2024b), making these models still suffer from high generation latency as well as missing inter-agent communication and coordination.
+Zong et al.
+(2022) propose a multi-agent pickup and delivery model with parallel decoding, using distinct decoders for each agent.
+However, this approach exhibits limited generalizability due to inflexible fixed decoders for specific agents, lacking a powerful communication mechanism, and conflict resolution handled naively by assigning random precedence to agents, restricting the robustness of the model.
+PARCO aims at addressing the shortcomings of previous works by leveraging parallel solution construction for any number of agents efficiently with a Multiple Pointer Mechanism, enhancing coordination via Communication Layers, and solving conflicts in a principled manner via Priority-based Conflict handlers.
+Finally, unlike previous works, PARCO tackles a general framework tackling multi-agent CO without restricting to a single class of problems.
+3 Preliminaries
 3.1 Markov Decision Processes
-CO problems can be framed as Markov Decision Processes (MDPs).
-In this formulation, the problem is defined by a set of states $\mathcal{S}$, where each state $s_{t}\in\mathcal{S}$ represents the configuration of the problem at time step $t$.
-At each step, an agent selects an action $a_{t}$ from the action space $\mathcal{A}$ according to a policy $\pi:\mathcal{S}\rightarrow\mathcal{A}$, which maps states to actions.
-The system then transitions from state $s_{t}$ to state $s_{t+1}$ according to a transition function $T:\mathcal{S}\times\mathcal{A}\rightarrow\mathcal{S}$.
-The agent reaches the terminal state once it has generated a viable solution for the problem instance $\bm{x}$.
-Typically, a reward $R$ is only obtained in the terminal state and takes the form of the objective function of the respective CO problem.
-3.2 Autoregressive Methods for NCO
-Given the sequential nature of MDPs, autoregressive (AR) methods pose a natural choice for the policy $\pi$.
+Multi-agent CO problems can be formulated as Markov Decision Processes (MDPs) and solved autoregressively using RL (Mazyavkina et al., 2021).
+In this framework, a solution $\bm{a}$ to a CO problem instance $\bm{x}$ is represented as a sequence of actions.
+Actions $a_{t}$ are selected sequentially from the action space $\mathcal{A}$ of size $N$ based on the current state $s_{t}\in\mathcal{S}$, which encodes the problem’s configuration at step $t$.
+In multi-agent problems, at each step one agent $m\in\mathcal{M}=\{1,\dots,M\}$ selects an action $a^{m}_{t}$ according to a policy $\pi_{\theta}$, usually represented by a $\theta$-parametrized neural network, mapping states to actions.
+Agents are selected either by some predefined precedence rule as in the sequential planning of Son et al.
+(2024), where agent solutions are constructed one after another, or by the policy itself, in which case $\pi_{\theta}:\mathcal{S}\rightarrow\mathcal{A}\times\mathcal{M}$.
+Given the agent and its corresponding action, the problem then transitions from state $s_{t}$ to state $s_{t+1}$ according to a transition function $\tau:\mathcal{S}\times\mathcal{A}\times\mathcal{M}\rightarrow\mathcal{S}$.
+This process reaches the terminal state once it has generated a feasible solution $\bm{a}=(a_{1},...,a_{T})$ for the problem instance $\bm{x}$ in $T$ construction steps.
+The (sparse) reward $R(\bm{a},\bm{x})$ is usually obtained only in the terminal state and takes the form of the negative of the cost function of the respective CO problem.
+3.2 AR Models for CO
+Given the sequential nature of MDPs, AR methods pose a natural choice for the policy $\pi_{\theta}$.
 AR methods construct a viable solution by sequentially generating actions based on the current state and previously selected actions.
-Formally, the process can be represented as:
-$p_{\theta}(\bm{a}|\bm{x})$ $\triangleq\prod_{t=1}^{T}g_{\theta}(a_{t}|a_{t-1},...,a_{0},\bm{h})$ (1)
-where $\bm{h}=f_{\theta}(\bm{x})$ is the encoding of problem instance $\bm{x}$ via encoder network $f$, which is used to decode actions autoregressively via decoder $g_{\theta}$.
-$p_{\theta}$ is a solver that maps $\bm{x}$ to a solution $\bm{a}$.
-$\bm{a}=(a_{1},...,a_{T})$ represents the (optimal) actions executed in $T$ construction steps, resulting in a feasible solution to CO problems.
-Notably, the same AR scheme has been applied to solve multi-agent CO problems, such as vehicle routing problems (Kool et al., 2018; Kwon et al., 2020), in which a single agent tasked with visiting all nodes under specific route length constraints, distinguishing it from a real multi-agent scenario by focusing on optimizing a solitary agent’s path within set limits.
-To tackle such problems, AR methods sequentially construct multiple routes as if handling a single agent at a time.
-However, inducing collaborative behavior among agents using the AR method poses a challenge, as it inherently focuses on one agent at a time.
-This limitation restricts the method’s ability to directly foster interactions among the agents, which is crucial for optimizing collective outcomes in multi-agent scenarios.
-3.3 Training AR Models via Reinforcement Learning
-In this work, we focus on Reinforcement Learning (RL) solvers for training $p_{\theta}$ as they can be trained without relying on (often hard-to-obtain) labeled solutions.
-We can thus cast the optimization problem as follows:
+Without loss of generality, the process can be represented in an encoder-decoder framework as:
+$p_{\theta}(\bm{a}|\bm{x})$ $\triangleq\prod_{t=1}^{T}g_{\theta}(a_{t}|\bm{a}_{<t},\bm{h})$ (1)
+where $\bm{a}_{<t}=(a_{1},\ldots,a_{t-1})$ is the sequence of actions taken prior to $t$ and $\bm{h}=f_{\theta}(\bm{x})$ is an encoding of the problem instance $\bm{x}$ obtained via the encoder network $f$.
+The decoder $g_{\theta}$ then autoregressively generates the sequence of actions, conditioned on $\bm{h}$ and the previously generated actions.
+The parameters $\theta$ encompass both the encoder and decoder components, which together define the policy as the mechanism for producing the joint distribution $p_{\theta}(\bm{a}|\bm{x})$.
+3.3 RL Training
+RL leverages the MDP formulation to learn the parameters $\theta$ of the policy $\pi_{\theta}$ without relying on labeled optimal solutions. We can thus cast the optimization problem as follows:
 […]
-where $P(\bm{x})$ is problem distribution, $R(\bm{a},\bm{x})$ is reward of $\bm{a}$ given $\bm{x}$.
-We can solve Eq.
-2 by employing various RL methods such as variations of policy gradients (REINFORCE) (Kwon et al., 2020; Kim et al., 2022).
+where $P(\bm{x})$ is the distribution over problem instances and $R(\bm{a},\bm{x})$ is the reward of a solution.
+We can solve the optimization problem in Equation 2 using various RL methods, including policy gradient (REINFORCE) variations (Kool et al., 2019b; Kwon et al., 2020; Kim et al., 2022).
 4 Methodology
-4.1 Parallel Multi-Agent Environments
-We propose to step multiple actions over the environment simultaneously to enhance efficiency in practice, which can reduce the number of steps required compared to single-agent stepping.
-In the general multi-agent CO setting, agents $k=1,...,m$ are selecting actions $\bm{a}_{t}=(a_{t}^{1},...,a_{t}^{m})$ from a shared action space $\mathcal{A}$ in parallel at a given decoding step $t$.
-Given the agent actions $\bm{a}_{t}$, the state of the problem instance transitions from $s_{t}$ to $s_{t+1}$ via some transition function $T:\mathcal{S}\times\mathcal{A}_{1}\times...\times\mathcal{A}_{m}\rightarrow\mathcal{S}$, which usually follows deterministic rules in case of non-stochastic CO problems.
-After reaching the terminal state, the agents receive a shared reward $R(A,\bm{x})$, with $\bm{x}$ the problem instance and $A=(\bm{a}_{1},...,\bm{a}_{T})$ the sequence of agents actions, which is the objective of the respective CO problem.
-Fig.
-1 illustrates an example of AR and Parallel AR solution construction.
-In this work, we propose a model architecture, PARCO, to solve such cooperative multi-agent CO problems efficiently.
-Drawing on studies on multi-agent RL (MARL) which promotes parameter sharing in cooperative MARL settings with a shared action space (Yu et al., 2022), PARCO establishes a central auto-regressive policy $\pi_{\theta}:\mathcal{S}\rightarrow P(\mathcal{A}_{1}\times...\times\mathcal{A}_{m})$ to decode the next action for the $m$ agents in parallel.
-This not only makes PARCO agnostic to the number of agents prevalent in the respective CO problem, but also enhances both solution quality and construction speed.
-4.2 Parallel Autoregressive Model
-In this section, we present our general architecture to solve CO problems efficiently cast as multi-agent sequential decision-making problems.
-PARCO follows the general encoder-decoder architecture prevalent in autoregressive NCO (Kool et al., 2018; Bello et al., 2016; Vinyals et al., 2015).
-Figure 2: Overview of the PARCO model.
-Encoder
-The job of the encoder in autoregressive NCO approaches is to learn a mapping from the graph representing the CO problem to a higher dimensional embedding space.
-This typically involves a problem depending on the initial embedding layer, projecting the graph nodes using their feature representations.
-This is also the case for PARCO, whose encoder first maps features of the graph instance to an embedding space to learn deep relationships between the nodes through a stack of transformer blocks, similar to Kool et al.
+We now outline the general structure of PARCO as shown in Figure 2.
+First, we formally define parallel multi-agent MDPs for CO (Section 4.1), which we use as a basis to derive our overall parallel autoregressive approach (Section 4.2).
+We then describe in detail the components of our model: Multi-Agent Encoder (Section 4.3), Communication Layers (Section 4.4), Decoder with Multiple Pointer Mechanism (Section 4.5) and Conflict Handlers (Section 4.6).
+Finally, we outline the training scheme (Section 4.7).
+Figure 2: Overview of PARCO.
+Our model encodes multi-agent CO problems into separate agent and node embeddings.
+Communication Layers allow for coordination among agents during decoding, which enhances solution quality.
+Actions are decoded efficiently autoregressively in parallel through a Multiple Pointer Mechanism enhanced by a Priority-based Conflict Handler.
+4.1 Parallel Multi-Agent MDPs
+We reformulate the multi-agent MDPs of Section 3.1 as parallel multi-agent MDPs by selecting multiple actions from a joint action space simultaneously to enhance efficiency and coordination (Yu et al., 2022).
+At each step $t$, ${M}$ agents select actions $\bm{a}_{t}=(a_{t}^{1},...,a_{t}^{M})$ according to a policy $\pi_{\theta}:\mathcal{S}\rightarrow\mathcal{A}_{1}\times...\times\mathcal{A}_{M}$, which maps the state space $\mathcal{S}$ to the joint action space of agents.
+A conflict handling function $\psi:(\mathcal{A}_{1}\times...\times\mathcal{A}_{M})\rightarrow(\mathcal{A}_{1}\times...\times\mathcal{A}_{M})$ ensures action compatibility by allowing only one agent to execute when multiple agents select mutually conflicting actions (e.g., the same customer location), and assigning a fallback action (e.g., staying at current position) to the others.
+Given the resolved agent actions $\tilde{\bm{a}}_{t}=\psi(\bm{a}_{t})$, the state of the problem $s_{t}$ progresses to $s_{t+1}$ according to the transition function $\tau:\mathcal{S}\times(\mathcal{A}_{1}\times...\times\mathcal{A}_{M})\rightarrow\mathcal{S}$.
+The agents receive a shared reward $R(\bm{a},\bm{x})$, with $\bm{a}=(\tilde{\bm{a}}_{1},...,\tilde{\bm{a}}_{T})$ the sequence of joint agent actions.
+In the following, we refer to the entities corresponding to actions in the MDP formulation of a CO problem (e.g., customer locations in VRPs) as nodes, following the convention of Kool et al.
+(2019a) and Kwon et al.
 (2020).
-The core of each transformer block is a multi-head attention (MHA) layer, which enables message passing between the nodes in the graph and can be defined as follows (Vaswani et al., 2017):
-$\text{MHA}(Q,K,V)$ $=\left(\mathop{\big{\|}}_{i=1}^{h}\text{Attn}(QW^{Q}_{i},KW^{K}_{i}VW^{V}_{i})\right)W^{O}$
-where
-$\text{Attn}(Q,K,V)$ $=\text{softmax}\left(\frac{QK^{\top}}{\sqrt{d_{k}}}\right)V\,.$
+4.2 Parallel AR Models for CO
+Motivated by the nature of multi-agent MDPs, PARCO introduces a Parallel AR model for the policy $\pi_{\theta}$.
+PARCO constructs feasible solutions by simultaneously generating multiple agent actions based on the current state.
+We formulate the solution generation process in an encoder-decoder framework similarly to Equation 1:
+$p_{\theta}(\bm{a}|\bm{x})$ $\triangleq\prod_{t=1}^{T}\psi\left(\prod_{m=1}^{M}g_{\theta}(a_{t}^{m}|\bm{a}_{<t},\bm{h})\right)$ (3)
+where $\bm{h}=f_{\theta}(\bm{x})$ is the encoding of problem instance $\bm{x}$ via encoder network $f_{\theta}$, and decoder $g_{\theta}$ implements the policy $\pi_{\theta}$ to autoregressively generate actions efficiently for all agents in parallel.
+At each step $t$, the decoder outputs joint actions $\bm{a}_{t}=(a_{t}^{1},...,a_{t}^{M})$ for all $M$ agents with $a_{t}^{m}$ being the sampled action of agent $m$.
+The conflict resolution function $\psi$ ensures action compatibility by allowing only one agent to execute when multiple agents select mutually conflicting actions, resulting in resolved actions $\tilde{\bm{a}}_{t}=\psi(\bm{a}_{t})$.
+$p_{\theta}$ is thus a solver that maps $\bm{x}$ to a solution $\bm{a}=(\tilde{\bm{a}}_{1},...,\tilde{\bm{a}}_{T})$.
+A benefit of our parallel formulation is that the total number of construction steps $T$ can be substantially lower compared to purely AR methods.
+While the latter require $\sum_{m=1}^{M}T_{m}$ total actions to construct a solution with $T_{m}$ being the number of steps required by agent $m$ to finish its task, PARCO’s Parallel AR needs only $\max_{m}T_{m}$ steps as agents effectively divide the solution space and act concurrently.
+This leads to a faster solution construction as illustrated in Figures 1 and 5 and significantly reduced training times.
+4.3 Multi-Agent Encoder
+The multi-agent encoder $f_{\theta}$ transforms an input instance $\bm{x}$ into a hidden representation $\bm{h}$.
+In PARCO, we explicitly model agents and employ separate agent and node embedding layers similar to Son et al.
+(2024) to project agents and nodes into the same embedding space.
+The agent embedding layer projects $k_{a}$ agent features – such as vehicle locations and capacities (routing) or machine characteristics (scheduling) – into a $d$-dimensional space using a linear projection ${\mathbf{W}}_{a}\in\mathbb{R}^{k_{a}\times d}$.
+Let $\bm{h}_{a}^{(0)}=\bm{x}_{a}{\mathbf{W}}_{a}$ denote the initial agent embeddings, where $\bm{x}_{a}\in\mathbb{R}^{M\times k_{a}}$ denotes the matrix of agent features.
+Similarly, the node embedding layer projects $k_{n}$ node features – such as customer demands in vehicle routing or job durations in scheduling problems – into the same $d$-dimensional space using a linear projection ${\mathbf{W}}_{n}\in\mathbb{R}^{k_{n}\times d}$.
+The initial embeddings of the nodes are defined as $\bm{h}_{n}^{(0)}=\bm{x}_{n}{\mathbf{W}}_{n}$ where $\bm{x}_{n_{i}}\in\mathbb{R}^{N\times k_{n}}$ represents the node feature matrix.
+Depending on the problem structure, the initial agent and node embeddings might either be concatenated as $\bm{h}^{(0)}=\text{Concat}(\bm{h}_{a}^{(0)},\bm{h}_{n}^{(0)})$ and passed through $L$ transformer blocks, consisting of multi-head self-attention $\text{MHA}(\bm{h}^{(0)},\bm{h}^{(0)},\bm{h}^{(0)})$ and multi-layer perceptrons (MLPs) as defined in Vaswani et al.
+(2017).
+Or, agent and node embeddings are used separately as query and keys/values, respectively, in a cross-attention mechanism $\text{MHA}(\bm{h}_{a}^{(0)},\bm{h}_{n}^{(0)},\bm{h}_{n}^{(0)})$ akin to MatNet (Kwon et al., 2021).
+The final embeddings $\bm{h}=\{\bm{h}_{a},\bm{h}_{n}\}$ emitted by the last encoder layer contain processed representations of both agents and nodes that capture their interactions as well as the overall problem structure.
+4.4 Communication Layers
+At each step $t$ of the decoding process of Equation 3, given the encoded representations $\bm{h}$, we construct dynamic agent queries that capture the current state of both agents and the environment.
+For each agent $m$, we form a context embedding $\bm{d}_{m}=\text{Concat}(\bm{h}_{a^{m}},\bm{h}_{\delta^{m}_{t}},\bm{h}_{e})$ consisting of the following components: (1) the (static) embedding of the agent $\bm{h}_{a^{m}}$; (2) a projection $\bm{h}_{\delta^{m}_{t}}=\delta^{m}_{t}{\mathbf{W}}_{\delta}\in\mathbb{R}^{d}$ of the agent’s dynamics $\delta^{m}_{t}$ like its current location and capacity (routing) or the time until the agent becomes idle (scheduling); (3) a projection $\bm{h}_{e}=e_{t}{\mathbf{W}}_{\delta}\in\mathbb{R}^{d}$ of the dynamic environment features $e_{t}$ that encode the current problem’s state.
+These dynamic embeddings are then projected into query vectors $\bm{q}_{m}=\bm{d}_{m}{\mathbf{W}}_{q}$ where ${\mathbf{W}}_{q}\in\mathbb{R}^{3d\times d}$ is a learnable projection matrix.
+The resulting queries $\bm{q}=[\bm{q}_{1},\dots,\bm{q}_{M}]\in\mathbb{R}^{M\times d}$ are then processed through communication layers comprising multi-head self-attention followed by an MLP:
+$\bm{q}^{\prime}$ $=\text{Norm}(\text{MHA}(\bm{q},\bm{q},\bm{q})+\bm{q})$ (4)
+$\bm{q}$ $=\text{Norm}(\text{MLP}((\bm{q}^{\prime}))+\bm{q}^{\prime})$ (5)
+where Norm denotes a normalization layer (Ioffe & Szegedy, 2015; Zhang & Sennrich, 2019).
+These layers are inherently agent-count agnostic, which allows PARCO to handle an arbitrary number of agents, making it more flexible and generalizable across different problems.
+Communication Layers allow agents to coordinate their actions by attending to both the problem structure and other agents’ states while maintaining efficiency through parallel processing.
+4.5 Decoder with Multiple Pointer Mechanism
+PARCO’s decoder extends the attention-based pointer mechanism of Kool et al.
+(2019a) – originally designed for single-agent scenarios and recently applied to sequential multi-agent planning with a single agent at a time (Son et al., 2024; Zheng et al., 2024a) – to handle multiple agents operating in parallel via a Multiple Pointer Mechanism.
+Starting with the processed agent queries $\bm{q}$ that underwent communication, we first compute agent-specific representations through masked cross MHA:
 […]
-PARCO additionally defines an agent projection layer that maps agents into the same embedding space as the problem nodes, whose exact definition heavily depends on the problem.
-In routing problems, for example, agent embeddings can be represented by the depot location they are starting from and their capacities, while an agent in scheduling problems can be represented by the machine whose schedule it is constructing.
-Lastly, encoded actions and agents can further be enhanced by performing cross-attention between them, depending on the structure of the problem (Kwon et al., 2021).
-We describe the specific encoder details for the different problems covered in this work in the Appendix.
-Having a separate agent encoder contrasts other multi-agent (NCO) approaches like Zhang et al.
-(2020b), which use separate policies $\pi^{k}_{\theta_{m}}$ for each agent.
-This has a substantial drawback, as a model trained for a given number of agents cannot be used to solve problems with a different number of agents.
-The PARCO encoder, on the other hand, can generate representations for arbitrary agents, making it agnostic to the problem size.
-Decoder
-The decoder is the core of the PARCO architecture. We adopt the pointer mechanism from Kool et al. (2018), which yields unnormalized log-probabilities $\bm{z}\in\mathbb{R}^{N}$ for each action/node:
-$\bm{z}=C\cdot tanh\left(\frac{\bm{u}^{\top}L}{\sqrt{d_{h}}}\right)$ (3)
-where $C$ is a scale parameter, $L$ is a projection of the node embeddings $\bm{h}$ and $\bm{u}$ is the output of an MHA layer, in which the query is defined by a single context embedding $\bm{q}^{c}_{t}\in\mathbb{R}^{d_{h}}$, specifying the current state of the problem at decoding step $t$ and keys and values are the node embeddings from the encoder.
-Each attention head of the MHA layer can further fuse the projection of the encoder embeddings with dynamic embeddings representing current information about the particular node, i.e.
-$\text{Attn}(\bm{q}^{c}_{t}W^{Q}_{i},\,\bm{h}W^{K}_{i}+W^{K}_{\delta}\delta_{t},\,\bm{h}W^{V}_{i}+W^{V}_{\delta}\delta_{t})$ (4)
-with $\delta_{t}$ encapsulating the dynamic elements of the problem nodes and $W^{K}_{\delta},W^{V}_{\delta}\in\mathbb{R}^{|\delta|\times d_{m}}$ projecting them into embedding space.
-For PARCO we introduce a novel Multiple Pointer Mechanism based on Eq. 3 of the following form:
-$Z=C\cdot tanh\left(\frac{UL^{\top}}{\sqrt{d_{h}}}\right)$ (5)
-Here, $U\in\mathbb{R}^{m\times d_{h}}$ is the output of an equivalent MHA layer than the one used in Kool et al.
-(2018).
-However, instead of a single context embedding, PARCO uses the $m$ agent embeddings of the encoder, enriched by the current state’s context and individual agent’s state embeddings, as queries $Q^{c}_{t}\in\mathbb{R}^{m\times d_{h}}$.
-The resulting $Z\in\mathbb{R}^{m\times N}$ log-probabilities are used to sample $m$ actions - one for each agent - in parallel.
-Communication Layers
-To enable effective agent coordination, we introduce a communication layer before the Multiple Pointer Mechanism, which updates the agent embeddings through messages exchanged with other agents.
-This component is vital for parallel decoding of multiple actions, as the quality of an action chosen by one agent is significantly influenced by how it affects the actions of other agents.
-Given the $m$ agent queries at decoding step $t$, the communication layer applies a transformer block (Vaswani et al., 2017) to capture intra-agent relationships and spot potential conflicts between agents:
-$H^{\prime}$ $=\text{Norm}(\text{MHA}(Q^{c}_{t},Q^{c}_{t},Q^{c}_{t})+Q^{c}_{t})$ (6)
-$Q^{\prime}_{t}$ $=\text{Norm}(\text{MLP}(H^{\prime})+H^{\prime})$ (7)
-where Norm denotes a normalization layer (Ioffe and Szegedy, 2015; Zhang and Sennrich, 2019) and MLP represents the multi-layer perceptron.
-Self-attention in the MHA layer enables message passing between agents based on their embeddings.
-By applying the communication layer after fusing the agent embeddings with the dynamic context embeddings, it can capture dynamic relationships that could not be resolved during encoding, thereby enabling the decoder to make more informed decisions.
-The updated query $Q^{\prime}_{s}$ is finally passed into the Multiple Pointer Mechanism.
-4.2.1 Conflicts Handlers
-When sampling from the probability distribution generated by the Multiple Pointer Mechanism, it is possible for multiple agents to select the same node simultaneously, resulting in a conflict.
-These conflicts must be resolved to ensure the generation of a feasible solution.
-To address this, we introduce a Priority-based Conflict Handler $\mathcal{H}$, which resolves such conflicts by leveraging a predefined priority scheme.
-The conflict handler can be defined as a function $\mathcal{H}:\mathbb{N}^{m}\times\mathbb{R}^{m}\rightarrow\mathbb{N}^{m}$ with number of agents $m$.
-The priorities are determined based on the current actions and states.
-Given an input vector of actions $\bm{a}=(a_{1},a_{2},\dots,a_{m})$ where $a_{i}\in\mathbb{N}$, the corresponding priority vector $\bm{p}=(p_{1},p_{2},\dots,p_{m})$ where $p_{i}\in\mathbb{R}$ and the fallback actions (i.e., the previous node) $\bm{r}=(r_{1},r_{2},\dots,r_{m})$ where $r_{i}\in\mathbb{N}$, the output is another vector of actions $\bm{a}^{\prime}=(a^{\prime}_{1},a^{\prime}_{2},\dots,a^{\prime}_{m})$ such that any conflicts are resolved based on the priority order as illustrated in Algorithm 1.
+where $\xi_{t}\in\mathbb{R}^{N\times k_{\xi}}$ are dynamic node features which are projected via ${\mathbf{W}}^{K}_{\xi},{\mathbf{W}}^{V}_{\xi}\in\mathbb{R}^{k_{\xi}\times d}$ for keys and values of the MHA, respectively.
+Further, $\mathbf{M}_{t}\in\mathbb{R}^{M\times N}$ is the current action mask at step $t$, avoiding agent representations to attend to infeasible actions.
+We then obtain a joint logit space $\bm{u}$ across all agents:
+$\bm{u}={\beta}\cdot\text{tanh}\left(\frac{\bm{q}(\bm{h}_{n}{\mathbf{W}}^{L}+\xi_{t}{\mathbf{W}}^{L}_{\xi})^{\top}}{\sqrt{d}}\right)$ (7)
+with learnable parameters ${\mathbf{W}}^{L}\in\mathbb{R}^{d\times d}$, ${\mathbf{W}}^{L}_{\xi}\in\mathbb{R}^{k_{\xi}\times d}$ and ${\beta}$ is a scale parameter, set to $10$ following Bello et al.
+(2016) to enhance exploration.
+The output logits $\bm{u}\in\mathbb{R}^{M\times N}$ are masked by setting infeasible actions given mask $\textbf{M}_{t}$ to minus infinity.The joint probability distribution over all agent actions is then computed as:
+$p(\bm{a}_{t}|\bm{a}_{<t},\bm{h})=\prod_{m=1}^{M}\frac{\exp(\bm{u}_{m,a_{t}^{m}})}{\sum_{j=1}^{N}\exp(\bm{u}_{m,j})}$ (8)
+where $\bm{a}_{t}=(a_{t}^{1},\ldots,a_{t}^{M})$ represents the joint action across all agents at step $t$.
+4.6 Conflict Handlers
+When sampling from the probability distribution $\bm{p}$ generated by the Multiple Pointer Mechanism, multiple agents may select the same action simultaneously, which can result in an infeasible solution in several CO problems: for instance, in vehicle routing problems, usually only one agent is allowed to visit a customer node.
+Thus, it is essential to choose how to deal with such a situation effectively.
+Conflict handling (i.e., tie-breaking) can be achieved by allowing a single agent among a number of agents that are in conflict to continue with its action while others revert to fallback actions, e.g., staying in their current position.
+A naive approach consists of randomly selecting an agent to perform the new action (Zong et al., 2022).
+However, this can be suboptimal since it excludes inductive biases that can be leveraged, such as learned representations.
 Algorithm 1 Priority-based Conflict Handler
-0:  Actions $\bm{a}\in\mathbb{N}^{m}$, Priorities $\bm{p}\in\mathbb{R}^{m}$, Fallback actions $\bm{r}\in\mathbb{N}^{m}$
-0:  Resolved Actions $\bm{a^{\prime}}\in\mathbb{N}^{m}$
-1:  $\sigma\leftarrow\text{argsort}(\bm{p},\text{descending}=\text{True})$Sort indices by priority}
-2:  $\hat{\bm{a}}\leftarrow\bm{a}[\sigma]$Reorder actions according to priority}
-3:  Initialize conflict mask $M\leftarrow\bm{0}^{m}$
+0:  Actions $\bm{a}\in\mathbb{N}^{M}$, Priorities $\bm{p}\in\mathbb{R}^{M}$, Fallback actions $\bm{r}\in\mathbb{R}^{M}$
+0:  Resolved Actions $\bm{a}^{\prime}\in\mathbb{N}^{M}$ // Sort indices based on priorities in descending order.
+1:  $\sigma\leftarrow\text{argsort}(\bm{p},\text{descending}=\text{True})$ // Reorder actions according to priority.
+2:  $\hat{\bm{a}}\leftarrow\bm{a}[\sigma]$ // Initialize conflict mask (all zeros).
+3:  $C\leftarrow\bm{0}^{M}$ // Check for conflicts in reordered actions.
 4:  for $i=2$ to ${M}$ do
 5:     if $\hat{\bm{a}}_{i}\in\{\hat{\bm{a}}_{1},\ldots,\hat{\bm{a}}_{i-1}\}$ then
-6:        $M_{i}\leftarrow 1$Identify conflicts}
+6:        $C_{i}\leftarrow 1$
 7:     end if
-8:  end for
-9:  $\hat{\bm{a}}\leftarrow(1-M)\odot\hat{\bm{a}}+M\cdot r$Resolve conflicts by assigning fallback action $r$}
-10:  $\bm{a}^{\prime}\leftarrow\hat{\bm{a}}[\sigma^{-1}]$Reorder actions back to original sequence}
-11:  return  $\bm{a^{\prime}}$
-4.2.2 Step Definition
-We define a single parallel step of the model similarly to Eq. 1 as follows:
-$p_{\theta}(\bm{a}|\bm{x})$ $=\prod_{i=1}^{m}g_{\theta}(a_{t,i}^{\prime}|a_{t-1,i}^{\prime},\dots,a_{0,i}^{\prime},\bm{h})$ (8)
-Here, $a_{t,i}^{\prime}$ denotes the action executed by agent $i$ at time $t$, after passing through the conflict handler $\mathcal{H}$:
-$a_{t,i}^{\prime}=\mathcal{H}(a_{t,1},a_{t,2},\dots,a_{t,m};\bm{p}_{t},\bm{r}_{t}),$
-where $\bm{p}_{t}$ and $\bm{r}_{t}$ are priorities and fallback actions, respectively, at the current decoding step $t$.
-The conflict handler $\mathcal{H}$ ensures that multiple agents do not select the same node, thus maintaining the integrity and feasibility of the solution.
-4.3 Training
-PARCO is a centralized multi-agent sequential decision-making model, with a shared policy $\pi_{\theta}$ for all agents and a global reward $R$.
-Thus, PARCO can be trained using any of the training algorithms adopted in the single-agent NCO literature.
-We train PARCO via the REINFORCE gradient estimator (Williams, 1992) with a shared baseline as outlined by Kwon et al.
+8:  end for// Resolve conflicts by assigning fallback actions.
+9:  $\hat{\bm{a}}\leftarrow(1-C)\odot\hat{\bm{a}}+C\odot\bm{r}$ // Reorder resolved actions back to original order.
+10:  $\bm{a}^{\prime}\leftarrow\hat{\bm{a}}[\sigma^{-1}]$
+In PARCO, we propose Priority-based Conflict Handlers that leverage priorities as a tie-breaking rule.
+Such priorities can be based on heuristics – such as giving priority to agents close to completion or whose action results in the smallest immediate cost – or on learned priorities.
+In the latter case, the probability values of the selected actions $p(\bm{a}_{t})$ serve as an indicator for prioritizing certain agents: the higher its value, the higher the priority learned by the model to have that agent win the tie-break.
+Algorithm 1 shows an efficient vectorized implementation of the Priority-based Conflict Handler algorithm.
+In practice, we augment the conflict handler $\psi$ from Section 4.2 with priorities $\bm{p}\coloneq p(\bm{a}_{t})$ and fallback actions $\bm{r}$, i.e.
+$\psi(\cdot)\coloneq\psi(\bm{a}_{t},\bm{p},\bm{r})$.
+Fallback actions in PARCO correspond to “do nothing” operations – maintaining an agent’s current position (routing) or keeping a machine idle (scheduling) – which may result in slightly more solution construction steps but do not affect the final solution $\bm{a}$.
+This approach effectively handles conflicts by allowing the affected agents to reconsider their choices given the actions of (preceding) agents in the next decoding step.
+4.7 Training Scheme
+PARCO is a centralized multi-agent decision-making framework and can thus be trained by RL algorithms proposed in the single-agent NCO literature.
+We train PARCO by employing the REINFORCE gradient estimator (Williams, 1992) with a shared baseline as outlined by Kwon et al.
 (2020) and Kim et al.
 (2022):
-$\nabla_{\theta}\mathcal{L}\approx\frac{1}{B\cdot L}\sum_{i=1}^{B}\sum_{j=1}^{L}G_{ij}\nabla_{\theta}\,\text{log}\,p_{\theta}(A_{ij}|\bm{x}_{i})$ (9)
-where $B$ is the size of the mini-batch and $G_{ij}$ is the advantage $R(A_{ij},\bm{x}_{i})-b^{\text{shared}}(\bm{x}_{i})$ of a solution $A_{ij}$ compared to the shared baseline $b^{\text{shared}}_{i}$ of problem instance $\bm{x}_{i}$.
+$\nabla_{\theta}\mathcal{L}\approx\frac{1}{B\cdot S}\sum_{i=1}^{B}\sum_{j=1}^{S}G_{ij}\nabla_{\theta}\,\text{log}\,p_{\theta}(\bm{a}_{ij}|\bm{x}_{i})$ (9)
+where $B$ is the batch size, $S$ the number of shared baseline samples, and $G_{ij}=R(\bm{a}_{ij},\bm{x}_{i})-b^{\text{shared}}(\bm{x}_{i})$ is the advantage of a solution $\bm{a}_{ij}$ compared to the shared baseline $b^{\text{shared}}_{i}$ of problem instance $\bm{x}_{i}$.
 5 Experiments
-In this section, we present the experimental results of PARCO in two routing problems, the min-max heterogenous capacitated vehicle routing problem (HCVRP) and the open multi-depot capacitated pickup and delivery problem (OMDCPDP), and a scheduling problem, namely the flexible flow shop problem (FFSP).
-We provide more details about the problem and experimental setups in Appendix A and Appendix B respectively.
-5.1 Problem Descriptions
-HCVRP
-The min-max HCVRP consists of $m$ agents sequentially visiting customers to satisfy their demands, with constraints including each customer can be visited exactly once and the amount of demand satisfied by a single vehicle in a trip cannot exceed its capacity, which can be reloaded by going back to the depot.
-The goal is to minimize the makespan, i.e., the worst route.
-Baselines include SISR (Christiaens and Vanden Berghe, 2020), Genetic Algorithm (GA) (Karakatič and Podgorelec, 2015), Simulated Annealing (SA) (İlhan, 2021), the Attention Model (AM) (Kool et al., 2018), Equity Transformer (ET) (Son et al., 2024), the model from Li et al.
-(2022) (DRLLi), and the state-of-the-art neural baseline 2D-Ptr (Liu et al., 2024c).
-OMDCPDP
-The OMDCPDP problem is a practical variant of the pickup and delivery problem in which agents have a stacking limit of orders that can be carried at any given time.
-Pickup and delivery locations are paired, and pickups must be visited before deliveries.
-Multiple agents start from different depots with no need to go back (open).
-The goal is to minimize the sum of arrival times to delivery locations, i.e.
-minimizing the cumulative lateness.
-We include ORTools (Furnon and Perron, 2024) as a classical baseline, the Heterogeneous Attention Model (HAM) (Li et al., 2021a) for sequential decision-making and MAPDP (Zong et al., 2022) for parallel decision-making.
-FFSP
-In the flexible flow shop problem (FFSP), $N$ jobs must be processed across $S$ stages, each with multiple machines ($m>1$).
-Jobs follow a specified sequence through these stages, but within each stage, any available machine can process the job, with the key constraint that no machine can handle more than one job simultaneously.
-The goal is to schedule the jobs so that all jobs are finished in the shortest time possible.
-Notable benchmarks include the MatNet model (Kwon et al., 2021), the Random and Shortest Job First (SJF) dispatching rules, as well as the evolutionary algorithms Particle Swarm Optimization (PSO), and Genetic Algorithm (GA) (Hejazi and Saghafian, 2005).
-Table 1: Benchmarks and results of our model for min-max HCVRP problems of varying sizes and agent numbers.
-Highlighting cost ($\downarrow$) and average gaps ($\downarrow$) from the best-known solutions of classical heuristic solvers.
-Inference times are shown in seconds in parentheses ()($\cdot$)( ⋅ ).
-(g.) refers to greedy performance while (s.) refers to sampling 1280 solutions.
-[…]
-$m$ 3 5 7 3 5 7 3 5 7 avg.
-SISR 6.57 (271) 4.00 (274) 2.91 (276) 8.52 (425) 5.10(430) 3.69 (434) 10.29 (615) 6.17 (623) 4.45 (625) 0.00
-GA 9.21 (233) 6.89 (320) 5.98 (405) 12.32 (347) 8.95 (465) 7.58 (578) 15.33 (479) 10.93 (623) 9.10 (772) 74.90
-SA 7.04 (130) 4.39 (289) 3.30 (362) 9.17 (318) 5.61 (417) 4.17 (515) 11.13 (434) 6.80 (557) 5.01 (678) 10.21
-AM (g.) 8.49 (0.08) 5.51 (0.08) 4.15 (0.09) 10.81 (0.10) 6.87 (0.11) 5.18 (0.10) 12.68 (0.14) 8.10 (0.13) 6.13 (0.13) 33.80
-ET (g.) 7.58 (0.15) 4.76 (0.17) 3.58 (0.16) 9.76 (0.21) 6.01 (0.20) 4.43 (0.23) 11.74 (0.25) 7.25 (0.25) 5.23 (0.26) 17.67
-DRLLi (g.) 7.43 (0.19) 4.71 (0.22) 3.60 (0.25) 9.64 (0.25) 5.97 (0.30) 4.52 (0.33) 11.44 (0.32) 7.06 (0.37) 5.38 (0.43) 17.08
-2D-Ptr (g.) 7.20 (0.11) 4.48 (0.11) 3.31 (0.11) 9.24 (0.15) 5.65 (0.15) 4.14 (0.14) 11.12 (0.18) 6.75 (0.18) 4.92 (0.17) 10.54
-PARCO (g.) 7.12 (0.12) 4.40 (0.11) 3.25 (0.11) 9.14 (0.15) 5.53 (0.15) 4.04 (0.15) 10.98 (0.19) 6.61 (0.18) 4.79 (0.18) 8.56
-AM (s.) 7.62 (0.14) 4.82 (0.13) 3.63 (0.14) 9.92 (0.20) 6.19 (0.21) 4.64 (0.22) 11.82 (0.29) 7.45 (0.28) 5.58 (0.28) 20.69
-ET (s.) 7.14 (0.21) 4.46 (0.22) 3.33 (0.22) 9.19 (0.30) 5.64 (0.30) 4.17 (0.31) 11.20 (0.41) 6.85 (0.38) 4.98 (0.40) 10.89
-DRLLi (s.) 6.97 (0.30) 4.34 (0.36) 3.25 (0.43) 9.10 (0.45) 5.54 (0.55) 4.13 (0.65) 10.90 (0.60) 6.65 (0.76) 4.98 (0.92) 8.78
-2D-Ptr (s.) 6.82 (0.13) 4.20 (0.13) 3.09 (0.14) 8.85 (0.17) 5.36 (0.18) 3.90 (0.19) 10.71 (0.22) 6.46 (0.23) 4.68 (0.24) 4.84
-PARCO (s.) 6.82 (0.31) 4.17 (0.31) 3.06 (0.31) 8.79 (0.36) 5.28 (0.36) 3.83 (0.35) 10.61 (0.40) 6.36 (0.40) 4.58 (0.40) 3.65
-5.2 Experimental Setup
-We perform all experiments on a machine equipped with two Intel(R) Xeon(R) Gold 6338 CPU @ 2.00GHz CPUs with a total $128$ threads and $8$ NVIDIA RTX 4090 graphic cards with $24$ GB of VRAM.
+We assess the effectiveness of PARCO on representative multi-agent CO problems, spanning both routing and scheduling domains.
+Specifically, we evaluate its performance on two challenging routing problems – the min-max heterogeneous capacitated vehicle routing problem (HCVRP) and the open multi-depot capacitated pickup and delivery problem (OMDCPDP) – as well as a scheduling problem, the flexible flow shop problem (FFSP).
+5.1 Experimental Settings
+We experiment on a workstation equipped with 2 Intel(R) Xeon(R) Gold 6338 CPUs and $8$ NVIDIA RTX 4090 graphic cards with $24$ GB of VRAM each.
 Training runs of PARCO take less than 24 hours each.
 During inference, we employ only one CPU and a single GPU.
-We report key metrics such as solution cost, inference times, and gaps in best-known solutions.
-We keep most settings of PARCO consistent across experiments, i.e., we use a Communication Layer and Priority-based Conflict Handling based on the highest probability action from the model output.
-We provide additional details regarding training and testing setups in the Appendix.
-(a) Effect of Communication Layers.
-(b) Effect of Conflict Handlers.
-(c) PARCO vs AR decoding speed.
-Figure 3: Ablation studies on PARCO components. PARCO scales better than AR models as ET.
-Table 2: Benchmarks and results of our model for OMDCPDP problems of varying sizes and agent numbers.
-[…]
-$N$ 50 100 200 1000 avg.
-$m$ 10 18 25 20 35 50 40 70 100 100 250 500
-OR-Tools 23.73 18.64 16.92 43.34 34.80 31.89 78.35 64.93 60.16 562.11 333.01 310.55 0.00%
-HAM (g.) 34.85 24.79 18.52 65.95 51.56 35.41 124.16 98.32 67.45 740.34 475.76 326.07 33.52%
-MAPDP (g.) 27.95 20.13 17.54 52.44 38.23 $3$ 10.80%
-PARCO (g.) 24.70 19.19 17.44 44.88 35.51 32.79 81.14 66.95 61.95 531.95 344.67 302.09 1.96%
-HAM (s.) 32.84 25.45 18.12 63.59 49.12 34.90 122.77 96.32 67.01 739.77 471.09 321.50 31.02%
-MAPDP (s.) 25.49 20.02 17.17 49.55 37.37 33.25 - - - - - - 7.04%
-PARCO (s.) 24.27 18.71 17.00 43.36 34.82 32.19 79.23 65.83 61.12 526.52 338.78 298.71 -0.01%
-Table 3: FFSP results. PARCO improves the MatNet performance and is also more than 4×4$\times$4 × faster overall.
-Method FFSP20 FFSP50 FFSP100
-MS Gap Time (training) MS Gap Time (training) MS Gap Time (training)
-CPLEX 46.4 21.0 60s - -
-CPLEX 36.6 6.4 600s - -
-Random 47.8 22.9 0.1s 93.2 44 0.2s 167.2 78.0 0.3s
-Shortest Job First 31.3 6.4 0.1s 57.0 7.8 0.1s 99.3 10.1 0.2s
-Genetic Algorithm 30.6 5.7 25.4s 56.4 7.2 57.8s 98.7 9.5 104.5s
-Particle Swarm Opt. 29.1 4.2 46.8s 55.1 5.9 92.9s 97.3 8.1 173.4s
-MatNet 27.26 3.0 0.9s (2h) 51.52 0.5 2.1s (5.25h) 91.58 0.0 4.9s (28h)
-PARCO (w/o CL) 27.02 2.1 0.2s (0.25h) 52.29 2.0 0.4s (1.25h) 92.15 2.0 0.9s (5.25h)
-PARCO 26.47 0.0 0.2s (0.5h) 51.25 0.0 0.5s (2h) 91.55 0.0 1.1s (6.25h)
+Additional details are available throughout Appendix B.
+HCVRP
+Problem.
+The min-max HCVRP involves ${M}$ agents serving customer demands while adhering to heterogeneous vehicle capacity constraints.
+Each vehicle can replenish its load by returning to the depot.
+The objective is to minimize the longest route taken by any agent (min-max), ensuring balanced workload distribution.
+Traditional solvers.
+We include state-of-the-art SISRs (Christiaens & Vanden Berghe, 2020), Genetic Algorithm (GA) (Karakatič & Podgorelec, 2015) and Simulated Annealing (SA) (İlhan, 2021).
+Neural baselines.
+We evaluate the sequential planning baselines Attention Model (AM) (Kool et al., 2019a), Equity Transformer (ET) (Son et al., 2024) and Decoupling Partition and Navigation (DPN) (Zheng et al., 2024a), and autoregressive models with agent selection DRLLi (Li et al., 2022) and state-of-the-art learning method 2D-Ptr (Liu et al., 2024b).
+Additional problem and experimental details are available in Section A.1 and Section B.1, respectively.
+OMDCPDP
+Problem.
+The OMDCPDP is a challenging problem arising in last-mile delivery settings where $M$ agents starting from different locations (i.e., multiple depots) must pick up and deliver parcels without returning to their starting point (i.e., open).
+Agents have a capacity constraint for orders that can be carried out as a stacking limit: a tour can include more pickups than the constraint, but the agent must deliver corresponding orders so that carrying capacity is freed.
+The goal is to minimize the lateness, i.e., the sum of delivery arrival times.
+Traditional solvers.
+We include the popular and efficient optimization suite Google OR-Tools (Furnon & Perron, 2024) as a classical baseline.
+Neural baselines.
+We evaluate models specializing in pickup and delivery problems, including the autoregressive Heterogeneous Attention Model (HAM) (Li et al., 2021a) for sequential planning and MAPDP (Zong et al., 2022) for parallel planning.
+Additional problem and experimental details are available in Section A.2 and Section B.2, respectively.
+FFSP
+Problem.
+In FFSP, $N$ jobs must be processed by $M$ machines divided equally in $S$ stages.
+Jobs follow a specified sequence through these stages.
+Within each, any available machine can process the job, with the key constraint that no machine can handle multiple jobs simultaneously.
+The goal is to schedule the jobs so that all jobs are finished in the shortest time possible.
+Traditional solvers.
+We incorporate the widely used and powerful Gurobi solver (Gurobi Optimization, LLC, 2024) as a baseline.
+Furthermore, we include dispatching rules Random and Shortest Job First (SJF), Particle Swarm Optimization (PSO) (Singh & Mahapatra, 2012) and Genetic Algorithm (GA) (Hejazi & Saghafian, 2005).
+Neural baselines.
+Notable benchmarks include the Matrix Encoding Network (MatNet) (Kwon et al., 2021) which demonstrates superior performance on FFSP.
+Additional problem and experimental details are available in Section A.3 and Section B.3, respectively.
+Table 1: Main results with different configurations in terms of problem size $N$ and number of agents $M$. The lower the better ($\downarrow$).
+HCVRP
+$N$ 60 100
+$M$ 3 5 7 3 5 7
+Metric Obj. Gap Time Obj. Gap Time Obj. Gap Time Obj. Gap Time Obj. Gap Time Obj. Gap Time
+SISRs 6.57 0.00% 271s 4.00 0.00% 274s 2.91 0.00% 276s 10.29 0.00% 615s 6.17 0.00% 623s 4.45 0.00% 625s
+GA 9.21 40.18% 233s 6.89 72.25% 320s 5.98 105.50% 405s 15.33 48.98% 479s 10.93 77.15% 623s 9.10 104.49% 772s
+SA 7.04 7.15% 130s 4.39 9.75% 289s 3.30 13.40% 362s 11.13 8.16% 434s 6.80 10.21% 557s 5.01 12.58% 678s
+AM (g.) 8.49 29.22% 0.08s 5.51 37.75% 0.08s 4.15 42.61% 0.09s 12.68 23.23% 0.14s 8.10 31.28% 0.13s 6.13 37.75% 0.13s
+ET (g.) 7.58 15.37% 0.15s 4.76 19.00% 0.17s 3.58 23.02% 0.16s 11.74 14.09% 0.25s 7.25 17.50% 0.25s 5.23 17.53% 0.26s
+DPN (g.) 7.50 14.16% 0.18s 4.60 15.00% 0.19s 3.45 18.56% 0.26s 11.54 12.15% 0.30s 6.94 12.48% 0.40s 4.98 11.91% 0.43s
+DRLLi (g.) 7.43 13.09% 0.19s 4.71 17.75% 0.22s 3.60 23.71% 0.25s 11.44 11.18% 0.32s 7.06 14.42% 0.37s 5.38 20.90% 0.43s
+2D-Ptr (g.) 7.20 9.59% 0.11s 4.48 12.00% 0.11s 3.31 13.75% 0.11s 11.12 8.07% 0.18s 6.75 9.40% 0.18s 4.92 10.56% 0.17s
+PARCO (g.) 7.12 8.37% 0.04s 4.40 10.00% 0.05s 3.25 11.68% 0.05s 10.98 6.71% 0.06s 6.61 7.13% 0.05s 4.79 7.64% 0.05s
+AM (s.) 7.62 15.98% 0.14s 4.82 20.50% 0.13s 3.63 24.74% 0.14s 11.82 14.87% 0.29s 7.45 20.75% 0.28s 5.58 25.39% 0.28s
+ET (s.) 7.14 8.68% 0.21s 4.46 11.50% 0.22s 3.33 14.43% 0.22s 11.20 8.84% 0.41s 6.85 11.02% 0.38s 4.98 11.91% 0.40s
+DPN (s.) 7.08 7.76% 0.25s 4.35 8.75% 0.28s 3.20 9.97% 0.38s 11.04 7.29% 0.48s 6.66 7.94% 0.52s 4.79 7.64% 0.78s
+DRLLi (s.) 6.97 6.09% 0.30s 4.34 8.50% 0.36s 3.25 11.68% 0.43s 10.90 5.93% 0.60s 6.65 7.78% 0.76s 4.98 11.91% 0.92s
+2D-Ptr (s.) 6.82 3.81% 0.13s 4.20 5.00% 0.13s 3.09 6.19% 0.14s 10.71 4.08% 0.22s 6.46 4.70% 0.23s 4.68 5.17% 0.24s
+PARCO (s.) 6.82 3.81% 0.05s 4.17 4.25% 0.05s 3.06 5.15% 0.07s 10.61 3.11% 0.08s 6.36 3.08% 0.08s 4.58 2.92% 0.09s
+OMDCPDP
+$N$ 50 100
+$M$ 5 7 10 10 15 20
+Metric Obj. Gap Time Obj. Gap Time Obj. Gap Time Obj. Gap Time Obj. Gap Time Obj. Gap Time
+OR-Tools 37.61 4.61% 30s 30.18 3.38% 30s 24.48 1.79% 30s 66.78 4.99% 60s 51.49 3.15% 60s 43.90 1.53% 60s
+HAM (g.) 39.67 10.30% 0.11s 31.49 7.85% 0.11s 29.24 21.23% 0.13s 71.12 11.60% 0.24s 54.31 8.69% 0.27s 53.73 23.93% 0.29s
+MAPDP (g.) 37.36 4.09% 0.02s 30.36 4.08% 0.01s 24.88 3.46% 0.01s 66.54 4.71% 0.01s 52.08 4.34% 0.01s 44.71 3.40% 0.01s
+PARCO (g.) 37.27 3.84% 0.02s 30.12 3.27% 0.01s 24.72 2.81% 0.01s 65.85 3.67% 0.02s 51.45 3.11% 0.01s 44.46 2.84% 0.01s
+HAM (s.) 36.26 1.17% 1.18s 29.54 1.38% 1.31s 27.77 15.24% 1.37s 66.91 5.29% 2.33s 51.60 3.42% 2.53s 52.24 20.52% 2.72s
+MAPDP (s.) 35.64 0.02% 0.02s 29.03 0.23% 0.02s 23.97 0.35% 0.01s 63.64 0.65% 0.03s 50.07 0.75% 0.03s 43.57 1.13% 0.02s
+PARCO (s.) 35.64 0.00% 0.03s 28.96 0.00% 0.02s 23.89 0.00% 0.02s 63.20 0.00% 0.04s 49.69 0.00% 0.03s 43.08 0.00% 0.03s
+FFSP
+$N$ 20 50 100 50
+$M$ 12 18 24 30
+Metric Obj. Gap Time Obj. Gap Time Obj. Gap Time Obj. Gap Time Obj. Gap Time Obj. Gap Time
+Gurobi (1m) 35.29 42.4% 60s - - 60s - - 60s - - 60s - - 60s - - 60s
+Gurobi (10m) 31.61 27.6% 600s - - 600s - - 600s - - 600s - - 600s - - 600s
+Random 47.89 93.3% 0.18s 93.34 89.4% 0.37s 167.22 86.9% 0.72s 67.03 112.1% 0.33s 54.48 130.9% 0.33s 46.84 137.0% 0.37s
+SJF 31.27 26.2% 0.13s 56.94 15.6% 0.34s 99.27 11.0% 0.62s 38.01 20.3% 0.25s 29.39 24.6% 0.25s 24.62 24.6% 0.29s
+GA 31.15 25.7% 21s 56.92 15.5% 44s 99.25 10.9% 89s 38.26 21.1% 47s 29.05 16.7% 50s 24.52 24.1% 55s
+PSO 29.10 17.4% 46s 55.10 11.8% 82s 97.3 8.8% 154s 36.83 16.6% 85s 28.06 12.7% 89s 23.44 18.6% 95s
+MatNet (g.) 27.26 10.0% 1.22s 51.52 4.6% 2.17s 91.58 2.4% 4.97s 34.82 10.2% 2.42s 27.52 16.7% 2.65s 23.65 19.7% 3.09s
+PARCO (g.) 26.31 6.2% 0.26s 51.19 3.9% 0.52s 91.29 2.0% 0.89s 32.88 4.1% 0.50s 24.89 5.5% 0.44s 20.29 2.7% 0.41s
+MatNet (s.) 25.44 2.7% 3.88s 49.68 0.8% 8.91s 89.72 0.3% 18s 33.45 5.9% 9.23s 26.00 10.2% 9.81s 22.51 13.9% 11s
+PARCO (s.) 24.78 0.0% 0.99s 49.27 0.0% 1.97s 89.46 0.0% 4.04s 31.60 0.0% 1.89s 23.59 0.0% 1.68s 19.76 0.0% 1.54s
+Figure 3: Analysis on PARCO components.
+(a) Full Communication Layers provide the best performance.
+(b) Average number of decoding steps required to construct solutions (bar heights) and training times (lines) for different models.
+(c) Random and Priority-based Conflict Handlers variants with heuristics and learned priorities with learned providing the best performance.
 5.2 Experimental Results
-Main experiments
-The main experiments showcasing the performance of PARCO are in Table 1, Table 2, Table 3 for HCVRP, OMDCPDP, and FFSP, respectively.
-Our PARCO consistently outperforms SotA neural baselines across a variety of experiments.
-We additionally note another property of PARCO: unlike neural baselines in HCVRP and MAPDP in the OMDCPDP, which are trained specifically for a single size and number of agents, PARCO is a single model trained on multiple location and agent distributions at the same time; nonetheless, our single model can outperform baselines trained ad hoc on specific distributions, demonstrating our method’s flexibility.
-Generalization
-We additionally study the generalization performance in Table 2 on the right, which are the sizes and number of agents unseen during training in OMDCPDP.
-Unlike MAPDP, which is constrained to a specific number of agents, PARCO can successfully generalize to unseen sizes and $m$.
-Notably, PARCO can even outperform Google ORTools for larger-scale instances, demonstrating remarkable scalability.
-Table 4: Effect of different ##\## of agents in FFSP.
-$J\times S\times m$ Metric MatNet PARCO
-$50\times 3\times 6$ Obj. 35.69 33.58
-Gap 6.3% 0.0%
-Duration 1.1h 10.3m
-Avg. Steps 200.05 19.10
-$50\times 3\times 8$ Obj. 28.59 25.62
-Gap 11.6% 0.0%
-Duration 1.5h 9.4m
-Avg. Steps 205.95 16.80
-$50\times 3\times 10$ Obj. 25.05 21.46
-Gap 16.7% 0.0%
-Duration 1.8h 7.5m
-Avg. Steps 215.85 13.45
+Table 2: Large-scale generalization for unseen numbers of nodes $N$ and agents $M$ (up to 10×10$\times$10 × of those seen during training).
+$N$ 500 1000
+$M$ 50 75 100 100 150 200
+Metric Obj. Gap Time Obj. Gap Time Obj. Gap Time Obj. Gap Time Obj. Gap Time Obj. Gap Time
+OR-Tools 290.79 7.81% 300s 223.72 4.68% 300s 192.59 1.51% 300s 780.89 33.93% 600s 642.54 36.47% 600s 584.27 36.81% 600s
+HAM (g.) 410.95 55.44% 1.03s 310.17 48.72% 1.13s 204.00 9.91% 1.23s 710.64 40.92% 2.43s 554.15 39.41% 2.66s 388.56 8.88% 2.91s
+PARCO (g.) 268.56 1.41% 0.02s 211.21 1.13% 0.02s 187.37 0.83% 0.01s 510.61 0.81% 0.02s 401.46 0.64% 0.02s 359.98 0.49% 0.02s
+HAM (s.) 409.67 54.95% 1.19s 305.16 46.32% 1.31s 203.50 9.64% 1.39s 708.55 40.50% 2.94s 552.76 39.06% 3.22s 384.10 7.63% 3.46s
+PARCO (s.) 264.38 0.00% 0.03s 208.56 0.00% 0.02s 185.61 0.00% 0.02s 504.30 0.00% 0.79s 397.51 0.00% 0.91s 356.87 0.00% 1.26s
+We present the main empirical results for HCVRP, OMDCPDP, and FFSP in Table 1, reporting the average objective function values (Obj.), gaps to the best-known solutions, and inference times for solving each single instance of each problem.
+For neural baselines, we evaluate both greedy (g.) and sampling (s.) performance, using 1280 sampled solutions for routing problems and 128 for FFSP.
+In HCVRP, PARCO outperforms all neural baselines in solution quality and speed while providing solutions at a fraction of the solving time required by traditional solvers.
+In OMDCPDP, our model surpasses all baselines, including OR-Tools.
+Notably, while the AR baseline HAM struggles with a larger number of agents $M$, PARCO’s parallel AR method maintains strong performance across all scales.
+In FFSP, PARCO outperforms traditional solvers (e.g., Gurobi cannot find solutions in time for $N>20$), dispatching rules, and MatNet in all tested scenarios while being more than 4×4$\times$4 × faster.
+Furthermore, similar to our results in routing, PARCO’s advantage becomes even more pronounced in instances with a larger number of agents where PARCO generates higher-quality schedules through effective agent coordination at a fraction of the cost of MatNet.
+5.3 Analysis
 Effect of Communication Layers
-We showcase the importance of Communication Layers in Fig.
-3 (a).
-We benchmark 1) No Communication (only context features), 2) communication via an MLP, 3) communication via an MHA layer, and 4) Our Communication layer.
-Our Communication Layer consistently outperforms other methods.
+We showcase the importance of Communication Layers in Figure 3(a).
+We benchmark different ways to modify decoder queries (see Section 4.5) with 1) No communication (only context features), 2) MLP, 3) MHA, 4) our transformer-based Communication Layers.
+Our Communication Layers consistently outperform other methods.
+Figure 3(b) further shows the effect of communication, which can not only improve performance but also reduce training time by reducing the number of steps, thereby offsetting the additional latency introduced by the Communication Layers.
 Effect of Conflict Handlers
-Fig.
-3 (b) shows the effect of the Priority-based Conflict Handler with different priorities methods $\bm{p}$ in the OMDCPDP.
-We consider the following: 1) Random: the priority is chosen randomly as in Zong et al.
-(2022), 2) First: the priority is chosen by the agent index $k=1,\dots,m$, 3) Smallest: gives priority to the agent that has traveled the least, 4) Closest: priority is given to the closest agent and 5) High Probability: priority is given to the agent whose probability of selecting the conflicting action is the highest.
-High Probability can consistently outperform other methods.
-We note that this is also the most general and problem-agnostic since methods 2-4 are problem-specific.
-Scalability
-Finally, we showcase PARCO’s scalability in terms of speed in large-scale instances in Fig.
-3 (c) compared to the constructive autoregressive ET (Son et al., 2024).
-Interestingly, PARCO’s gap grows with more agents.
-This is because our method can fully exploit agent parallelism and can thus be a strong candidate for large-scale real-time CO applications.
+We study the effect of conflict handlers in Figure 3(c).
+We compare 1) the random handler from MAPDP and our proposed Priority-based Conflict Handlers in different configurations, namely with priorities based on heuristics, including 2) “smallest” prioritizing the agent with the lowest cost so far, and 3) “closest”, prioritizing agents closer to the corresponding node, and finally 4) “learned” with learned priorities (see Section 4.6).
+PARCO’s Priority-based Conflict Handlers with learned priorities consistently outperform other methods, which also enjoy a relative reduction in the number of steps for constructing a solution, e.g., with a $4\%$ reduction in conflict rates.
+Large-Scale Generalization
+We study the zero-shot large-scale generalization performance of PARCO on OMDCPDP and report the results in Table 2 for out-of-distribution numbers of nodes $N$ and agents $M$, both up to 10×10$\times$10 × those seen in training.
+The autoregressive HAM baseline cannot generalize well to such scales due to the lack of communication and robust parallel construction.
+We note that MAPDP cannot be applied to an unseen number of agents because of its inflexible decoder structure.
+Conversely, our method outperforms all baselines, even including OR-Tools with a 10-minute solving time per instance for $N=1000$, making PARCO a powerful candidate for real-time deployment.
+Figure 4: PARCO vs AR inference time. PARCO can construct solutions faster with more agents $M$ thanks to parallel decoding.
+PARCO vs AR Models Scalability
+Finally, we showcase PARCO’s scalability against autoregressive methods in Figure 4.
+Notably, compared to AR models (e.g., DPN, HAM), PARCO achieves significantly faster solution speeds, with efficiency improving as the number of agents $M$ increases.
+Thanks to parallel decoding, fewer solution construction steps for larger $M$ lead to substantially lower latency.
 6 Conclusion
-In this paper, we introduced a novel approach, PARCO - Parallel AutoRegressive Combinatorial Optimization - that utilizes parallel autoregressive (AR) policies to tackle multi-agent NCO problems.
-We introduced a new Communication Layer to allow for the agents to effectively coordinate their next steps during decoding, alongside a Multiple Pointer Mechanism coupled with a Priority-based Conflicts Handler that can generate feasible solutions efficiently.
-We validate the effectiveness of PARCO through extensive experiments with different CO problems from the routing and scheduling domain and demonstrate its competitive performance against classical heuristic and neural baselines.
-While PARCO is already faster than most evaluated neural baselines and decreases the number of required decoding steps by more than 90% in FFSP50 instances with 10 agents as shown in Table 4, it still does not achieve the potential reduction from $\mathcal{O}(N)$ to $\mathcal{O}(\frac{N}{m})$ in many cases.
-This is mainly attributed to the fact that agents can have conflicts that we currently resolve with the priority-based conflict handler.
-Giving one agent precedence and forcing others to stay in their current position results in more than necessary decoding steps.
-In future work, we will try to address this limitation by constructing a custom loss function, which penalizes conflicts and makes agents learn to avoid them during training.
-Further, we plan to extend our work with learnable improvement methods to enable fast first solutions coupled with powerful local search.
-Given PARCO’s solution quality and efficiency, a further exciting avenue of future research is to couple our model with population-based approaches (Grinsztajn et al., 2022; Chalumeau et al., 2023; Hottung et al., 2024) that could search the latent space for diverse and better solutions.
+We introduced PARCO, a learning model to tackle multi-agent combinatorial optimization problems efficiently via parallel autoregressive solution construction.
+By integrating transformer-based Communication Layers, a Multiple Pointer Mechanism, and Priority-based Conflict Resolution, PARCO enables effective agent coordination and significantly reduces computational latency.
+Our extensive experiments on multi-agent vehicle routing and scheduling demonstrate that PARCO consistently outperforms state-of-the-art learning-based solvers, achieving higher solution quality while maintaining superior efficiency.
+Limitations & Future Work
+Although PARCO can efficiently solve multi-agent CO problems with a defined number of agents $M$, it cannot be directly applied to CO tasks where solutions are constructed via an unspecified $M$.
+In future work, we plan to explore multi-agent CO problems with an unspecified number of agents, which could be achieved by either rolling out a batch of several values of $M$ until an optimal solution is reached or by employing a prediction module to predict an optimal agent number.
 Acknowledgements
-We gratefully acknowledge OMELET for supporting us with computing. We also thank people in the AI4CO open research community who have contributed to RL4CO, which this work is based on.
-Appendix
-[…]
-Appendix A Problem Descriptions
+We thank people in the AI4CO open research community who have contributed to RL4CO, which this work is based on.
+This work was supported by the Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korean government(MSIT)[2022-0-01032, Development of Collective Collaboration Intelligence Framework for Internet of Autonomous Things].
+Impact Statement
+This paper presents work whose goal is to advance the field of Machine Learning.
+There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
+Appendix A Problem Definitions
 A.1 HCVRP
-A.1.1 Problem Definition
 The min-max HCVRP (Heterogeneous Capacitated Vehicle Routing Problem) consists of ${M}$ agents sequentially visiting customers to satisfy their demands, with constraints including each customer can be visited exactly once and the amount of demand satisfied by a single vehicle in a trip cannot exceed its capacity, which can be reloaded by going back to the depot.
 The goal is to minimize the makespan, i.e., the worst route.
-A.1.2 Mathematical Formulation
 Consider a problem with $N+1$ nodes (including $N$ customers and a depot) and ${M}$ vehicles. The depot is indexed as 0, and customers are indexed from 1 to $N$.
@@
 A.2 OMDCPDP
-A.2.1 Problem Definition
 The OMDCPDP (Open Multi-Depot Capacitated Pickup and Delivery Problem) is a practical variant of the pickup and delivery problem in which agents have a stacking limit of orders that can be carried at any given time.
@@
 The goal is to minimize the sum of arrival times to delivery locations, i.e., minimizing the cumulative lateness.
-A.2.2 Mathematical Formulation
 Indices
@@
 A.3 FFSP
-A.3.1 Problem Definition
-The flexible flow shop problem (FFSP) is a challenging and extensively studied optimization problem in production scheduling, involving $N$ jobs that must be processed across $i=1\dots S$ stages, each with multiple machines ($m_{i}>1$).
+The flexible flow shop problem (FFSP) is a challenging and extensively studied optimization problem in production scheduling, involving $N$ jobs that must be processed by a total of $M$ machines divided into $i=1\dots S$ stages, each with multiple machines ($m_{i}>1$).
 Jobs follow a specified sequence through these stages, but within each stage, any available machine can process the job, with the key constraint that no machine can handle more than one job simultaneously.
@@
 The process repeats until all jobs for each stage have been scheduled, and the ultimate goal is to minimize the makespan, i.e., the total time required to complete all jobs.
-A.3.2 Mathematical Formulation
+Mathematical Model
 We use the mathematical model outlined in Kwon et al. (2021) to define the FFSP:
@@
 $Y_{iij}$ $=0$ $i=1,\dots,S;\;j=1,\dots,N$ (31)
-$\sum_{j=1}^{N}\sum_{l=1}^{N}Y_{ilj}$ $=\sum_{k=1}^{m_{i}}\max\left(\sum_{j=1}^{n}(X_{ijk}-1),0\right)$ $i=1,\dots,S$ (32)
+$\sum_{j=1}^{N}\sum_{l=1}^{N}Y_{ilj}$ $=\sum_{k=1}^{m_{i}}\max\left(\sum_{j=1}^{n}(X_{ijk})-1,0\right)$ $i=1,\dots,S$ (32)
 $Y_{ilj}$ $\leq\max\left(\max_{k=1\dots m_{i}}\{X_{ijk}+X_{ilk}\}-1,0\right)$ $i=1,\dots,S;\;j,l=1,2,\dots,N$ (33)
@@
 $C_{ij}+M(1-Y_{ilj})$ $\geq C_{il}+\sum_{k=1}^{m_{i}}p_{ijk}\cdot X_{ijk}$ $i=1,2,\dots,S;\;j,l=1,2,\dots,N$ (38)
-Constraint Explanations: Here, the objective function Eq.
-29 minimizes the makespan of the resulting schedule, that is, the completetion time of the job that finishes last.
-The schedule has to adhere to several constraints: First, constraint set Eq.
-30 ensures that each job is assigned to exactly one machine at each stage.
-Constraint sets Eq.
-31 through Eq.
-35 define the precedence relationships between jobs within a stage.
-Specifically, constraint set Eq.
-31 ensures that a job has no precedence relationship with itself.
-Constraint set Eq.
-32 ensures that the total number of precedence relationships in a stage equals $N-m_{i}$ minus the number of machines with no jobs assigned.
-Constraint set Eq.
-33 dictates that precedence relationships can only exist among jobs assigned to the same machine.
-Additionally, constraint sets Eq.
-34 and Eq.
-35 restrict a job to having at most one preceding job and one following job.
-Moving on, constraint set Eq.
-36 specifies that the completion time of a job in the first stage must be at least as long as its processing time in that stage.
-The relationship between the completion times of a job in consecutive stages is described by constraint set Eq.
-37.
-Finally, constraint set Eq.
-38 ensures that no more than one job can be processed on the same machine simultaneously.
+Constraint Explanations: Here, the objective function Equation 29 minimizes the makespan of the resulting schedule, that is, the completetion time of the job that finishes last.
+The schedule has to adhere to several constraints: First, constraint set Equation 30 ensures that each job is assigned to exactly one machine at each stage.
+Constraint sets Equation 31 through Equation 35 define the precedence relationships between jobs within a stage.
+Specifically, constraint set Equation 31 ensures that a job has no precedence relationship with itself.
+Constraint set Equation 32 ensures that the total number of precedence relationships in a stage equals $N-m_{i}$ minus the number of machines with no jobs assigned.
+Constraint set Equation 33 dictates that precedence relationships can only exist among jobs assigned to the same machine.
+Additionally, constraint sets Equation 34 and Equation 35 restrict a job to having at most one preceding job and one following job.
+Moving on, constraint set Equation 36 specifies that the completion time of a job in the first stage must be at least as long as its processing time in that stage.
+The relationship between the completion times of a job in consecutive stages is described by constraint set Equation 37.
+Finally, constraint set Equation 38 ensures that no more than one job can be processed on the same machine simultaneously.
 Appendix B Experimental Details
@@
 B.1.1 Baselines
+We follow the experimental setup of Liu et al. (2024b) for baselines, with additional baselines hyperparameter details reported in their respective papers.
 SISR
-The Slack Induction by String Removals (SISR) approach (Christiaens and Vanden Berghe, 2020) offers a heuristic method for addressing vehicle routing problems (VRPs), focusing on simplifying the optimization process.
+The Slack Induction by String Removals (SISR) approach (Christiaens & Vanden Berghe, 2020) offers a heuristic method for addressing vehicle routing problems (VRPs), focusing on simplifying the optimization process.
 It combines techniques for route dismantling and reconstruction, along with vehicle fleet minimization strategies.
@@
 GA
-The Genetic Algorithm (GA) (Karakatič and Podgorelec, 2015) is used to address vehicle routing problems (VRPs) and other NP-hard challenges by simulating natural evolutionary processes.
-Particularly effective in Multi-Depot Vehicle Routing Problems (MDVRP), GA quickly generates adequate solutions with reasonable computational resources.
-In our experiment, we follow the same carefully tuned hyperparameters from Liu et al.
-(2024c) with $n=200,\mathrm{iter}=40\times N,P_{m}=0.8,P_{c}=1$.
+The Genetic Algorithm (GA) (Karakatič & Podgorelec, 2015) is used to address vehicle routing problems (VRPs) and other NP-hard challenges by simulating natural evolutionary processes.
+GA generates adequate solutions with reasonable computational resources.
+Our experiment follows the same carefully tuned hyperparameters from (Liu et al., 2024b) with $n=200,\mathrm{iter}=40\times N,P_{m}=0.8,P_{c}=1$.
 SA
@@
 It incorporates local search and the improved 2-opt algorithm to refine routes alongside crossover techniques to speed up convergence.
-In our experiment, we follow the same carefully tuned hyperparameters from Liu et al.
-(2024c) with $T_{0}=100,T_{f}=10^{-7},L=20\times N,\alpha=0.98$.
+In our experiment, we follow the same carefully tuned hyperparameters from (Liu et al., 2024b) with $T_{0}=100,T_{f}=10^{-7},L=20\times N,\alpha=0.98$.
 AM
-The Attention Model (AM) (Kool et al., 2018) applies the attention mechanism to tackle combinatorial optimization problems like the Traveling Salesman and Vehicle Routing Problems.
+The Attention Model (AM) (Kool et al., 2019a) applies the attention mechanism to tackle combinatorial optimization problems like the Traveling Salesman and Vehicle Routing Problems.
 It utilizes attention layers for model improvement and trains using REINFORCE with deterministic rollouts.
@@
 It focuses on equitable workload distribution among multiple agents, applying this strategy to challenges like the min-max multi-agent traveling salesman and pickup and delivery problems.
-In our experiments, we modify the decoder mask in ET to generate feasible solutions for HCVRP, and integrate vehicle features into both the input layer and the context encoder.
+In our experiments, we modify the decoder mask in ET to generate feasible solutions for HCVRP and integrate vehicle features into both the input layer and the context encoder, similarly to the setting of Liu et al.
+(2024b).
+DPN
+The Decoupling-Partition-Navigation (DPN) approach (Zheng et al., 2024a) is a SOTA sequential planning AR baseline that tackles min-max vehicle routing problems (min-max VRPs) by explicitly separating the tasks of customer partitioning and route navigation.
+It introduces a Partition-and-Navigation (P&N) Encoder to learn distinct embeddings for these tasks, an Agent-Permutation-Symmetric (APS) loss to leverage routing symmetries, and a Rotation-Based Positional Encoding to enhance generalization across different depot locations.
+We employ a similar setting as ET.
 DRLLi
@@
 (2022) employs a transformer architecture similar to Kool et al.
-(2018) in which the vehicle and node selection happens in two steps via a two selection decoder, thus requiring two actions.
+(2019a) in which the vehicle and node selection happens in two steps via a two selection decoder, thus requiring two actions.
 We employ their original model with additional context of variable vehicle speeds, noting that in the original setting each model was trained on a single distribution of number of agents ${M}$, each with always the same characteristics.
 2D-Ptr
-The 2D Array Pointer network (2D-Ptr) (Liu et al., 2024c) addresses the heterogeneous capacitated vehicle routing problem (HCVRP) by using a dual-encoder setup to map vehicles and customer nodes effectively.
+The 2D Array Pointer network (2D-Ptr) (Liu et al., 2024b) addresses the heterogeneous capacitated vehicle routing problem (HCVRP) by using a dual-encoder setup to map vehicles and customer nodes effectively.
 This approach facilitates dynamic, real-time decision-making for route optimization.
 Its decoder employs a 2D array pointer for action selection, prioritizing actions over vehicles.
-The model is designed to adapt to changes in vehicle and customer numbers, ensuring robust performance across different scenarios.
-For our study, 2D-Ptr was configured following the original study’s parameters for consistent comparison with 8 heads, 128 hidden dimensions for model structure; $\mathrm{lr}=10^{-4},\lambda_{\mathrm{Adam}}=0.995,\kappa_{\mathrm{L2}}=3$ for turning; 50 epochs, 2500 batches per epoch, 512 instances per batch for training.
+The model is designed to adapt to vehicle and customer numbers changes, ensuring robust performance across different scenarios.
 B.1.2 Datasets
@@
 In PARCO, we select a varying size and number of customer training schemes: at each training step, we sample $N\sim\mathcal{U}(60,100)$ and $m\sim\mathcal{U}(3,7)$.
-As we show in the Table 1, a single PARCO model can outperform baseline models even when they were fitted on a specific distribution.
+As we show in Table 1, a single PARCO model can outperform baseline models even when they were fitted on a specific distribution.
 The coordinates of each customer location $(x_{i},y_{i})$, where $i=1,\ldots,N$, are sampled from a uniform distribution $\mathcal{U}(0.0,1.0)$ within a two-dimensional space.
@@
 The demand $d_{i}$ for each customer $i$ is also drawn from a uniform distribution $\mathcal{U}(1,10)$, with the depot having no demand, i.e., $d_{0}=0$.
-Each vehicle $k$, where $k=1,\ldots,m$, is assigned a capacity $Q_{k}$ sampled from a uniform distribution $\mathcal{U}(20,41)$.
-The speed $f_{k}$ of each vehicle is uniformly distributed within the range $\mathcal{U}(0.5,1.0)$.
+Each vehicle $m$, where $m=1,\ldots,{M}$, is assigned a capacity $Q_{m}$ sampled from a uniform distribution $\mathcal{U}(20,41)$.
+The speed $f_{m}$ of each vehicle is uniformly distributed within the range $\mathcal{U}(0.5,1.0)$.
 Testing
-Testing is performed on the 1280 instances per test setting from Liu et al.
-(2024c).
+Testing is performed on the 1280 instances per $(N,M)$ test setting from Liu et al.
+(2024b).
 In Table 1, (g.) refers to the greedy performance of the model, i.e., taking a single trajectory by taking the maximum action probability; (s.) refers to sampling $1280$ solutions in the latent space and selecting the one with the lowest cost (i.e., highest reward).
-We report optimality gaps to ORTools and solution time in seconds in parentheses.
 B.1.3 PARCO Network Hyperparameters
@@
 Main Encoder.
-we employ $L=3$ attention layers in the encoder, with hidden dimension $d_{h}=128$, $8$ attention heads in the MHA, MLP hidden dimension set to $512$, with RMSNorm (Zhang and Sennrich, 2019) as normalization before the MHA and the MLP.
+we employ $L=3$ attention layers in the encoder, with hidden dimension $d_{h}=128$, $8$ attention heads in the MHA, MLP hidden dimension set to $512$, with RMSNorm (Zhang & Sennrich, 2019) as normalization before the MHA and the MLP.
 Decoder
@@
 The context is the embedding for the depot states, current node states, current time, remaining capacities, time of backing to the depot, and number of visited nodes.
-These features are then employed to update multiple queries $q_{k},~{}k=1,\dots,m$ simultaneously.
-Main Decoder.
+Multiple Pointer Mechanism.
 Similarly to the encoder, we employ the same hidden dimension and number of attention heads for the Multiple Pointer Mechanism.
 Communication Layer
-We employ a single transformer layer with hidden dimension $d_{h}=128$, $8$ attention heads in the MHA, MLP hidden dimension set to $512$, with RMSNorm (Zhang and Sennrich, 2019) as normalization before the MHA and the MLP.
-Unlike the encoder layer, which acts between all $m+N$ nodes, communication layers are lighter because they communicate between $m$ agents.
+We employ a single transformer layer with hidden dimension $d_{h}=128$, $8$ attention heads in the MHA, MLP hidden dimension set to $512$, with RMSNorm (Zhang & Sennrich, 2019) as normalization before the MHA and the MLP.
+Unlike the encoder layer, which acts between all $M+N$ problem tokens, communication layers are lighter because they communicate between ${M}$ agents.
 Agent Handler
-We use the High Probability Handler for managing conflicts: priority is given to the agent whose (log-) probability of selecting the conflicting action $a$ is the highest.
-In other words, priorities $\mathbf{p}_{k}=\log p_{\theta}(a_{k}|x)$ for $k=1,\dots,m$.
+We use the Priority-based Conflict Handler guided by the model output probability for managing conflicts: priority is given to the agent whose probability of selecting the conflicting action is the highest (see Section 4.6).
 B.1.4 PARCO Training Hyperparameters
-For each problem size, we train a single PARCO model that can effectively generalize over multiple size and agent distributions.
-We train PARCO with RL via SymNCO (Kim et al., 2022) with $K=10$ symmetric augmentations as shared REINFORCE baseline for $100$ epochs using the Adam optimizer (Kingma and Ba, 2014) with a total batch size $512$ (using 4 GPUs in Distributed Data Parallel configuration) and an initial learning rate of $10^{-4}$ with a step decay factor of $0.1$ after the $80$th and $95$th epochs.
+Unlike baselines, which are trained and tested on the same distribution, we train a single PARCO model that can effectively generalize over multiple size and agent distributions thanks to our flexible structure.
+We train PARCO with RL via SymNCO (Kim et al., 2022) with $K=10$ symmetric augmentations as shared REINFORCE baseline for $100$ epochs using the Adam optimizer (Kingma & Ba, 2014) with a total batch size $512$ (using 4 GPUs in Distributed Data Parallel configuration) and an initial learning rate of $10^{-4}$ with a step decay factor of $0.1$ after the $80$th and $95$th epochs.
 For each epoch, we sample $4\times 10^{5}$ randomly generated data.
@@
 B.2 OMDCPDP
-B.2.1 Datasets
+The setting introduced in OMDCPDP is a more general and realistic setting than the one introduced in Zong et al.
+(2022), particularly due to the multiple depots and the global lateness objective function which is harder to optimize than vanilla min-sum.
+B.2.1 Baselines
+OR-Tools
+Google OR-Tools (Furnon & Perron, 2024) is an open-source software suite designed to address various combinatorial optimization problems.
+This toolkit offers a comprehensive selection of solvers suitable for linear programming, mixed-integer programming, constraint programming, and routing and scheduling challenges.
+Specifically for routing problems like the OMDCPDP, OR-Tools can integrate additional constraints to enhance solution accuracy.
+For our experiments, we maintained consistent parameters across various problem sizes and numbers of agents.
+We configured the global span cost coefficient to $10,000$, selected PATH_CHEAPEST_ARC as the initial solution strategy, followed by GUIDED_LOCAL_SEARCH for local optimization.
+The solving time was set as $\{30,60,300,600\}$ seconds for $N=\{50,100,500,1000\}$, respectively.
+HAM
+The Heterogeneous Attention Model (HAM) (Li et al., 2021a) utilizes a neural network-integrated with a heterogeneous attention mechanism that distinguishes between the roles of nodes and enforces precedence constraints, ensuring the correct sequence of pickup and delivery nodes.
+This approach helps the deep reinforcement learning model to make informed node selections during route planning.
+We adapt the original model to handle OMDCPDP with multiple agents akin to the sequential planning of Son et al.
+(2024).
+MAPDP
+The Multi-Agent Reinforcement Learning-based Framework for Cooperative Pickup and Delivery Problem (MAPDP) (Zong et al., 2022) introduces a cooperative PDP with multiple vehicle agents.
+This framework is trained via a centralized (MA)RL architecture to generate cooperative decisions among agents, incorporating a paired context embedding to capture the inter-dependency of heterogeneous nodes.
+We adapted the MAPDP to fit our OMDCPDP task, utilizing the same encoder as PARCO to ensure a fair comparison.
+For the decoder and training phases, we kept the same random conflict handler, and we retained the hyperparameters detailed in the original study overall.
+B.2.2 Datasets
 Train data generation
-Neural baselines were trained with the specific number of nodes $2N$ and number of agents $m$ they were tested on.
-In PARCO, we select a varying size and number of customer training schemes: at each training step, we sample $2N\sim\mathcal{U}(50,100)$ and $m\sim\mathcal{U}(10,50)$.
+Neural baselines were trained with the specific number of nodes $N$ and number of agents ${M}$ they were tested on.
+In PARCO, we select a varying size and number of customer training schemes: at each training step, we sample $N\sim\mathcal{U}(50,100)$ and $m\sim\mathcal{U}(5,20)$.
 As we show in the Table 1, a single PARCO model can outperform baseline models even when they were fitted on a specific distribution.
-The coordinates of each customer location $(x_{i},y_{i})$, where $i=1,\ldots,2N$, are sampled from a uniform distribution $\mathcal{U}(0.0,1.0)$ within a two-dimensional space.
+The coordinates of each customer location $(x_{i},y_{i})$, where $i=1,\ldots,N$, are sampled from a uniform distribution $\mathcal{U}(0.0,1.0)$ within a two-dimensional space.
 Similarly, we sample ${M}$ initial vehicle locations from the same distribution.
@@
 Testing
-Testing is performed on 100 new instances for each setting of $N$ and $m$ in Table 2 with the distributions from the training settings.
-We finally test PARCO on large-scale, real-world data from complex, realistic distributions with $32$ instances for each vehicle, pickup, and delivery location in Seoul City, South Korea.
-This dataset is provided in our source code.
-Fig.
-4 provides an example of such data.
-In Table 2, (g.) refers to the greedy performance of the model, i.e., taking a single trajectory by taking the maximum action probability; (s.) refers to sampling $1280$ solutions in the latent space and selecting the one with the lowest cost (i.e., highest reward).
-Figure 4: Real-world instance for the OMDCPDP problem in Seoul City, South Korea, with $N=1000$ locations and $m=100$ agents (green) showing relations (gray) of pickups (red triangles) and their respective deliveries (black crosses).
-B.2.2 Baselines
-OR-Tools
-The OR-Tools (Furnon and Perron, 2024) is open-source software designed to address various combinatorial optimization problems.
-This toolkit offers a comprehensive suite of solvers suitable for linear programming, mixed-integer programming, constraint programming, and routing and scheduling challenges.
-Specifically for routing problems like the OMDCPDP, OR-Tools can integrate additional constraints to enhance solution accuracy.
-Documentation for the Python OR-Tools library, along with example codes for diverse optimization challenges, is readily accessible on their official website.
-For our experiments, we maintained consistent parameters across various problem sizes and numbers of agents.
-We configured the global span cost coefficient to $10,000$, selected PATH_CHEAPEST_ARC as the initial solution strategy, followed by GUIDED_LOCAL_SEARCH for local optimization.
-The solving time was capped at 300 seconds to standardize comparison and measure efficiency.
-HAM
-The Heterogeneous Attention Model (HAM) (Li et al., 2021a) utilizes a neural network integrated with a heterogeneous attention mechanism that distinguishes between the roles of nodes and enforces precedence constraints, ensuring the correct sequence of pickup and delivery nodes.
-This approach helps the deep reinforcement learning model to make informed node selections during route planning.
-We used the same hyperparameter settings from the original HAM experiments.
-MAPDP
-The Multi-Agent Reinforcement Learning-based Framework for Cooperative Pickup and Delivery Problem (MAPDP) (Zong et al., 2022) pioneers the use of multi-agent reinforcement learning (MARL) for the cooperative PDP with multiple vehicle agents.
-This innovative framework introduces a centralized MARL architecture to generate cooperative decisions among agents, incorporating a paired context embedding to capture the inter-dependency of heterogeneous nodes.
-We adapted the MAPDP to fit our OMDCPDP task, utilizing the same encoder as PARCO to ensure a fair comparison.
-For the decoder and training phases, we kept the same random conflict handler and retained the same hyperparameters as detailed in the original study with 8 heads, 128 hidden dimensions for model structure; $\mathrm{lr}=10^{-3},L_{\mathrm{Adam}}=3$ for turning.
+Testing is performed on $1000$ new instances for each setting of in-distribution $N$ and ${M}$ in Table 1 with the distributions from the training settings.
+For large-scale generalization in Table 2, we generate $100$ new instances.
 B.2.3 PARCO Network Hyperparameters
-Most hyperparameters are kept similar to Section B.1.3; we report all of them nonetheless below.
+Most hyperparameters are kept similar to Section B.1.3.
 Encoder
@@
 This layer projects initial raw features to hidden space.
-For depots, the initial embeddings encode the location $o_{k}$ and the respective vehicle’s capacity $Q_{k}$.
+For depots, the initial embeddings encode the location $o_{m}$ and the respective vehicle’s capacity $Q_{m}$.
 For pickup nodes, the initial embeddings encode the location and paired delivery nodes’ location.
@@
 Main Encoder.
-We employ $l=3$ attention layers in the encoder, with hidden dimension $d_{h}=128$, $8$ attention heads in the MHA, MLP hidden dimension set to $512$, with RMSNorm (Zhang and Sennrich, 2019) as normalization before the MHA and the MLP.
+We employ $l=3$ attention layers in the encoder, with hidden dimension $d_{h}=128$, $8$ attention heads in the MHA, MLP hidden dimension set to $512$, with RMSNorm (Zhang & Sennrich, 2019) as normalization before the MHA and the MLP.
 Decoder
@@
 This layer projects dynamic raw features to hidden space.
-The context is the embedding for the depot states $o_{k}$, current node states, current length, remaining capacity, and number of visited nodes.
-These features are then employed to update multiple queries $q_{k},~{}k=1,\dots,m$ simultaneously.
+The context is the embedding for the depot states $o_{m}$, current node states, current length, remaining capacity, and number of visited nodes.
+These features are then employed to update multiple queries $q_{m},~{}m=1,\dots,{M}$ simultaneously.
 Main Decoder.
@@
 Communication Layer
-We employ a single transformer layer with hidden dimension $d_{h}=128$, $8$ attention heads in the MHA, MLP hidden dimension set to $512$, with RMSNorm (Zhang and Sennrich, 2019) as normalization before the MHA and the MLP.
-Note that unlike the encoder layer, which acts between all $m+N$ nodes, communication layers are lighter because they communicate between $m$ agents.
+We employ a single transformer layer with hidden dimension $d_{h}=128$, $8$ attention heads in the MHA, MLP hidden dimension set to $512$, with RMSNorm (Zhang & Sennrich, 2019) as normalization before the MHA and the MLP.
+Note that unlike the encoder layer, which acts between all $M+N$ tokens, communication layers are lighter because they communicate between ${M}$ agents.
 Agent Handler
-We use the High Probability Handler for managing conflicts: priority is given to the agent whose (log-) probability of selecting the conflicting action $a$ is the highest.
-In other words, priorities $\mathbf{p}_{k}=\log p_{\theta}(a_{k}|x)$ for $k=1,\dots,m$.
+We employ the Priority-based Conflict Handler guided by the model output probability for managing conflicts with priority given to the agent whose probability of selecting the conflicting action is the highest (see Section 4.6).
 B.2.4 PARCO Training Hyperparameters
 For each problem size, we train a single PARCO model that can effectively generalize over multiple size and agent distributions.
-We train PARCO with RL via SymNCO (Kim et al., 2022) with $K=8$ symmetric augmentations as shared REINFORCE baseline for $100$ epochs using the Adam optimizer (Kingma and Ba, 2014) with a total batch size $128$ on a single GPU and an initial learning rate of $10^{-4}$ with a step decay factor of $0.1$ after the $80$th and $95$th epochs.
+We train PARCO with RL via SymNCO (Kim et al., 2022) with $K=8$ symmetric augmentations as shared REINFORCE baseline for $100$ epochs using the Adam optimizer (Kingma & Ba, 2014) with a total batch size $128$ on a single GPU and an initial learning rate of $10^{-4}$ with a step decay factor of $0.1$ after the $80$th and $95$th epochs.
 For each epoch, we sample $10^{5}$ randomly generated data.
-Training takes less than 4 hours in our configuration.
+Training takes less than 5 hours in our configuration.
 B.3 FFSP
 B.3.1 Baselines
-MatNet
-The comparison of PARCO with other FFSP baselines is shown in table Table 3.
-Being a neural solver, we benchmark PARCO mainly against MatNet (Kwon et al., 2021), the SotA NCO architecture for the FFSP.
-MatNet is an encoder-decoder architecture, which is inspired by the attention model (Kool et al., 2018).
-It extends the encoder of the attention model with a dual graph attention layer, a horizontal stack of two transformer blocks, capable to encode nodes of different types in a bipartite graph - like machines and jobs in the FFSP.
-Kwon et al.
-(2021) train MatNet using POMO (Kwon et al., 2020).
-CPLEX
-Besides the aforementioned MatNet model, we implement the mathematical model described above in the exact solver CPLEX with a time budget of 60 and 600 seconds per instance.
-However, with both time budgets, CPLEX is only capable of generating solutions to the FFSP20 instances.
-Random and Shortest Job First
-Third, we use different (meta-)heuristics to benchmark PARCO on the FFSP.
+Gurobi
+We implement the mathematical model described above in the exact solver Gurobi (Gurobi Optimization, LLC, 2024) with a time budget of 60 and 600 seconds per instance.
+However, with both time budgets, Gurobi is only capable of generating solutions to the FFSP20 instances, similar to the findings made by Kwon et al.
+(2021) for the CPLEX solver.
+Random and Shortest Job First (SJF)
 The Random and Shortest Job First (SJF) heuristics are simple construction strategies that build valid schedules in an iterative manner.
@@
 Likewise, the SJF proceeds by assigning job-machine pairs with the shortest processing time first.
-Genetic Algorithm
-The Genetic Algorithms (GA) are metaheuristics widely used by the OR community to tackle the FFSP (Kahraman et al., 2008).
+[^2] To obtain results for the heuristics and metaheuristics, we used the implementation of Kwon et al.
+(2021), provided in the official GitHub repository of the paper: https://github.com/yd-kwon/MatNet
+Genetic Algorithm (GA)
+Genetic Algorithms are metaheuristics widely used by the OR community to tackle the FFSP (Kahraman et al., 2008).
 The GA iteratively improves multiple candidate solutions called chromosomes.
@@
 Each instance runs for 1,000 iterations.
-Particle Swarm Optimization
-Finally, Particle Swarm Optimization (PSO) iteratively updates multiple candidate solutions called particles, which are updated by the weighted sum of the inertial value, the local best, and the global best at each iteration (Singh and Mahapatra, 2012).
+Particle Swarm Optimization (PSO)
+Finally, Particle Swarm Optimization iteratively updates multiple candidate solutions called particles, which are updated by the weighted sum of the inertial value, the local best, and the global best at each iteration (Singh & Mahapatra, 2012).
 In this implementation, particles use the same representation as GA chromosomes.
@@
 Like GA, PSO runs for 1,000 iterations per instance.
+MatNet
+We benchmark PARCO mainly against MatNet (Kwon et al., 2021), a state-of-the-art NCO architecture for the FFSP.
+MatNet is an encoder-decoder architecture, which is inspired by the attention model (Kool et al., 2019a).
+It extends the encoder of the attention model with a dual graph attention layer, a horizontal stack of two transformer blocks, capable of encoding nodes of different types in bipartite graph-like machines and jobs in the FFSP.
+Kwon et al.
+(2021) train MatNet using POMO (Kwon et al., 2020).
 B.3.2 Datasets
@@
 (2021) sample processing times for job-machine pairs independently from a uniform distribution within the bounds $[2,10]$.
-For the main experiments on the FFSP shown in Table 3, we also use the same instance sizes as Kwon et al.
-(2021) with $N=20,50\text{ and }100$ jobs (denoted FFSP20, FFSP50 and FFSP100, respectively) that need to be processed in $S=3$ stages on $m_{i}=4$ machines each (i.e.
-$m=12$ machines in total).
-For the agent sensitivity analysis shown in Table 4, we fix the number of jobs to 50 but alter the number of agents.
+For the first three FFSP instance types shown in Table 1 we also use the same instance sizes as Kwon et al.
+(2021) with $N=20,50\text{ and }100$ jobs and $M=12$ machines which are spread evenly over $S=3$ stages.
+To test for agent sensitivity in the FFSP, we fix the number of jobs to $N=50$ but alter the number of agents for the last three instance types shown in Table 1.
 Still, we use $S=3$ for this experiment, but alter the number of machines per stage to $M_{i}=6,8\text{ and }10$, yielding a total of 18, 24 and 30 agents, respectively.
 Testing
-Testing is performed on 1,000 test instances generated according to the above sampling scheme.
-We use the same instances as Kwon et al.
-(2021) for the experiments of Table 3 and generate new ones for the experiments of Table 4.
+Testing is performed on 100 separate test instances generated randomly according to the above generation scheme.
 B.3.3 PARCO Network Hyperparameters
@@
 The machines are regarded as the agents in our PARCO framework.
-As such, their embeddings are used as queries $Q_{t}^{s}$ in the Multiple Pointer Mechanism Eq.
-5, while job embeddings are used as the keys and values.
+As such, their embeddings are used as queries $\bm{q}$ in the Multiple Pointer Mechanism Equation 7, while job embeddings are used as the keys and values.
 In each decoding step, the machine embeddings are fused with a projection of the time the respective machine becomes idle.
-Similarly, job embeddings are augmented with a linear transformation of the time they become available in the respective stage before entering the attention head in Eq.
-4.
+Similarly, job embeddings are augmented with a linear transformation of the time they become available in the respective stage before entering the attention head in Equation 6.
 Communication Layer
@@
 We use the High Probability Handler for managing conflicts: priority is given to the agent whose (log-) probability of selecting the conflicting action is the highest.
-Formally, priorities $\mathbf{p}_{k}=\log p_{\theta}(a_{k}|x)$ for $k=1,\dots,m$.
+Formally, priorities $\mathbf{p}_{m}=\log p_{\theta}(a_{m}|x)$ for $m=1,\dots,{M}$.
 B.3.4 PARCO Training Hyperparameters
-Regarding the training setup, each training instance $i$ is augmented by a factor of 24, and the average makespan over the augmented instances is used as a shared baseline $b^{\text{shared}}_{i}$ for the REINFORCE gradient estimator of Eq.
-9.
-We use the ADAM optimizer (Kingma and Ba, 2014) with a learning rate of $4\times 10^{-4}$, which we alter during training using a cosine annealing scheme.
-We train separate models for FFSP20, FFSP50 and FFSP100 instances for 100, 150 and 200 epochs, respectively, using 1,000 randomly generated instances per epoch split into batches of size 50.111Note: to avoid OOMs, For FFSP100 instances, batches are further split into mini-batches of size 25 whose gradients are accumulated.
-For the learned methods, the times given in parenthesis indicate the time required to train the respective models, whereas the other time information indicates the time required to solve a single instance.
-For the agent sensitivity analysis in Table 4, Matnet, as well as our PARCO model, are trained for 20 epochs on 1,000 instances per epoch.
-Duration in this table indicates the time required for training the respective models on the hardware specified in Section B.4.
+Regarding the training setup, each training instance $i$ is augmented by a factor of 24, and the average makespan over the augmented instances is used as a shared baseline $b^{\text{shared}}_{i}$ for the REINFORCE gradient estimator of Equation 9.
+We use the Adam optimizer (Kingma & Ba, 2014) with a learning rate of $4\times 10^{-4}$, which we alter during training using a cosine annealing scheme.
+We train separate models for the environment configurations used in Table 1.
+We train models corresponding to environments with 20 jobs for 100, with 50 jobs for 150 and with 100 jobs for 200 epochs.
+In each epoch, we train the models using 1,000 randomly generated instances split into batches of size 50.333Note: to avoid OOMs, For FFSP100 instances, batches are further split into mini-batches of size 25 whose gradients are accumulated.
 B.3.5 Diagram for MatNet Decoding vs. PARCO Decoding for the FFSP
 The following figures visualize the decoding for the machines of a given stage using MatNet and PARCO.
-As one can see in Fig.
-6, MatNet requires a decoder forward pass for each machine to schedule a job on each of them.
-In contrast, as detailed in Fig.
-6, PARCO can schedule jobs on all machines simultaneously through its Multiple Pointer Mechanism and Agent Handler, leading to significant efficiency gains.
-Figure 5: A FFSP Decoding Step with MatNet
-Figure 6: A FFSP Decoding Step with PARCO
+As one can see in Figure 5(a), MatNet requires a decoder forward pass for each machine to schedule a job on each of them.
+In contrast, as detailed in Figure 5(b), PARCO can schedule jobs on all machines simultaneously through its Multiple Pointer Mechanism and Agent Handler, leading to significant efficiency gains.
+(a) An FFSP Decoding Step with MatNet
+(b) An FFSP Decoding Step with PARCO
+Figure 5: Comparison of decoding steps for FFSP with MatNet and PARCO.
 B.4 Hardware and Software
 B.4.1 Hardware
-Experiments were carried out on a machine equipped with two Intel(R) Xeon(R) Gold 6338 CPU @ 2.00GHz CPUs with a total $128$ threads and $8$ NVIDIA RTX 4090 graphic cards with $24$ GB of VRAM.
+We experiment on a workstation equipped with 2 Intel(R) Xeon(R) Gold 6338 CPUs and $8$ NVIDIA RTX 4090 graphic cards with $24$ GB of VRAM each.
+Training runs of PARCO take less than 24 hours each.
 During inference, we employ only one CPU and a single GPU.
 B.4.2 Software
-We used Python 3.12, PyTorch 2.4 (Paszke et al., 2019) coupled with PyTorch Lightning (Falcon and The PyTorch Lightning team, 2019) with most code based on the RL4CO library (Berto et al., 2024a).
-The main OS is Ubuntu 22.04.4 LTS.
-There were no major difficulties when trying the code on other platforms, so we expect experiments to be reproducible in other architectures.
+We used Python 3.12, PyTorch 2.5 (Paszke et al., 2019) coupled with PyTorch Lightning (Falcon & The PyTorch Lightning team, 2019) with most code based on the RL4CO library (Berto et al., 2024a).
+The operating system is Ubuntu 24.04 LTS.
 B.5 Source Code
-We value open reproducibility. We provide the source code to reproduce our experiments on Github at https://github.com/ai4co/parco.
-Appendix C Additional Discussion
-Following Section 6, we outline further discussion here.
-We additionally remark that several areas of multi-agent planning, including multi-agent pathfinding (Lin and Ma, 2023; Wang et al., 2023; Tang et al., 2024a) and social robot navigation (Li et al., 2024a; He et al., 2024) employ communication as a tool with great importance, from which we also borrow ideas such as prioritized conflict resolution (Tang et al., 2024b).
-An exciting future direction going forward is the development of foundation models, in which PARCO could aid in improving the efficiency and performance in solving multi-agent CO problems.
-Recent examples in the literature have demonstrated the possibility of scaling up solvers with (self)supervision (Drakulic et al., 2023; Luo et al., 2023; Corsini et al., 2024; Pirnay and Grimm, 2024; Luo et al., 2024), developing multi-task solvers for VRPs (Liu et al., 2024a; Lin et al., 2024; Zhou et al., 2024; Berto et al., 2024b) as well as more general CO tasks (Drakulic et al., 2024; Jiang et al., 2024).
-Such approaches model CO problems as sequential autoregressive decision-making; parallelizing this process with PARCO could enrich solution speed/quality.
-We note that most NCO approaches in e.g., VRPs, mostly focus on solving without explicitly modeling (heterogeneous) agents as in PARCO, which is of practical relevance in several real-world tasks.
-Better heuristics for prioritized planning might also be found utilizing automated discovery via recent LLM approaches for CO (Liu et al., 2024b; Ye et al., 2024b).
-Finally, integrating PARCO in online optimization (Hottung et al., 2021; Choo et al., 2022) and end-to-end pipelines for construction and improvement, as done recently in (Berto et al., 2024a; Kong et al., 2024), would further improve the performance of PARCO given larger computational time budgets.
+B.5.1 Licenses
+We provide the summary of licenses Table 3.
+Table 3: Summary of licenses for used assets.
+Resource Type License
+OR-Tools (Furnon & Perron, 2024) Code Apache License, Version 2.0
+AM (Kool et al., 2019a) Code MIT License
+ET (Son et al., 2024) Code Available on Github
+DPN (Zheng et al., 2024a) Code MIT License
+DRLLi (Li et al., 2022) Code Available on Github
+2D-Ptr (Liu et al., 2024b) Code/Dataset Available on Github
+HAM (Li et al., 2021a) Code MIT License
+Gurobi (Gurobi Optimization, LLC, 2024) Code Commercial license (free for academic use)
+MatNet (Kwon et al., 2021) Code/Dataset MIT License
+RL4CO (Berto et al., 2024a) Library MIT License
+B.5.2 Code Release
+We value open reproducibility and release our source code at https://github.com/ai4co/parco.

### v3

@@
-Parallel AutoRegressive Models for Multi-Agent Combinatorial Optimization
+PARCO: Parallel AutoRegressive Models for Multi-Agent Combinatorial Optimization
 Abstract
 Combinatorial optimization problems involving multiple agents are notoriously challenging due to their NP-hard nature and the necessity for effective agent coordination.
-Despite advancements in learning-based methods, existing approaches often face critical limitations, including suboptimal agent coordination, poor generalizability, and high computational latency.
-To address these issues, we propose Parallel AutoRegressive Combinatorial Optimization (PARCO), a reinforcement learning framework designed to construct high-quality solutions for multi-agent combinatorial tasks efficiently.
-To this end, PARCO integrates three key components: (1) transformer-based communication layers to enable effective agent collaboration during parallel solution construction, (2) a multiple pointer mechanism for low-latency, parallel agent decision-making, and (3) priority-based conflict handlers to resolve decision conflicts via learned priorities.
-We evaluate PARCO in multi-agent vehicle routing and scheduling problems where our approach outperforms state-of-the-art learning methods and demonstrates strong generalization ability and remarkable computational efficiency.
-Machine Learning, ICML
+Despite advancements in learning-based methods, existing approaches often face critical limitations, including suboptimal agent coordination, poor generalization, and high computational latency.
+To address these issues, we propose PARCO (Parallel AutoRegressive Combinatorial Optimization), a general reinforcement learning framework designed to construct high-quality solutions for multi-agent combinatorial tasks efficiently.
+To this end, PARCO integrates three key novel components: (1) transformer-based communication layers to enable effective agent collaboration during parallel solution construction, (2) a multiple pointer mechanism for low-latency, parallel agent decision-making, and (3) priority-based conflict handlers to resolve decision conflicts via learned priorities.
+We evaluate PARCO in multi-agent vehicle routing and scheduling problems, where our approach outperforms state-of-the-art learning methods, demonstrating strong generalization ability and remarkable computational efficiency.
+We make our source code publicly available to foster future research: https://github.com/ai4co/parco.
 1 Introduction
-Combinatorial optimization (CO) problems involve determining an optimal sequence of actions in discrete spaces with several crucial domains, including logistics and supply chain management (Weinand et al., 2022).
+Figure 1: PARCO generates better solutions with higher efficiency through parallel decision making.
+Combinatorial optimization (CO) problems involve determining an optimal sequence of actions in discrete spaces with several crucial domains, including logistics and supply chain management [91].
 Many practical CO problems require a solution to be constructed by coordinating multiple distinct entities (i.e., agents), each with unique characteristics.
 We call such problems multi-agent CO.
-This class of problems naturally arises in real-world applications such as coordinated vehicle routing for disaster management (Liberatore et al., 2014), manufacturing (Jungwattanakit et al., 2008) and last-mile delivery optimization (Archetti & Bertazzi, 2021), where heterogeneous agents must operate under complex constraints.
-CO problems are notoriously hard to solve and cannot generally be solved optimally in polynomial time due to their NP-hardness (Jünger et al., 1995; Peres & Castelli, 2021).
-While traditional methods such as exact and heuristic methods have been developed to solve a variety of problems (Laporte & Osman, 1995; Hejazi & Saghafian, 2005), these approaches often concentrate on single-agent scenarios and require long execution times.
-Moreover, multi-agent CO poses numerous additional challenges such as additional constraints and different optimization objectives including minimizing lateness or the makespan among agents (Soman & Patil, 2020; Mahmoudinazlou & Kwon, 2024).
-Recently, Neural Combinatorial Optimization (NCO) has emerged as a powerful alternative by learning efficient neural solvers (Bengio et al., 2021).
-In particular, Reinforcement Learning (RL) has shown promise due to its ability to learn directly from interactions with CO problems instead of relying on costly labeled datasets in the shape of optimal solutions, and could even outperform traditional approaches by automatically discovering better (neural) heuristics (Bello et al., 2016; Kwon et al., 2020; Kim et al., 2023a).
-Among NCO methods, Autoregressive (AR) models—constructing solutions step-by-step—have garnered attention for their ability to generate solutions for a variety of problems with hard constraints (Kool et al., 2019a; Kwon et al., 2021; Bi et al., 2024).
-This capability is crucial for addressing complex problems with multiple constraints, such as heterogeneous capacity (Baldacci et al., 2008) or precedences in pickup and delivery (Parragh et al., 2008) and machine scheduling problems (Zhang et al., 2020a).
+This class of problems naturally arises in real-world applications such as coordinated vehicle routing for disaster management [60], manufacturing [37] and last-mile delivery optimization [1], where heterogeneous agents must operate under complex constraints.
+CO problems are notoriously hard to solve and cannot generally be solved optimally in polynomial time due to their NP-hardness [36, 79].
+While traditional methods such as exact and heuristic methods have been developed to solve a variety of problems [51, 25], these approaches often concentrate on single-agent scenarios and require long execution times.
+Moreover, multi-agent CO poses additional challenges including additional constraints and different optimization objectives, such as minimizing lateness or the makespan among agents [82, 72].
+Recently, Neural Combinatorial Optimization (NCO) has emerged as a powerful alternative by learning efficient neural solvers [4].
+In particular, Reinforcement Learning (RL) has shown promise due to its ability to learn directly from interactions with CO problems instead of relying on costly labeled datasets in the shape of optimal solutions, and could even outperform traditional approaches by automatically discovering better (neural) heuristics [3, 49, 40].
+Among NCO methods, Autoregressive (AR) models—constructing solutions step-by-step—have garnered attention for their ability to generate solutions for a variety of problems with hard constraints [47, 50, 7].
+This capability is crucial for addressing complex problems with multiple constraints, such as heterogeneous capacity [2] or precedences in pickup and delivery [77] and machine scheduling problems [101].
 However, despite their promising results in single-agent settings, AR-CO methods pose two challenges that hinder their practical adoption in solving complex multi-agent CO problems.
-First, existing AR methods for multi-agent CO suffer from poor coordination, resulting in unsatisfactory solution quality and poor generalization across varying problem sizes and agent configurations (Zong et al., 2022; Son et al., 2024; Zheng et al., 2024a; Liu et al., 2024b).
-Moreover, AR sequence generation is associated with high latency due to each single action (or “token”) depending on each single previous one, akin to large language models (Leviathan et al., 2023; Gao et al., 2024).
+First, existing AR methods for multi-agent CO suffer from poor coordination, resulting in unsatisfactory solution quality and poor generalization across varying problem sizes and agent configurations [108, 84, 104, 63].
+Moreover, AR sequence generation is associated with high latency due to each single action (or “token”) depending on each previous one, akin to large language models [52, 21].
 This issue becomes increasingly pronounced when dealing with large problem instances involving numerous agents.
-Figure 1: Example multi-agent vehicle routing problem.
-[Top] Autoregressive policies construct solutions one agent at a time, with inefficient coordination and requiring several steps.
-[Bottom] Our parallel autoregressive policies ensure not only improved decision-making thanks to inter-agent communication that enables effective coordination but also faster construction in fewer steps.
-This paper introduces PARCO (Parallel AutoRegressive Combinatorial Optimization), a novel learning framework to address multi-agent combinatorial problems effectively via parallel solution construction as illustrated in Figure 1.
+This paper introduces PARCO (Parallel AutoRegressive Combinatorial Optimization), a novel learning framework to address multi-agent combinatorial problems effectively via parallel solution construction as illustrated in Fig.˜1.
 We design specialized transformer-based Communication Layers to enhance coordination among agents during decision-making, enabling collaborative behaviors.
@@
   • We enhance solution quality with Priority-based Conflict Handlers that tie break with learned priorities.
-  • We evaluate PARCO on multi-agent vehicle routing and scheduling, where we outperform state-of-the-art learning methods across diverse problems [^1] Code at: https://github.com/ai4co/parco.
+  • We evaluate PARCO on multi-agent vehicle routing and scheduling, where we outperform state-of-the-art learning methods in solution quality, generalization, and efficiency.
 2 Related Work
 Neural Combinatorial Optimization
-Recent advancements in Neural Combinatorial Optimization (NCO) have shown promising end-to-end solutions for combinatorial optimization problems (Bengio et al., 2021; Yang & Whinston, 2023).
-NCO has led to the development of a variety of methods for diverse problems, including the incorporation of problem-specific biases (Jin et al., 2023; Luo et al., 2023; Kim et al., 2023c, b; Son et al., 2023; Gao et al., 2023), bi-level solution pipelines (Li et al., 2021b; Ye et al., 2023b; Zheng et al., 2024b), learning-guided search (Ye et al., 2023a; Kim et al., 2024; Sun & Yang, 2023; Li et al., 2024b, 2025), improvement methods (Hottung et al., 2021; Ma et al., 2021, 2024), effective training algorithms (Kim et al., 2022; Grinsztajn et al., 2023; Drakulic et al., 2023; Corsini et al., 2024; Pirnay & Grimm, 2024; Luo et al., 2024), downstream applications (Chen et al., 2023; Zhou et al., 2023; Ma et al., 2023; Luttmann & Xie, 2024; Yan & Wu, 2024), and the recent development of end-to-end foundation models (Liu et al., 2024a; Zhou et al., 2024; Drakulic et al., 2024; Berto et al., 2024b; Jiang et al., 2024; Li et al., 2024a).
-Among such methods, RL-based end-to-end autoregressive (AR) models present several advantages, including eliminating the need for labeled solutions, reducing reliance on handcrafted heuristics, and achieving high efficiency in generating high-quality solutions (Kool et al., 2019a; Kwon et al., 2020, 2021; Berto et al., 2024a; Bonnet et al., 2024).
+Recent advancements in Neural Combinatorial Optimization (NCO) have shown promising end-to-end solutions for combinatorial optimization problems [4, 94].
+NCO has led to the development of a variety of methods for diverse problems, including the incorporation of problem-specific biases [35, 64, 44, 41, 83, 20, 31], bi-level solution pipelines [56, 96, 105], learning-guided search [95, 45, 85, 57, 58, 59, 42], improvement methods [26, 69, 70, 28], effective training algorithms [43, 23, 14, 13, 80, 65, 103], downstream applications [11, 106, 71, 66, 93], and the recent development of end-to-end foundation models [61, 107, 15, 6, 34, 53, 22].
+Among such methods, RL-based end-to-end autoregressive (AR) models present several advantages, including eliminating the need for labeled solutions, reducing reliance on handcrafted heuristics, and achieving high efficiency in generating high-quality solutions [47, 49, 50, 5, 8].
 Multi-Agent AR Methods for CO
 While the seminal works in AR-CO methods of Vinyals et al.
-(2015); Kool et al.
-(2019a); Kwon et al.
-(2020, 2021) propose models that can be used in loose multi-agent settings, such methods cannot be employed directly to model heterogeneous agents with different attributes and constraints.
+[89], Kool et al.
+[47], Kwon et al.
+[49, 50] propose models that can be used in loose multi-agent settings, such methods cannot be employed directly to model heterogeneous agents with different attributes and constraints.
 Building on Kool et al.
-(2019a), Son et al.
-(2024) and Zheng et al.
-(2024a) introduce attention-based policies for multi-agent min-max routing.
+[47], Son et al.
+[84] and Zheng et al.
+[104] introduce attention-based policies for multi-agent min-max routing.
 These models adopt a sequential AR construction strategy, solving for one agent at a time and switching agents only after completing a single-agent solution.
-While this approach outperforms decentralized methods (Cao et al., 2022; Park et al., 2023) it remains inherently sequential.
-In contrast, other multi-agent AR methods determine a location for every agent in each solution construction step, but the different agents select their actions sequentially in either random (Zhang et al., 2020b) or learned order (Falkner & Schmidt-Thieme, 2020; Li et al., 2022; Liu et al., 2024b), making these models still suffer from high generation latency as well as missing inter-agent communication and coordination.
+While this approach outperforms decentralized methods [10, 76] it remains inherently sequential.
+In contrast, other multi-agent AR methods determine a location for every agent in each solution construction step, but the different agents select their actions sequentially in either random [102] or learned order [18, 55, 63], making these models still suffer from high generation latency as well as missing inter-agent communication and coordination.
 Zong et al.
-(2022) propose a multi-agent pickup and delivery model with parallel decoding, using distinct decoders for each agent.
+[108] propose a multi-agent pickup and delivery model with parallel decoding, using distinct decoders for each agent.
 However, this approach exhibits limited generalizability due to inflexible fixed decoders for specific agents, lacking a powerful communication mechanism, and conflict resolution handled naively by assigning random precedence to agents, restricting the robustness of the model.
-PARCO aims at addressing the shortcomings of previous works by leveraging parallel solution construction for any number of agents efficiently with a Multiple Pointer Mechanism, enhancing coordination via Communication Layers, and solving conflicts in a principled manner via Priority-based Conflict handlers.
-Finally, unlike previous works, PARCO tackles a general framework tackling multi-agent CO without restricting to a single class of problems.
+PARCO addresses the shortcomings of previous works by leveraging parallel solution construction for any number of agents efficiently with a Multiple Pointer Mechanism, enhancing coordination via Communication Layers, and solving conflicts in a principled manner via Priority-based Conflict handlers.
+Finally, unlike previous works, PARCO is a general framework tackling multi-agent CO without restricting to a single class of problems.
 3 Preliminaries
 3.1 Markov Decision Processes
-Multi-agent CO problems can be formulated as Markov Decision Processes (MDPs) and solved autoregressively using RL (Mazyavkina et al., 2021).
+Multi-agent CO problems can be formulated as Markov Decision Processes (MDPs) and solved autoregressively using RL [74].
 In this framework, a solution $\bm{a}$ to a CO problem instance $\bm{x}$ is represented as a sequence of actions.
@@
 Agents are selected either by some predefined precedence rule as in the sequential planning of Son et al.
-(2024), where agent solutions are constructed one after another, or by the policy itself, in which case $\pi_{\theta}:\mathcal{S}\rightarrow\mathcal{A}\times\mathcal{M}$.
+[84], where agent solutions are constructed one after another, or by the policy itself, in which case $\pi_{\theta}:\mathcal{S}\rightarrow\mathcal{A}\times\mathcal{M}$.
 Given the agent and its corresponding action, the problem then transitions from state $s_{t}$ to state $s_{t+1}$ according to a transition function $\tau:\mathcal{S}\times\mathcal{A}\times\mathcal{M}\rightarrow\mathcal{S}$.
@@
 3.2 AR Models for CO
-Given the sequential nature of MDPs, AR methods pose a natural choice for the policy $\pi_{\theta}$.
+Given the sequential nature of MDPs, autoregressive (AR) models pose a natural choice for the policy $\pi_{\theta}$.
 AR methods construct a viable solution by sequentially generating actions based on the current state and previously selected actions.
@@
 The parameters $\theta$ encompass both the encoder and decoder components, which together define the policy as the mechanism for producing the joint distribution $p_{\theta}(\bm{a}|\bm{x})$.
-3.3 RL Training
-RL leverages the MDP formulation to learn the parameters $\theta$ of the policy $\pi_{\theta}$ without relying on labeled optimal solutions. We can thus cast the optimization problem as follows:
-[…]
-where $P(\bm{x})$ is the distribution over problem instances and $R(\bm{a},\bm{x})$ is the reward of a solution.
-We can solve the optimization problem in Equation 2 using various RL methods, including policy gradient (REINFORCE) variations (Kool et al., 2019b; Kwon et al., 2020; Kim et al., 2022).
+Thus, the RL objective becomes finding the optimal set of parameters $\theta^{*}$ that maximizes the reward function $R$ [48, 49, 43].
 4 Methodology
-We now outline the general structure of PARCO as shown in Figure 2.
-First, we formally define parallel multi-agent MDPs for CO (Section 4.1), which we use as a basis to derive our overall parallel autoregressive approach (Section 4.2).
-We then describe in detail the components of our model: Multi-Agent Encoder (Section 4.3), Communication Layers (Section 4.4), Decoder with Multiple Pointer Mechanism (Section 4.5) and Conflict Handlers (Section 4.6).
-Finally, we outline the training scheme (Section 4.7).
+We now outline the general structure of PARCO as shown in Fig.˜2.
+First, we formally define parallel multi-agent MDPs for CO (§˜4.1), which we use as a basis to derive our overall parallel autoregressive approach (§˜4.2).
+We then describe in detail the components of our model: Multi-Agent Encoder (§˜4.3), Communication Layers (§˜4.4), Decoder with Multiple Pointer Mechanism (§˜4.5) and Conflict Handlers (§˜4.6).
+Finally, we outline the training scheme (§˜4.7).
 Figure 2: Overview of PARCO.
@@
 Actions are decoded efficiently autoregressively in parallel through a Multiple Pointer Mechanism enhanced by a Priority-based Conflict Handler.
-4.1 Parallel Multi-Agent MDPs
-We reformulate the multi-agent MDPs of Section 3.1 as parallel multi-agent MDPs by selecting multiple actions from a joint action space simultaneously to enhance efficiency and coordination (Yu et al., 2022).
-At each step $t$, ${M}$ agents select actions $\bm{a}_{t}=(a_{t}^{1},...,a_{t}^{M})$ according to a policy $\pi_{\theta}:\mathcal{S}\rightarrow\mathcal{A}_{1}\times...\times\mathcal{A}_{M}$, which maps the state space $\mathcal{S}$ to the joint action space of agents.
+4.1 Cooperative Multi-Agent MDPs
+We reformulate the MDPs of §˜3.1 as cooperative multi-agent MDPs [9], often termed fully cooperative Markov games [73], by selecting multiple actions from a joint action space simultaneously to enhance efficiency and coordination [98, 68].
+At each step $t$, M{M} agents select actions $\bm{a}_{t}=(a_{t}^{1},...,a_{t}^{M})$ according to a policy $\pi_{\theta}:\mathcal{S}\rightarrow\mathcal{A}_{1}\times...\times\mathcal{A}_{M}$, which maps the state space $\mathcal{S}$ to the joint action space of agents.
 A conflict handling function $\psi:(\mathcal{A}_{1}\times...\times\mathcal{A}_{M})\rightarrow(\mathcal{A}_{1}\times...\times\mathcal{A}_{M})$ ensures action compatibility by allowing only one agent to execute when multiple agents select mutually conflicting actions (e.g., the same customer location), and assigning a fallback action (e.g., staying at current position) to the others.
@@
 In the following, we refer to the entities corresponding to actions in the MDP formulation of a CO problem (e.g., customer locations in VRPs) as nodes, following the convention of Kool et al.
-(2019a) and Kwon et al.
-(2020).
+[47] and Kwon et al.
+[49].
 4.2 Parallel AR Models for CO
@@
 PARCO constructs feasible solutions by simultaneously generating multiple agent actions based on the current state.
-We formulate the solution generation process in an encoder-decoder framework similarly to Equation 1:
+We formulate the solution generation process in an encoder-decoder framework similarly to Eq.˜1:
 $p_{\theta}(\bm{a}|\bm{x})$ $\triangleq\prod_{t=1}^{T}\psi\left(\prod_{m=1}^{M}g_{\theta}(a_{t}^{m}|\bm{a}_{<t},\bm{h})\right)$ (2)
@@
 While the latter require $\sum_{m=1}^{M}T_{m}$ total actions to construct a solution with $T_{m}$ being the number of steps required by agent $m$ to finish its task, PARCO’s Parallel AR needs only $\max_{m}T_{m}$ steps as agents effectively divide the solution space and act concurrently.
-This leads to a faster solution construction as illustrated in Figures 1 and 5 and significantly reduced training times.
+This leads to a faster solution construction as illustrated in Figs.˜1 and 6 and significantly reduced training times.
 4.3 Multi-Agent Encoder
@@
 In PARCO, we explicitly model agents and employ separate agent and node embedding layers similar to Son et al.
-(2024) to project agents and nodes into the same embedding space.
+[84] to project agents and nodes into the same embedding space.
 The agent embedding layer projects $k_{a}$ agent features – such as vehicle locations and capacities (routing) or machine characteristics (scheduling) – into a $d$-dimensional space using a linear projection ${\mathbf{W}}_{a}\in\mathbb{R}^{k_{a}\times d}$.
@@
 Depending on the problem structure, the initial agent and node embeddings might either be concatenated as $\bm{h}^{(0)}=\text{Concat}(\bm{h}_{a}^{(0)},\bm{h}_{n}^{(0)})$ and passed through $L$ transformer blocks, consisting of multi-head self-attention $\text{MHA}(\bm{h}^{(0)},\bm{h}^{(0)},\bm{h}^{(0)})$ and multi-layer perceptrons (MLPs) as defined in Vaswani et al.
-(2017).
-Or, agent and node embeddings are used separately as query and keys/values, respectively, in a cross-attention mechanism $\text{MHA}(\bm{h}_{a}^{(0)},\bm{h}_{n}^{(0)},\bm{h}_{n}^{(0)})$ akin to MatNet (Kwon et al., 2021).
+[88].
+Or, agent and node embeddings are used separately as query and keys/values, respectively, in a cross-attention mechanism $\text{MHA}(\bm{h}_{a}^{(0)},\bm{h}_{n}^{(0)},\bm{h}_{n}^{(0)})$ akin to MatNet [50].
 The final embeddings $\bm{h}=\{\bm{h}_{a},\bm{h}_{n}\}$ emitted by the last encoder layer contain processed representations of both agents and nodes that capture their interactions as well as the overall problem structure.
 4.4 Communication Layers
-At each step $t$ of the decoding process of Equation 3, given the encoded representations $\bm{h}$, we construct dynamic agent queries that capture the current state of both agents and the environment.
+At each step $t$ of the decoding process of Eq.˜2, given the encoded representations $\bm{h}$, we construct dynamic agent queries that capture the current state of both agents and the environment.
 For each agent $m$, we form a context embedding $\bm{d}_{m}=\text{Concat}(\bm{h}_{a^{m}},\bm{h}_{\delta^{m}_{t}},\bm{h}_{e})$ consisting of the following components: (1) the (static) embedding of the agent $\bm{h}_{a^{m}}$; (2) a projection $\bm{h}_{\delta^{m}_{t}}=\delta^{m}_{t}{\mathbf{W}}_{\delta}\in\mathbb{R}^{d}$ of the agent’s dynamics $\delta^{m}_{t}$ like its current location and capacity (routing) or the time until the agent becomes idle (scheduling); (3) a projection $\bm{h}_{e}=e_{t}{\mathbf{W}}_{\delta}\in\mathbb{R}^{d}$ of the dynamic environment features $e_{t}$ that encode the current problem’s state.
@@
 $\bm{q}$ $=\text{Norm}(\text{MLP}((\bm{q}^{\prime}))+\bm{q}^{\prime})$ (4)
-where Norm denotes a normalization layer (Ioffe & Szegedy, 2015; Zhang & Sennrich, 2019).
-These layers are inherently agent-count agnostic, which allows PARCO to handle an arbitrary number of agents, making it more flexible and generalizable across different problems.
+where Norm denotes a normalization layer [33, 100].
+These layers are inherently agent-count agnostic, allowing PARCO to handle arbitrary numbers of agents, making it more flexible and generalizable across different problems.
 Communication Layers allow agents to coordinate their actions by attending to both the problem structure and other agents’ states while maintaining efficiency through parallel processing.
 4.5 Decoder with Multiple Pointer Mechanism
-PARCO’s decoder extends the attention-based pointer mechanism of Kool et al.
-(2019a) – originally designed for single-agent scenarios and recently applied to sequential multi-agent planning with a single agent at a time (Son et al., 2024; Zheng et al., 2024a) – to handle multiple agents operating in parallel via a Multiple Pointer Mechanism.
+PARCO’s decoder improves the AR pointer mechanism [89, 47] – originally designed for single-agent scenarios and recently applied to sequential multi-agent planning with a single agent at a time [84, 104] – to handle multiple agents operating in parallel via a Multiple Pointer Mechanism.
 Starting with the processed agent queries $\bm{q}$ that underwent communication, we first compute agent-specific representations through masked cross MHA:
-[…]
+$\bm{q}^{\prime}=\text{MHA}(\bm{q},\,\bm{h}_{n}+\xi_{t}{\mathbf{W}}^{K}_{\xi},\,\bm{h}_{n}+\xi_{t}{\mathbf{W}}^{V}_{\xi};\,\mathbf{M}_{t})$ (5)
 where $\xi_{t}\in\mathbb{R}^{N\times k_{\xi}}$ are dynamic node features which are projected via ${\mathbf{W}}^{K}_{\xi},{\mathbf{W}}^{V}_{\xi}\in\mathbb{R}^{k_{\xi}\times d}$ for keys and values of the MHA, respectively.
@@
 We then obtain a joint logit space $\bm{u}$ across all agents:
-$\bm{u}={\beta}\cdot\text{tanh}\left(\frac{\bm{q}(\bm{h}_{n}{\mathbf{W}}^{L}+\xi_{t}{\mathbf{W}}^{L}_{\xi})^{\top}}{\sqrt{d}}\right)$ (7)
+$\bm{u}={\beta}\cdot\text{tanh}\left(\frac{\bm{q}^{\prime}(\bm{h}_{n}{\mathbf{W}}^{L}+\xi_{t}{\mathbf{W}}^{L}_{\xi})^{\top}}{\sqrt{d}}\right)$ (6)
 with learnable parameters ${\mathbf{W}}^{L}\in\mathbb{R}^{d\times d}$, ${\mathbf{W}}^{L}_{\xi}\in\mathbb{R}^{k_{\xi}\times d}$ and $\beta$} is a scale parameter, set to $10$ following Bello et al.
-(2016) to enhance exploration.
-The output logits $\bm{u}\in\mathbb{R}^{M\times N}$ are masked by setting infeasible actions given mask $\textbf{M}_{t}$ to minus infinity.The joint probability distribution over all agent actions is then computed as:
+[3] to enhance exploration.
+The output logits $\bm{u}\in\mathbb{R}^{M\times N}$ are masked by setting infeasible actions given mask $\textbf{M}_{t}$ to -$\infty$.
+The joint probability distribution over all agent actions becomes:
 $p(\bm{a}_{t}|\bm{a}_{<t},\bm{h})=\prod_{m=1}^{M}\frac{\exp(\bm{u}_{m,a_{t}^{m}})}{\sum_{j=1}^{N}\exp(\bm{u}_{m,j})}$ (7)
@@
 4.6 Conflict Handlers
-When sampling from the probability distribution $\bm{p}$ generated by the Multiple Pointer Mechanism, multiple agents may select the same action simultaneously, which can result in an infeasible solution in several CO problems: for instance, in vehicle routing problems, usually only one agent is allowed to visit a customer node.
-Thus, it is essential to choose how to deal with such a situation effectively.
+When sampling from the probability distribution $\bm{p}$ generated by the Multiple Pointer Mechanism, multiple agents may select the same action simultaneously, which can result in an infeasible solution in several CO problems – for instance, in vehicle routing problems, usually only one agent is allowed to visit a customer node – and it becomes essential how to deal with such a situation effectively.
 Conflict handling (i.e., tie-breaking) can be achieved by allowing a single agent among a number of agents that are in conflict to continue with its action while others revert to fallback actions, e.g., staying in their current position.
-A naive approach consists of randomly selecting an agent to perform the new action (Zong et al., 2022).
+A simple approach introduced by Zong et al.
+[108] consists of randomly selecting an agent to perform the new action.
 However, this can be suboptimal since it excludes inductive biases that can be leveraged, such as learned representations.
 Algorithm 1 Priority-based Conflict Handler
-0:  Actions $\bm{a}\in\mathbb{N}^{M}$, Priorities $\bm{p}\in\mathbb{R}^{M}$, Fallback actions $\bm{r}\in\mathbb{R}^{M}$
-0:  Resolved Actions $\bm{a}^{\prime}\in\mathbb{N}^{M}$ // Sort indices based on priorities in descending order.
-1:  $\sigma\leftarrow\text{argsort}(\bm{p},\text{descending}=\text{True})$ // Reorder actions according to priority.
-2:  $\hat{\bm{a}}\leftarrow\bm{a}[\sigma]$ // Initialize conflict mask (all zeros).
-3:  $C\leftarrow\bm{0}^{M}$ // Check for conflicts in reordered actions.
-4:  for $i=2$ to ${M}$ do
-5:     if $\hat{\bm{a}}_{i}\in\{\hat{\bm{a}}_{1},\ldots,\hat{\bm{a}}_{i-1}\}$ then
-6:        $C_{i}\leftarrow 1$
-7:     end if
-8:  end for// Resolve conflicts by assigning fallback actions.
-9:  $\hat{\bm{a}}\leftarrow(1-C)\odot\hat{\bm{a}}+C\odot\bm{r}$ // Reorder resolved actions back to original order.
-10:  $\bm{a}^{\prime}\leftarrow\hat{\bm{a}}[\sigma^{-1}]$
+1:Actions $\bm{a}\in\mathbb{N}^{M}$, Priorities $\bm{p}\in\mathbb{R}^{M}$, Fallback actions $\bm{r}\in\mathbb{R}^{M}$
+2:Resolved Actions $\bm{a}^{\prime}\in\mathbb{N}^{M}$
+[…]
+[…]
+[…]
+6:for $i=2$ to M{M} do // Check for conflicts in reordered actions
+7:  if $\hat{\bm{a}}_{i}\in\{\hat{\bm{a}}_{1},\ldots,\hat{\bm{a}}_{i-1}\}$ then
+8:   $C_{i}\leftarrow 1$ // $C_{i}=1$ indicates a conflict for index $i$
+9:  end if
+10:end for
+[…]
+[…]
 In PARCO, we propose Priority-based Conflict Handlers that leverage priorities as a tie-breaking rule.
 Such priorities can be based on heuristics – such as giving priority to agents close to completion or whose action results in the smallest immediate cost – or on learned priorities.
-In the latter case, the probability values of the selected actions $p(\bm{a}_{t})$ serve as an indicator for prioritizing certain agents: the higher its value, the higher the priority learned by the model to have that agent win the tie-break.
-Algorithm 1 shows an efficient vectorized implementation of the Priority-based Conflict Handler algorithm.
-In practice, we augment the conflict handler $\psi$ from Section 4.2 with priorities $\bm{p}\coloneq p(\bm{a}_{t})$ and fallback actions $\bm{r}$, i.e.
+In the latter case, the model output probability values of the selected actions $p(\bm{a}_{t})$ serve as an indicator for prioritizing certain agents: the higher their value, the higher the priority learned by the model to have those agents win the tie-break.
+Algorithm˜1 shows our efficient vectorized implementation of the Priority-based Conflict Handler algorithm.
+In practice, we augment the conflict handler $\psi$ from §˜4.2 with priorities $\bm{p}\coloneq p(\bm{a}_{t})$ and fallback actions $\bm{r}$, i.e.
 $\psi(\cdot)\coloneq\psi(\bm{a}_{t},\bm{p},\bm{r})$.
@@
 PARCO is a centralized multi-agent decision-making framework and can thus be trained by RL algorithms proposed in the single-agent NCO literature.
-We train PARCO by employing the REINFORCE gradient estimator (Williams, 1992) with a shared baseline as outlined by Kwon et al.
-(2020) and Kim et al.
-(2022):
+We train PARCO by employing the REINFORCE gradient estimator [92] with a shared baseline as outlined by Kwon et al.
+[49] and Kim et al.
+[43]:
 $\nabla_{\theta}\mathcal{L}\approx\frac{1}{B\cdot S}\sum_{i=1}^{B}\sum_{j=1}^{S}G_{ij}\nabla_{\theta}\,\text{log}\,p_{\theta}(\bm{a}_{ij}|\bm{x}_{i})$ (8)
@@
 Specifically, we evaluate its performance on two challenging routing problems – the min-max heterogeneous capacitated vehicle routing problem (HCVRP) and the open multi-depot capacitated pickup and delivery problem (OMDCPDP) – as well as a scheduling problem, the flexible flow shop problem (FFSP).
+Experimental details are available through Appendix˜B [^1] Source code is available at https://github.com/ai4co/parco.
 5.1 Experimental Settings
-We experiment on a workstation equipped with 2 Intel(R) Xeon(R) Gold 6338 CPUs and $8$ NVIDIA RTX 4090 graphic cards with $24$ GB of VRAM each.
-Training runs of PARCO take less than 24 hours each.
-During inference, we employ only one CPU and a single GPU.
-Additional details are available throughout Appendix B.
 HCVRP
 Problem.
-The min-max HCVRP involves ${M}$ agents serving customer demands while adhering to heterogeneous vehicle capacity constraints.
+The min-max HCVRP involves M{M} agents serving customer demands while adhering to heterogeneous vehicle capacity constraints.
 Each vehicle can replenish its load by returning to the depot.
@@
 Traditional solvers.
-We include state-of-the-art SISRs (Christiaens & Vanden Berghe, 2020), Genetic Algorithm (GA) (Karakatič & Podgorelec, 2015) and Simulated Annealing (SA) (İlhan, 2021).
+We include state-of-the-art SISRs [12], Genetic Algorithm (GA) [39] and Simulated Annealing (SA) [32].
 Neural baselines.
-We evaluate the sequential planning baselines Attention Model (AM) (Kool et al., 2019a), Equity Transformer (ET) (Son et al., 2024) and Decoupling Partition and Navigation (DPN) (Zheng et al., 2024a), and autoregressive models with agent selection DRLLi (Li et al., 2022) and state-of-the-art learning method 2D-Ptr (Liu et al., 2024b).
-Additional problem and experimental details are available in Section A.1 and Section B.1, respectively.
+We evaluate the sequential planning baselines Attention Model (AM) [47], Equity Transformer (ET) [84] and Decoupling Partition and Navigation (DPN) [104], and autoregressive models with agent selection DRLLi [55] and state-of-the-art learning method 2D-Ptr [63].
+Additional problem and experimental details are available in §˜A.1 and §˜B.1, respectively.
 OMDCPDP
@@
 Traditional solvers.
-We include the popular and efficient optimization suite Google OR-Tools (Furnon & Perron, 2024) as a classical baseline.
+We include the popular and efficient optimization suite Google OR-Tools [19] as a classical baseline.
 Neural baselines.
-We evaluate models specializing in pickup and delivery problems, including the autoregressive Heterogeneous Attention Model (HAM) (Li et al., 2021a) for sequential planning and MAPDP (Zong et al., 2022) for parallel planning.
-Additional problem and experimental details are available in Section A.2 and Section B.2, respectively.
+We evaluate models specializing in pickup and delivery problems, including the autoregressive Heterogeneous Attention Model (HAM) [54] for sequential planning and MAPDP [108] for parallel planning.
+Additional problem and experimental details are available in §˜A.2 and §˜B.2, respectively.
 FFSP
@@
 Traditional solvers.
-We incorporate the widely used and powerful Gurobi solver (Gurobi Optimization, LLC, 2024) as a baseline.
-Furthermore, we include dispatching rules Random and Shortest Job First (SJF), Particle Swarm Optimization (PSO) (Singh & Mahapatra, 2012) and Genetic Algorithm (GA) (Hejazi & Saghafian, 2005).
+We incorporate the widely used and powerful Gurobi solver [24] as a baseline.
+Furthermore, we include dispatching rules Random and Shortest Job First (SJF), Particle Swarm Optimization (PSO) [81] and Genetic Algorithm (GA) [25].
 Neural baselines.
-Notable benchmarks include the Matrix Encoding Network (MatNet) (Kwon et al., 2021) which demonstrates superior performance on FFSP.
-Additional problem and experimental details are available in Section A.3 and Section B.3, respectively.
-Table 1: Main results with different configurations in terms of problem size $N$ and number of agents $M$. The lower the better ($\downarrow$).
+Notable benchmarks include the Matrix Encoding Network (MatNet) [50] which demonstrates superior performance on FFSP.
+Additional problem and experimental details are available in §˜A.3 and §˜B.3, respectively.
+Table 1: Main results on different problems with different configurations for problem size $N$ and number of agents $M$. For all metrics, the lower the better ($\downarrow$).
 HCVRP
-$N$ 60 100
+NN 60 100
 $M$ 3 5 7 3 5 7
@@
 OMDCPDP
-$N$ 50 100
+NN 50 100
 $M$ 5 7 10 10 15 20
@@
 FFSP
-$N$ 20 50 100 50
+NN 20 50 100 50
 $M$ 12 18 24 30
@@
 PARCO (s.) 24.78 0.0% 0.99s 49.27 0.0% 1.97s 89.46 0.0% 4.04s 31.60 0.0% 1.89s 23.59 0.0% 1.68s 19.76 0.0% 1.54s
-Figure 3: Analysis on PARCO components.
-(a) Full Communication Layers provide the best performance.
-(b) Average number of decoding steps required to construct solutions (bar heights) and training times (lines) for different models.
-(c) Random and Priority-based Conflict Handlers variants with heuristics and learned priorities with learned providing the best performance.
 5.2 Experimental Results
-Table 2: Large-scale generalization for unseen numbers of nodes $N$ and agents $M$ (up to 10×10$\times$10 × of those seen during training).
-$N$ 500 1000
+We report the main empirical results for HCVRP, OMDCPDP, and FFSP in Table˜1, with average objective function values (Obj.), gaps to the best-known solutions, and inference times for solving each single problem instance.
+For neural baselines, we evaluate both greedy (g.) and sampling (s.) performance, using 1280 sampled solutions for routing problems and 128 for FFSP.
+In HCVRP, PARCO outperforms all neural baselines in solution quality and speed while providing solutions at a fraction of the solving time required by traditional solvers.
+In OMDCPDP, our model surpasses all baselines, including OR-Tools.
+Notably, while the AR baseline HAM struggles with a larger number of agents $M$, PARCO’s parallel AR method maintains strong performance across all scales.
+In FFSP, PARCO outperforms traditional solvers (e.g., Gurobi cannot find solutions in time for $N>20$), dispatching rules, and MatNet in all tested scenarios while being more than $4\times$ faster.
+Furthermore, similar to our results in routing, PARCO’s advantage becomes even more pronounced in instances with a larger number of agents where PARCO generates higher-quality schedules through effective agent coordination at a fraction of the cost of MatNet.
+5.3 Analysis
+Effect of Communication Layers
+We showcase the importance of Communication Layers in Fig.˜3(a).
+We benchmark different ways to obtain decoder queries (see §˜4.5) with 1) No communication (W/o Comm., i.e., with only context features), 2) MLP, 3) MHA, 4) our transformer-based Communication Layers (Comm.).
+Our Communication Layers consistently outperform other methods.
+Fig.˜3(b) shows decoding steps and training times on the FFSP for $N=50$.
+PARCO greatly reduces the number of steps and training times, with Communication Layers further reducing them through better coordination, especially at a higher number of agents $M$.
+Table 2: Generalization for unseen numbers of nodes $N$ and agents $M$ (up to $10\times$ those seen during training).
+NN 500 1000
 $M$ 50 75 100 100 150 200
@@
 PARCO (s.) 264.38 0.00% 0.03s 208.56 0.00% 0.02s 185.61 0.00% 0.02s 504.30 0.00% 0.79s 397.51 0.00% 0.91s 356.87 0.00% 1.26s
-We present the main empirical results for HCVRP, OMDCPDP, and FFSP in Table 1, reporting the average objective function values (Obj.), gaps to the best-known solutions, and inference times for solving each single instance of each problem.
-For neural baselines, we evaluate both greedy (g.) and sampling (s.) performance, using 1280 sampled solutions for routing problems and 128 for FFSP.
-In HCVRP, PARCO outperforms all neural baselines in solution quality and speed while providing solutions at a fraction of the solving time required by traditional solvers.
-In OMDCPDP, our model surpasses all baselines, including OR-Tools.
-Notably, while the AR baseline HAM struggles with a larger number of agents $M$, PARCO’s parallel AR method maintains strong performance across all scales.
-In FFSP, PARCO outperforms traditional solvers (e.g., Gurobi cannot find solutions in time for $N>20$), dispatching rules, and MatNet in all tested scenarios while being more than 4×4$\times$4 × faster.
-Furthermore, similar to our results in routing, PARCO’s advantage becomes even more pronounced in instances with a larger number of agents where PARCO generates higher-quality schedules through effective agent coordination at a fraction of the cost of MatNet.
-5.3 Analysis
-Effect of Communication Layers
-We showcase the importance of Communication Layers in Figure 3(a).
-We benchmark different ways to modify decoder queries (see Section 4.5) with 1) No communication (only context features), 2) MLP, 3) MHA, 4) our transformer-based Communication Layers.
-Our Communication Layers consistently outperform other methods.
-Figure 3(b) further shows the effect of communication, which can not only improve performance but also reduce training time by reducing the number of steps, thereby offsetting the additional latency introduced by the Communication Layers.
+(a) Effect of communication
+(b) Decoding steps and training time
+(c) Effect of conflict handlers
+Figure 3: Analysis of PARCO components.
 Effect of Conflict Handlers
-We study the effect of conflict handlers in Figure 3(c).
-We compare 1) the random handler from MAPDP and our proposed Priority-based Conflict Handlers in different configurations, namely with priorities based on heuristics, including 2) “smallest” prioritizing the agent with the lowest cost so far, and 3) “closest”, prioritizing agents closer to the corresponding node, and finally 4) “learned” with learned priorities (see Section 4.6).
-PARCO’s Priority-based Conflict Handlers with learned priorities consistently outperform other methods, which also enjoy a relative reduction in the number of steps for constructing a solution, e.g., with a $4\%$ reduction in conflict rates.
+In Fig.˜3(c), we compare 1) the random handler from MAPDP and our proposed Priority-based Conflict Handlers in different configurations, namely with priorities based on simple heuristics as 2) “smallest” prioritizing the agent with the lowest cost so far, and 3) “closest”, prioritizing agents closer to the corresponding node, and finally 4) “learned” based on model output probabilities.
+The latter consistently outperforms other methods, which also enjoy a relative reduction in the number of steps for constructing a solution, e.g., with a $4\%$ reduction in conflict rates.
 Large-Scale Generalization
-We study the zero-shot large-scale generalization performance of PARCO on OMDCPDP and report the results in Table 2 for out-of-distribution numbers of nodes $N$ and agents $M$, both up to 10×10$\times$10 × those seen in training.
-The autoregressive HAM baseline cannot generalize well to such scales due to the lack of communication and robust parallel construction.
-We note that MAPDP cannot be applied to an unseen number of agents because of its inflexible decoder structure.
-Conversely, our method outperforms all baselines, even including OR-Tools with a 10-minute solving time per instance for $N=1000$, making PARCO a powerful candidate for real-time deployment.
-Figure 4: PARCO vs AR inference time. PARCO can construct solutions faster with more agents $M$ thanks to parallel decoding.
+We study the zero-shot large-scale generalization performance of PARCO in the OMDCPDP and report the results in Table˜2 for out-of-distribution numbers of nodes $N$ and agents $M$, both up to $10\times$ those seen in training.
+We find that the AR HAM baseline cannot generalize well to such scales due to the lack of communication and robust parallel construction, while MAPDP cannot be applied to an unseen $M$ because of its inflexible decoder structure.
+Figure 4: PARCO vs AR inference time. PARCO constructs solutions faster with more agents $M$.
+Conversely, our method outperforms all baselines, including OR-Tools with a 10-minute solving time per instance for $N=1000$, making PARCO a strong candidate for real-time deployment.
 PARCO vs AR Models Scalability
-Finally, we showcase PARCO’s scalability against autoregressive methods in Figure 4.
-Notably, compared to AR models (e.g., DPN, HAM), PARCO achieves significantly faster solution speeds, with efficiency improving as the number of agents $M$ increases.
-Thanks to parallel decoding, fewer solution construction steps for larger $M$ lead to substantially lower latency.
+Finally, we showcase PARCO’s speedups against autoregressive methods in Fig.˜4.
+Notably, compared to AR models (e.g., DPN, HAM), PARCO achieves significant speedups of $3.3\times$ up to $24.7\times$, with inference time decreasing as the number of agents $M$ increases: thanks to parallel decoding, fewer solution construction steps for larger $M$ lead to substantially lower latency.
 6 Conclusion
@@
 By integrating transformer-based Communication Layers, a Multiple Pointer Mechanism, and Priority-based Conflict Resolution, PARCO enables effective agent coordination and significantly reduces computational latency.
-Our extensive experiments on multi-agent vehicle routing and scheduling demonstrate that PARCO consistently outperforms state-of-the-art learning-based solvers, achieving higher solution quality while maintaining superior efficiency.
+Our extensive experiments on multi-agent vehicle routing and scheduling demonstrate that PARCO consistently outperforms state-of-the-art learning-based solvers with better solution quality and higher efficiency.
 Limitations & Future Work
@@
 In future work, we plan to explore multi-agent CO problems with an unspecified number of agents, which could be achieved by either rolling out a batch of several values of $M$ until an optimal solution is reached or by employing a prediction module to predict an optimal agent number.
+We defer additional discussions to §˜C.1.
 Acknowledgements
-We thank people in the AI4CO open research community who have contributed to RL4CO, which this work is based on.
-This work was supported by the Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korean government(MSIT)[2022-0-01032, Development of Collective Collaboration Intelligence Framework for Internet of Autonomous Things].
-Impact Statement
-This paper presents work whose goal is to advance the field of Machine Learning.
-There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
+We are deeply grateful to the members of the AI4CO open research community for their invaluable contributions to PARCO and related projects, including RL4CO.
+We also thank the anonymous reviewers who greatly helped improve our paper with their constructive feedback.
+This work was supported by the Institute of Information & Communications Technology Planning & Evaluation (IITP) grant, funded by the Korean government (MSIT) [Grant No.
+2022-0-01032, Development of Collective Collaboration Intelligence Framework for Internet of Autonomous Things]; National Research Foundation of Korea(NRF) grants funded by the Korea government(MSIT) (No.
+RS-2024-00410082 and No.
+RS-2025-00563763), and by the InnoCORE program of the Ministry of Science and ICT(N10250154).
+Supplementary Material
 Appendix A Problem Definitions
 A.1 HCVRP
-The min-max HCVRP (Heterogeneous Capacitated Vehicle Routing Problem) consists of ${M}$ agents sequentially visiting customers to satisfy their demands, with constraints including each customer can be visited exactly once and the amount of demand satisfied by a single vehicle in a trip cannot exceed its capacity, which can be reloaded by going back to the depot.
+The min-max HCVRP (Heterogeneous Capacitated Vehicle Routing Problem) consists of M{M} agents sequentially visiting customers to satisfy their demands, with constraints including each customer can be visited exactly once and the amount of demand satisfied by a single vehicle in a trip cannot exceed its capacity, which can be reloaded by going back to the depot.
 The goal is to minimize the makespan, i.e., the worst route.
-Consider a problem with $N+1$ nodes (including $N$ customers and a depot) and ${M}$ vehicles. The depot is indexed as 0, and customers are indexed from 1 to $N$.
+Consider a problem with $N+1$ nodes (including $N$ customers and a depot) and M{M} vehicles. The depot is indexed as 0, and customers are indexed from 1 to $N$.
 Indices
@@
 Parameters
-$N$ Number of customer nodes (excluding depot)
-${M}$ Number of vehicles
+NN Number of customer nodes (excluding depot)
+M{M} Number of vehicles
 $X_{i}$ Location of node $i$
@@
 Decision Variables
-$x_{ijk}$$\left\{\begin{array}[]{ll}1&\text{if vehicle }k\text{ travels directly from node }i\text{ to node }j\\ 0&\text{otherwise}\end{array}\right.$
+[…]
 $l_{ijk}$ Remaining load of vehicle $k$ before travelling from node $i$ to node $j$
@@
 Constraint Explanations: The formulation is subject to several constraints that define the feasible solution space.
-Equation (11) ensures that each customer is visited exactly once by one vehicle.
+Eq.˜10 ensures that each customer is visited exactly once by one vehicle.
 The flow conservation constraint (11) guarantees that each vehicle that enters a node also leaves that node, maintaining route continuity.
@@
 Parameters
-$N$ Number of pickup-delivery pairs
-${M}$ Number of vehicles
+NN Number of pickup-delivery pairs
+M{M} Number of vehicles
 $c_{ij}$ Travel time between nodes $i$ and $j$
@@
 Decision Variables
-$x_{ijk}$$\left\{\begin{array}[]{ll}1&\text{if vehicle }k\text{ travels directly from node }i\text{ to node }j\\ 0&\text{otherwise}\end{array}\right.$
+[…]
 $y_{ik}$ $\left\{\begin{array}[]{ll}1&\text{if vehicle }k\text{ visits node }i\\ 0&\text{otherwise}\end{array}\right$.
@@
 $t_{i}$ $\geq 0$ $i=1,\ldots,2N$ (27)
-Constraints Explanations: Equation (18) ensures that each node is visited exactly once.
+Constraints Explanations: Eq.˜17 ensures that each node is visited exactly once.
 Constraint (18) guarantees that each vehicle starts from its designated depot.
 The flow conservation constraint (19) ensures route continuity for each vehicle.
-Equation (21) defines the relationship between x and y variables.
+Eq.˜20 defines the relationship between x and y variables.
 Time consistency is enforced by constraint (21), while (22) ensures that pickups are visited before their corresponding deliveries.
@@
 Finally, the vehicle capacity constraint (25) ensures that the load never exceeds the vehicle’s stacking limit.
+Visualization
+We provide a visualization of a large-scale instance in Fig.˜5.
+Figure 5: Real-world instance for the OMDCPDP problem in Seoul City, South Korea, with $N=1000$ locations and $m=100$ agents ( ) showing relations () of pickups ( ) and their respective deliveries ( ).
 A.3 FFSP
@@
 Mathematical Model
-We use the mathematical model outlined in Kwon et al. (2021) to define the FFSP:
+We use the mathematical model outlined in Kwon et al. [50] to define the FFSP:
 Indices
@@
 Parameters
-$N$ Number of jobs
+NN Number of jobs
 $S$ Number of stages
@@
 $C_{ij}$ Completion time of job $j$ in stage $i$
-$X_{ijk}$$\left\{\begin{array}[]{ll}1&\text{if job }j\text{ is assigned to machine }k\text{ in stage }i\\ 0&\text{otherwise}\end{array}\right.$
-$Y_{ilj}$$\left\{\begin{array}[]{ll}1&\text{if job }l\text{ is processed earlier than job }j\text{ in stage }i\\ 0&\text{otherwise}\end{array}\right.$
+[…]
+[…]
 Objective:
@@
 $C_{ij}+M(1-Y_{ilj})$ $\geq C_{il}+\sum_{k=1}^{m_{i}}p_{ijk}\cdot X_{ijk}$ $i=1,2,\dots,S;\;j,l=1,2,\dots,N$ (37)
-Constraint Explanations: Here, the objective function Equation 29 minimizes the makespan of the resulting schedule, that is, the completetion time of the job that finishes last.
-The schedule has to adhere to several constraints: First, constraint set Equation 30 ensures that each job is assigned to exactly one machine at each stage.
-Constraint sets Equation 31 through Equation 35 define the precedence relationships between jobs within a stage.
-Specifically, constraint set Equation 31 ensures that a job has no precedence relationship with itself.
-Constraint set Equation 32 ensures that the total number of precedence relationships in a stage equals $N-m_{i}$ minus the number of machines with no jobs assigned.
-Constraint set Equation 33 dictates that precedence relationships can only exist among jobs assigned to the same machine.
-Additionally, constraint sets Equation 34 and Equation 35 restrict a job to having at most one preceding job and one following job.
-Moving on, constraint set Equation 36 specifies that the completion time of a job in the first stage must be at least as long as its processing time in that stage.
-The relationship between the completion times of a job in consecutive stages is described by constraint set Equation 37.
-Finally, constraint set Equation 38 ensures that no more than one job can be processed on the same machine simultaneously.
+Constraint Explanations: Here, the objective function Eq.˜28 minimizes the makespan of the resulting schedule, that is, the completion time of the job that finishes last.
+The schedule has to adhere to several constraints: First, constraint set (29) ensures that each job is assigned to exactly one machine at each stage.
+Constraint sets (30) through (34) define the precedence relationships between jobs within a stage.
+Specifically, constraint set (30) ensures that a job has no precedence relationship with itself.
+Constraint set (31) ensures that the total number of precedence relationships in a stage equals $N-m_{i}$ minus the number of machines with no jobs assigned.
+Constraint set (32) dictates that precedence relationships can only exist among jobs assigned to the same machine.
+Additionally, constraint sets (33) and (34) restrict a job to having at most one preceding job and one following job.
+Moving on, constraint set (35) specifies that the completion time of a job in the first stage must be at least as long as its processing time in that stage.
+The relationship between the completion times of a job in consecutive stages is described by constraint set (36).
+Finally, constraint set (37) ensures that no more than one job can be processed on the same machine simultaneously.
 Appendix B Experimental Details
@@
 B.1.1 Baselines
-We follow the experimental setup of Liu et al. (2024b) for baselines, with additional baselines hyperparameter details reported in their respective papers.
+We follow the experimental setup of Liu et al. [63] for baselines, with additional baselines hyperparameter details reported in their respective papers.
 SISR
-The Slack Induction by String Removals (SISR) approach (Christiaens & Vanden Berghe, 2020) offers a heuristic method for addressing vehicle routing problems (VRPs), focusing on simplifying the optimization process.
+The Slack Induction by String Removals (SISR) approach [12] offers a heuristic method for addressing vehicle routing problems (VRPs), focusing on simplifying the optimization process.
 It combines techniques for route dismantling and reconstruction, along with vehicle fleet minimization strategies.
@@
 GA
-The Genetic Algorithm (GA) (Karakatič & Podgorelec, 2015) is used to address vehicle routing problems (VRPs) and other NP-hard challenges by simulating natural evolutionary processes.
+The Genetic Algorithm (GA) [39] is used to address vehicle routing problems (VRPs) and other NP-hard challenges by simulating natural evolutionary processes.
 GA generates adequate solutions with reasonable computational resources.
-Our experiment follows the same carefully tuned hyperparameters from (Liu et al., 2024b) with $n=200,\mathrm{iter}=40\times N,P_{m}=0.8,P_{c}=1$.
+Our experiment follows the same carefully tuned hyperparameters from [63] with $n=200,\mathrm{iter}=40\times N,P_{m}=0.8,P_{c}=1$.
 SA
-The Simulated Annealing (SA) method (İlhan, 2021) targets the capacitated vehicle routing problem (CVRP) using a population-based approach combined with crossover operators.
+The Simulated Annealing (SA) method [32] targets the capacitated vehicle routing problem (CVRP) using a population-based approach combined with crossover operators.
 It incorporates local search and the improved 2-opt algorithm to refine routes alongside crossover techniques to speed up convergence.
-In our experiment, we follow the same carefully tuned hyperparameters from (Liu et al., 2024b) with $T_{0}=100,T_{f}=10^{-7},L=20\times N,\alpha=0.98$.
+In our experiment, we follow the same carefully tuned hyperparameters from [63] with $T_{0}=100,T_{f}=10^{-7},L=20\times N,\alpha=0.98$.
 AM
-The Attention Model (AM) (Kool et al., 2019a) applies the attention mechanism to tackle combinatorial optimization problems like the Traveling Salesman and Vehicle Routing Problems.
+The Attention Model (AM) [47] applies the attention mechanism to tackle combinatorial optimization problems like the Traveling Salesman and Vehicle Routing Problems.
 It utilizes attention layers for model improvement and trains using REINFORCE with deterministic rollouts.
@@
 ET
-The Equity-Transformer (ET) approach (Son et al., 2024) addresses large-scale min-max routing problems by employing a sequential planning approach with sequence generators like the Transformer.
+The Equity-Transformer (ET) approach [84] addresses large-scale min-max routing problems by employing a sequential planning approach with sequence generators like the Transformer.
 It focuses on equitable workload distribution among multiple agents, applying this strategy to challenges like the min-max multi-agent traveling salesman and pickup and delivery problems.
 In our experiments, we modify the decoder mask in ET to generate feasible solutions for HCVRP and integrate vehicle features into both the input layer and the context encoder, similarly to the setting of Liu et al.
-(2024b).
+[63].
 DPN
-The Decoupling-Partition-Navigation (DPN) approach (Zheng et al., 2024a) is a SOTA sequential planning AR baseline that tackles min-max vehicle routing problems (min-max VRPs) by explicitly separating the tasks of customer partitioning and route navigation.
+The Decoupling-Partition-Navigation (DPN) approach [104] is a SOTA sequential planning AR baseline that tackles min-max vehicle routing problems (min-max VRPs) by explicitly separating the tasks of customer partitioning and route navigation.
 It introduces a Partition-and-Navigation (P&N) Encoder to learn distinct embeddings for these tasks, an Agent-Permutation-Symmetric (APS) loss to leverage routing symmetries, and a Rotation-Based Positional Encoding to enhance generalization across different depot locations.
@@
 The DRL approach for solving HCVRP by Li et al.
-(2022) employs a transformer architecture similar to Kool et al.
-(2019a) in which the vehicle and node selection happens in two steps via a two selection decoder, thus requiring two actions.
-We employ their original model with additional context of variable vehicle speeds, noting that in the original setting each model was trained on a single distribution of number of agents ${M}$, each with always the same characteristics.
+[55] employs a transformer architecture similar to Kool et al.
+[47] in which the vehicle and node selection happens in two steps via a two selection decoder, thus requiring two actions.
+We employ their original model with additional context of variable vehicle speeds, noting that in the original setting each model was trained on a single distribution of number of agents M{M}, each with always the same characteristics.
 2D-Ptr
-The 2D Array Pointer network (2D-Ptr) (Liu et al., 2024b) addresses the heterogeneous capacitated vehicle routing problem (HCVRP) by using a dual-encoder setup to map vehicles and customer nodes effectively.
+The 2D Array Pointer network (2D-Ptr) [63] addresses the heterogeneous capacitated vehicle routing problem (HCVRP) by using a dual-encoder setup to map vehicles and customer nodes effectively.
 This approach facilitates dynamic, real-time decision-making for route optimization.
@@
 Train data generation
-Neural baselines were trained with the specific number of nodes $N$ and number of agents ${M}$ they were tested on.
+Neural baselines were trained with the specific number of nodes $N$ and number of agents M{M} they were tested on.
 In PARCO, we select a varying size and number of customer training schemes: at each training step, we sample $N\sim\mathcal{U}(60,100)$ and $m\sim\mathcal{U}(3,7)$.
-As we show in Table 1, a single PARCO model can outperform baseline models even when they were fitted on a specific distribution.
+As we show in Table˜1, a single PARCO model can outperform baseline models even when they were fitted on a specific distribution.
 The coordinates of each customer location $(x_{i},y_{i})$, where $i=1,\ldots,N$, are sampled from a uniform distribution $\mathcal{U}(0.0,1.0)$ within a two-dimensional space.
@@
 Testing is performed on the 1280 instances per $(N,M)$ test setting from Liu et al.
-(2024b).
-In Table 1, (g.) refers to the greedy performance of the model, i.e., taking a single trajectory by taking the maximum action probability; (s.) refers to sampling $1280$ solutions in the latent space and selecting the one with the lowest cost (i.e., highest reward).
+[63].
+In Table˜1, (g.) refers to the greedy performance of the model, i.e., taking a single trajectory by taking the maximum action probability; (s.) refers to sampling $1280$ solutions in the latent space and selecting the one with the lowest cost (i.e., highest reward).
 B.1.3 PARCO Network Hyperparameters
@@
 Main Encoder.
-we employ $L=3$ attention layers in the encoder, with hidden dimension $d_{h}=128$, $8$ attention heads in the MHA, MLP hidden dimension set to $512$, with RMSNorm (Zhang & Sennrich, 2019) as normalization before the MHA and the MLP.
+we employ $L=3$ attention layers in the encoder, with hidden dimension $d_{h}=128$, $8$ attention heads in the MHA, MLP hidden dimension set to $512$, with RMSNorm [100] as normalization before the MHA and the MLP.
 Decoder
@@
 Communication Layer
-We employ a single transformer layer with hidden dimension $d_{h}=128$, $8$ attention heads in the MHA, MLP hidden dimension set to $512$, with RMSNorm (Zhang & Sennrich, 2019) as normalization before the MHA and the MLP.
-Unlike the encoder layer, which acts between all $M+N$ problem tokens, communication layers are lighter because they communicate between ${M}$ agents.
+We employ a single transformer layer with hidden dimension $d_{h}=128$, $8$ attention heads in the MHA, MLP hidden dimension set to $512$, with RMSNorm [100] as normalization before the MHA and the MLP.
+Unlike the encoder layer, which acts between all $M+N$ problem tokens, communication layers are lighter because they communicate between M{M} agents.
 Agent Handler
-We use the Priority-based Conflict Handler guided by the model output probability for managing conflicts: priority is given to the agent whose probability of selecting the conflicting action is the highest (see Section 4.6).
+We use the Priority-based Conflict Handler guided by the model output probability for managing conflicts: priority is given to the agent whose probability of selecting the conflicting action is the highest (see §˜4.6).
 B.1.4 PARCO Training Hyperparameters
 Unlike baselines, which are trained and tested on the same distribution, we train a single PARCO model that can effectively generalize over multiple size and agent distributions thanks to our flexible structure.
-We train PARCO with RL via SymNCO (Kim et al., 2022) with $K=10$ symmetric augmentations as shared REINFORCE baseline for $100$ epochs using the Adam optimizer (Kingma & Ba, 2014) with a total batch size $512$ (using 4 GPUs in Distributed Data Parallel configuration) and an initial learning rate of $10^{-4}$ with a step decay factor of $0.1$ after the $80$th and $95$th epochs.
+We train PARCO with RL via SymNCO [43] with $K=10$ symmetric augmentations as shared REINFORCE baseline for $100$ epochs using the Adam optimizer [46] with a total batch size $512$ (using 4 GPUs in Distributed Data Parallel configuration) and an initial learning rate of $10^{-4}$ with a step decay factor of $0.1$ after the $80$th and $95$th epochs.
 For each epoch, we sample $4\times 10^{5}$ randomly generated data.
@@
 The setting introduced in OMDCPDP is a more general and realistic setting than the one introduced in Zong et al.
-(2022), particularly due to the multiple depots and the global lateness objective function which is harder to optimize than vanilla min-sum.
+[108], particularly due to the multiple depots and the global lateness objective function which is harder to optimize than vanilla min-sum.
 B.2.1 Baselines
 OR-Tools
-Google OR-Tools (Furnon & Perron, 2024) is an open-source software suite designed to address various combinatorial optimization problems.
+Google OR-Tools [19] is an open-source software suite designed to address various combinatorial optimization problems.
 This toolkit offers a comprehensive selection of solvers suitable for linear programming, mixed-integer programming, constraint programming, and routing and scheduling challenges.
@@
 HAM
-The Heterogeneous Attention Model (HAM) (Li et al., 2021a) utilizes a neural network-integrated with a heterogeneous attention mechanism that distinguishes between the roles of nodes and enforces precedence constraints, ensuring the correct sequence of pickup and delivery nodes.
+The Heterogeneous Attention Model (HAM) [54] utilizes a neural network-integrated with a heterogeneous attention mechanism that distinguishes between the roles of nodes and enforces precedence constraints, ensuring the correct sequence of pickup and delivery nodes.
 This approach helps the deep reinforcement learning model to make informed node selections during route planning.
 We adapt the original model to handle OMDCPDP with multiple agents akin to the sequential planning of Son et al.
-(2024).
+[84].
 MAPDP
-The Multi-Agent Reinforcement Learning-based Framework for Cooperative Pickup and Delivery Problem (MAPDP) (Zong et al., 2022) introduces a cooperative PDP with multiple vehicle agents.
+The Multi-Agent Reinforcement Learning-based Framework for Cooperative Pickup and Delivery Problem (MAPDP) [108] introduces a cooperative PDP with multiple vehicle agents.
 This framework is trained via a centralized (MA)RL architecture to generate cooperative decisions among agents, incorporating a paired context embedding to capture the inter-dependency of heterogeneous nodes.
@@
 Train data generation
-Neural baselines were trained with the specific number of nodes $N$ and number of agents ${M}$ they were tested on.
+Neural baselines were trained with the specific number of nodes $N$ and number of agents M{M} they were tested on.
 In PARCO, we select a varying size and number of customer training schemes: at each training step, we sample $N\sim\mathcal{U}(50,100)$ and $m\sim\mathcal{U}(5,20)$.
-As we show in the Table 1, a single PARCO model can outperform baseline models even when they were fitted on a specific distribution.
+As we show in the Table˜1, a single PARCO model can outperform baseline models even when they were fitted on a specific distribution.
 The coordinates of each customer location $(x_{i},y_{i})$, where $i=1,\ldots,N$, are sampled from a uniform distribution $\mathcal{U}(0.0,1.0)$ within a two-dimensional space.
-Similarly, we sample ${M}$ initial vehicle locations from the same distribution.
+Similarly, we sample M{M} initial vehicle locations from the same distribution.
 We set the demand $d_{i}$ for each customer to $1$ and the capacity of each vehicle to $3$.
@@
 Testing
-Testing is performed on $1000$ new instances for each setting of in-distribution $N$ and ${M}$ in Table 1 with the distributions from the training settings.
-For large-scale generalization in Table 2, we generate $100$ new instances.
+Testing is performed on $1000$ new instances for each setting of in-distribution NN and M{M} in Table˜1 with the distributions from the training settings.
+For large-scale generalization in Table˜2, we generate $100$ new instances.
 B.2.3 PARCO Network Hyperparameters
-Most hyperparameters are kept similar to Section B.1.3.
+Most hyperparameters are kept similar to §˜B.1.3.
 Encoder
@@
 Main Encoder.
-We employ $l=3$ attention layers in the encoder, with hidden dimension $d_{h}=128$, $8$ attention heads in the MHA, MLP hidden dimension set to $512$, with RMSNorm (Zhang & Sennrich, 2019) as normalization before the MHA and the MLP.
+We employ $l=3$ attention layers in the encoder, with hidden dimension $d_{h}=128$, $8$ attention heads in the MHA, MLP hidden dimension set to $512$, with RMSNorm [100] as normalization before the MHA and the MLP.
 Decoder
@@
 Communication Layer
-We employ a single transformer layer with hidden dimension $d_{h}=128$, $8$ attention heads in the MHA, MLP hidden dimension set to $512$, with RMSNorm (Zhang & Sennrich, 2019) as normalization before the MHA and the MLP.
-Note that unlike the encoder layer, which acts between all $M+N$ tokens, communication layers are lighter because they communicate between ${M}$ agents.
+We employ a single transformer layer with hidden dimension $d_{h}=128$, $8$ attention heads in the MHA, MLP hidden dimension set to $512$, with RMSNorm [100] as normalization before the MHA and the MLP.
+Note that unlike the encoder layer, which acts between all $M+N$ tokens, communication layers are lighter because they communicate between M{M} agents.
 Agent Handler
-We employ the Priority-based Conflict Handler guided by the model output probability for managing conflicts with priority given to the agent whose probability of selecting the conflicting action is the highest (see Section 4.6).
+We employ the Priority-based Conflict Handler guided by the model output probability for managing conflicts with priority given to the agent whose probability of selecting the conflicting action is the highest (see §˜4.6).
 B.2.4 PARCO Training Hyperparameters
 For each problem size, we train a single PARCO model that can effectively generalize over multiple size and agent distributions.
-We train PARCO with RL via SymNCO (Kim et al., 2022) with $K=8$ symmetric augmentations as shared REINFORCE baseline for $100$ epochs using the Adam optimizer (Kingma & Ba, 2014) with a total batch size $128$ on a single GPU and an initial learning rate of $10^{-4}$ with a step decay factor of $0.1$ after the $80$th and $95$th epochs.
+We train PARCO with RL via SymNCO [43] with $K=8$ symmetric augmentations as shared REINFORCE baseline for $100$ epochs using the Adam optimizer [46] with a total batch size $128$ on a single GPU and an initial learning rate of $10^{-4}$ with a step decay factor of $0.1$ after the $80$th and $95$th epochs.
 For each epoch, we sample $10^{5}$ randomly generated data.
@@
 Gurobi
-We implement the mathematical model described above in the exact solver Gurobi (Gurobi Optimization, LLC, 2024) with a time budget of 60 and 600 seconds per instance.
+We implement the mathematical model described above in the exact solver Gurobi [24] with a time budget of 60 and 600 seconds per instance.
 However, with both time budgets, Gurobi is only capable of generating solutions to the FFSP20 instances, similar to the findings made by Kwon et al.
-(2021) for the CPLEX solver.
+[50] for the CPLEX solver.
 Random and Shortest Job First (SJF)
@@
 [^2] To obtain results for the heuristics and metaheuristics, we used the implementation of Kwon et al.
-(2021), provided in the official GitHub repository of the paper: https://github.com/yd-kwon/MatNet
+[50], provided in the official GitHub repository of the paper: https://github.com/yd-kwon/MatNet
 Genetic Algorithm (GA)
-Genetic Algorithms are metaheuristics widely used by the OR community to tackle the FFSP (Kahraman et al., 2008).
+Genetic Algorithms are metaheuristics widely used by the OR community to tackle the FFSP [38].
 The GA iteratively improves multiple candidate solutions called chromosomes.
@@
 Particle Swarm Optimization (PSO)
-Finally, Particle Swarm Optimization iteratively updates multiple candidate solutions called particles, which are updated by the weighted sum of the inertial value, the local best, and the global best at each iteration (Singh & Mahapatra, 2012).
+Finally, Particle Swarm Optimization iteratively updates multiple candidate solutions called particles, which are updated by the weighted sum of the inertial value, the local best, and the global best at each iteration [81].
 In this implementation, particles use the same representation as GA chromosomes.
@@
 MatNet
-We benchmark PARCO mainly against MatNet (Kwon et al., 2021), a state-of-the-art NCO architecture for the FFSP.
-MatNet is an encoder-decoder architecture, which is inspired by the attention model (Kool et al., 2019a).
+We benchmark PARCO mainly against MatNet [50], a state-of-the-art NCO architecture for the FFSP.
+MatNet is an encoder-decoder architecture, which is inspired by the attention model [47].
 It extends the encoder of the attention model with a dual graph attention layer, a horizontal stack of two transformer blocks, capable of encoding nodes of different types in bipartite graph-like machines and jobs in the FFSP.
 Kwon et al.
-(2021) train MatNet using POMO (Kwon et al., 2020).
+[50] train MatNet using POMO [49].
 B.3.2 Datasets
@@
 We follow the instance generation scheme outlined in Kwon et al.
-(2021) sample processing times for job-machine pairs independently from a uniform distribution within the bounds $[2,10]$.
-For the first three FFSP instance types shown in Table 1 we also use the same instance sizes as Kwon et al.
-(2021) with $N=20,50\text{ and }100$ jobs and $M=12$ machines which are spread evenly over $S=3$ stages.
-To test for agent sensitivity in the FFSP, we fix the number of jobs to $N=50$ but alter the number of agents for the last three instance types shown in Table 1.
+[50] sample processing times for job-machine pairs independently from a uniform distribution within the bounds $[2,10]$.
+For the first three FFSP instance types shown in Table˜1 we also use the same instance sizes as Kwon et al.
+[50] with $N=20,50\text{ and }100$ jobs and $M=12$ machines which are spread evenly over $S=3$ stages.
+To test for agent sensitivity in the FFSP, we fix the number of jobs to $N=50$ but alter the number of agents for the last three instance types shown in Table˜1.
 Still, we use $S=3$ for this experiment, but alter the number of machines per stage to $M_{i}=6,8\text{ and }10$, yielding a total of 18, 24 and 30 agents, respectively.
@@
 To solve the FFSP with our PARCO method, we use a similar encoder as Kwon et al.
-(2021).
+[50].
 The MatNet encoder generates embeddings for all machines of all stages and the jobs they need to process, plus an additional dummy job embedding, which can be selected by any machine in each decoding step to skip to the next step.
@@
 We use $L=3$ encoder layers, generating embeddings of dimensionality $d_{h}=256$, which are split over $h=16$ attention heads in the MHA layers.
-Further, we employ Instance Normalization (Ulyanov et al., 2016) and a feed-forward network with 512 neurons in the transformer blocks of the encoder.
+Further, we employ Instance Normalization [87] and a feed-forward network with 512 neurons in the transformer blocks of the encoder.
 Decoder
 The machines are regarded as the agents in our PARCO framework.
-As such, their embeddings are used as queries $\bm{q}$ in the Multiple Pointer Mechanism Equation 7, while job embeddings are used as the keys and values.
+As such, their embeddings are used as queries $\bm{q}$ in the Multiple Pointer Mechanism Eq.˜6, while job embeddings are used as the keys and values.
 In each decoding step, the machine embeddings are fused with a projection of the time the respective machine becomes idle.
-Similarly, job embeddings are augmented with a linear transformation of the time they become available in the respective stage before entering the attention head in Equation 6.
+Similarly, job embeddings are augmented with a linear transformation of the time they become available in the respective stage before entering the attention head in Eq.˜5.
 Communication Layer
@@
 B.3.4 PARCO Training Hyperparameters
-Regarding the training setup, each training instance $i$ is augmented by a factor of 24, and the average makespan over the augmented instances is used as a shared baseline $b^{\text{shared}}_{i}$ for the REINFORCE gradient estimator of Equation 9.
-We use the Adam optimizer (Kingma & Ba, 2014) with a learning rate of $4\times 10^{-4}$, which we alter during training using a cosine annealing scheme.
-We train separate models for the environment configurations used in Table 1.
+Regarding the training setup, each training instance $i$ is augmented by a factor of 24, and the average makespan over the augmented instances is used as a shared baseline $b^{\text{shared}}_{i}$ for the REINFORCE gradient estimator of Eq.˜8.
+We use the Adam optimizer [46] with a learning rate of $4\times 10^{-4}$, which we alter during training using a cosine annealing scheme.
+We train separate models for the environment configurations used in Table˜1.
 We train models corresponding to environments with 20 jobs for 100, with 50 jobs for 150 and with 100 jobs for 200 epochs.
@@
 The following figures visualize the decoding for the machines of a given stage using MatNet and PARCO.
-As one can see in Figure 5(a), MatNet requires a decoder forward pass for each machine to schedule a job on each of them.
-In contrast, as detailed in Figure 5(b), PARCO can schedule jobs on all machines simultaneously through its Multiple Pointer Mechanism and Agent Handler, leading to significant efficiency gains.
+As one can see in Fig.˜6(a), MatNet requires a decoder forward pass for each machine to schedule a job on each of them.
+In contrast, as detailed in Fig.˜6(b), PARCO can schedule jobs on all machines simultaneously through its Multiple Pointer Mechanism and Agent Handler, leading to significant efficiency gains.
 (a) An FFSP Decoding Step with MatNet
@@
 B.4.2 Software
-We used Python 3.12, PyTorch 2.5 (Paszke et al., 2019) coupled with PyTorch Lightning (Falcon & The PyTorch Lightning team, 2019) with most code based on the RL4CO library (Berto et al., 2024a).
-The operating system is Ubuntu 24.04 LTS.
+We used Python 3.12, PyTorch 2.5 [78] coupled with PyTorch Lightning [17] with most code based on the RL4CO library [5]. The operating system is Ubuntu 24.04 LTS.
 B.5 Source Code
-B.5.1 Licenses
-We provide the summary of licenses Table 3.
+We value open reproducibility and release our source code at https://github.com/ai4co/parco. We also provide the summary of licenses in Table˜3.
 Table 3: Summary of licenses for used assets.
 Resource Type License
-OR-Tools (Furnon & Perron, 2024) Code Apache License, Version 2.0
-AM (Kool et al., 2019a) Code MIT License
-ET (Son et al., 2024) Code Available on Github
-DPN (Zheng et al., 2024a) Code MIT License
-DRLLi (Li et al., 2022) Code Available on Github
-2D-Ptr (Liu et al., 2024b) Code/Dataset Available on Github
-HAM (Li et al., 2021a) Code MIT License
-Gurobi (Gurobi Optimization, LLC, 2024) Code Commercial license (free for academic use)
-MatNet (Kwon et al., 2021) Code/Dataset MIT License
-RL4CO (Berto et al., 2024a) Library MIT License
-B.5.2 Code Release
-We value open reproducibility and release our source code at https://github.com/ai4co/parco.
+OR-Tools [19] Code Apache License, Version 2.0
+AM [47] Code MIT License
+ET [84] Code Available on Github
+DPN [104] Code MIT License
+DRLLi [55] Code Available on Github
+2D-Ptr [63] Code/Dataset Available on Github
+HAM [54] Code MIT License
+Gurobi [24] Code Commercial license (free for academic use)
+MatNet [50] Code/Dataset MIT License
+RL4CO [5] Library MIT License
+Appendix C Additional Materials
+C.1 Further Discussions
+Follow-ups
+We acknowledge there are works based on or inspired by PARCO already out in the wild at the time of publication.
+These include extending to profiled VRP [29, 30], mixed-shelves picker-routing [67], and self-improvement for better performance [68].
+We hope to see even more in the future from diverse research groups, and we remain available for discussions, including through the AI4CO Slack.
+Possible future directions
+In this paragraph, we will share a short list of possible additional future directions for PARCO.
+One promising direction for NCO is to employ Large Language Models (LLMs) for automating the design of algorithms and particularly heuristics [62, 97, 27].
+Several components of PARCO could be designed by LLMs, including the conflict handling mechanism or attention biases dependent on specific problems [31, 90] as done by Tran et al.
+[86].
+Of interest would also be extensions to other problem variants [99].
+Integrating parallel decoding into search methods, as NDS [28], may also yield much more efficient models and enable them to capture different agents.
+C.2 Comparison with Decentralized and Graph-Based Communication Methods
+We additionally compare PARCO with decentralized methods and alternative communication models like Graph Neural Networks (GNNs).
+We experiment on the min-max traveling salesman problem (mTSP), benchmarking PARCO against notable decentralized methods like GNN-DisPN [16], DAN [10] and GNN-based communication models like ScheduleNet [75].
+The results, shown in Table˜4 with instances from Park et al.
+[75] as well as in the mTSPLib [^4] https://profs.info.uaic.ro/mihaela.breaban/mtsplib/MinMaxMTSP/ in Table˜5 – additionally adding the HGA solver for further comparison from Mahmoudinazlou and Kwon [72] – demonstrate that PARCO consistently outperforms these methods in solution quality.
+Centralized training with parallel decoding, as used in PARCO, offers distinct advantages by enabling global coordination and highly efficient solution construction.
+For example, GNN-based communication in models like ScheduleNet incurs significantly higher computational overhead, making our approach much faster than those methods.
+Table 4: Cost comparison on the mTSP with different numbers of salesmen $M$ and number of nodes NN.
+NN $50$ $100$ $200$ Gap(%)
+$M$ 5 7 10 5 10 15 10 15 20
+LKH3 (solver) 2.00 1.95 1.91 2.20 1.97 1.98 2.04 2.00 1.97 -
+OR-Tools (solver) 2.04 1.96 1.96 2.36 2.29 2.25 2.57 2.59 2.59 14.42
+GNN-DisPN (g.) 2.14 2.10 1.99 2.56 2.22 2.04 2.97 2.30 2.15 13.45
+DAN (g.) 2.29 2.11 2.03 2.72 2.17 2.09 2.40 2.20 2.15 11.75
+SchedNet (g.) 2.17 2.07 1.98 2.59 2.13 2.07 2.45 2.24 2.17 10.16
+PARCO (g.) 2.12 2.00 1.92 2.47 2.02 1.98 2.28 2.06 1.99 4.43
+DAN (s.) 2.12 1.99 1.95 2.55 2.05 2.00 2.29 2.13 2.07 6.13
+SchedNet (s.) 2.07 1.99 1.92 2.43 2.03 1.99 2.25 2.08 2.05 4.29
+PARCO (s.) 2.07 1.98 1.91 2.38 1.99 1.98 2.22 2.03 1.98 2.80
+Table 5: Results for the mTSPLib. CPLEX results with * are optimal solutions. Otherwise, the best-known upper bound of CPLEX results are reported.
+instance_N\_N eil51 berlin52 eil76 rat99 Gap $(\%)$
+$M$ 2 3 5 7 2 3 5 7 2 3 5 7 2 3 5 7
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+We note that these experiments refer to a previous version of PARCO in which the conflict handler was not properly working and was similar to the random handler – we would expect retraining PARCO on the newest implementation would perform even better.
+As the scope of this comparison is to compare against decentralized methods, we do not include more recent approaches as ET [84] or DPN [104] which would outperform this old PARCO version – albeit with slower decoding.
+We expect that recent follow-ups of our work, which allow for stopping actions by adding special tokens to the embedding, as MACSIM [68], would outperform ET and DPN, perhaps considerably.
+We leave this as an interesting direction for future work.
+C.3 Convergence Rates
+Table 6: Convergence rates in different problems: cost as a percentage of training budget.
+10% 25% 50% 75% 100%
+HCVRP 5.13 $\pm$ 0.09 5.00 $\pm$ 0.06 4.90 $\pm$ 0.05 4.88 $\pm$ 0.03 4.79 $\pm$ 0.03
+OMDCPDP 46.23 $\pm$ 0.21 45.57 $\pm$ 0.15 45.34 $\pm$ 0.11 44.87 $\pm$ 0.10 44.48 $\pm$ 0.09
+FFSP 95.45 $\pm$ 0.52 94.32 $\pm$ 0.31 92.88 $\pm$ 0.20 92.15 $\pm$ 0.12 91.48 $\pm$ 0.08
+Table˜6 shows the convergence rate of PARCO across different problems on validation datasets at different training budgets (in percentage). PARCO is robust during training and converges stably.
+C.4 XXL Instances
+To further strengthen our results, we have tested PARCO for even larger scales in both the number of nodes and the number of agents.
+We generated 16 new instances for $N=5000$ nodes and 3 different values of $M$ agents (a total of 48 instances) of OMDCPDP.
+We evaluated OR-Tools with 1 hour of runtime and greedy performance for HAM and PARCO.
+Table 7: Large-scale generalization results for OMDCPDP with $N=5000$.
+$M=500$ $M=750$ $M=1,000$
+Obj. Gap Time Obj. Gap Time Obj. Gap Time
+OR-Tools 5575.73 134.06% 3600s 5127.46 115.24% 3600s 4974.81 188.10% 3600s
+HAM 4813.99 102.08% 17.4s 3732.06 97.33% 19.5s 3258.26 88.69% 22.3s
+PARCO 2382.22 0.0% 0.21s 1891.28 0.0% 0.21s 1726.78 0.0% 0.22s
+As shown in Table˜7, PARCO excels at generalization at XXL scales with $50\times$ the number of nodes and agents seen during training and up to $1,000$ agents.
+Thanks to its massively parallel structure, PARCO can solve such instances in a fraction of a second with better results than OR-Tools – at a $10,000\times$ speedup.
+This makes PARCO ideal for real-world, real-time, large-scale complex problems.
+NeurIPS Paper Checklist
+  1. 1.
+    Claims
+  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
+  3. Answer: [Yes]
+  4. Justification: The main claims on the abstract and introduction accurately reflect the paper’s contribution and scope (see methodology of §˜4 and experimental results §˜5).
+  5. 2.
+    Limitations
+  6. Question: Does the paper discuss the limitations of the work performed by the authors?
+  7. Answer: [Yes]
+  8. Justification: We include a paragraph in §˜6 discussing limitations and future works.
+  9. 3.
+    Theory assumptions and proofs
+  10. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
+  11. Answer: [N/A]
+  12. Justification: This paper does not include theoretical results.
+  13. 4.
+    Experimental result reproducibility
+  14.
+Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
+  15. Answer: [Yes]
+  16.
+Justification: We give a detailed description of our method in §˜4 and discuss the experimental settings in §˜5 with further information in the Appendix.
+The source code of PARCO is also publicly available.
+  17. 5.
+    Open access to data and code
+  18. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
+  19. Answer: [Yes]
+  20. Justification: The code of PARCO is available at https://github.com/ai4co/parco. We provide detailed instructions on how to reproduce the main experimental results.
+  21. 6.
+    Experimental setting/details
+  22. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
+  23. Answer: [Yes]
+  24.
+Justification: We describe the settings of the training and test details in §˜5, supplemented by further details for each experiment in Appendix˜B.
+The original configurations are additionally available in the publicly available code to foster reproducibility and openness.
+  25. 7.
+    Experiment statistical significance
+  26. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
+  27. Answer: [Yes]
+  28.
+Justification: We report the standard deviation in Fig.˜3 with experiments run on 3 different seeds.
+For the main results including Table˜1, we adhere to the established standard in machine learning for combinatorial optimization papers, which involves reporting averaged results for several dataset instances each to ensure statistical significance: 1,280, 1,000, and 100 for HCVRP, OMDCPDP, and FFSP, respectively.
+  29. 8.
+    Experiments compute resources
+  30. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
+  31. Answer: [Yes]
+  32. Justification: We report our resources in §˜5.
+  33. 9.
+    Code of ethics
+  34. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
+  35. Answer: [Yes]
+  36. Justification: Our research conforms to the NeurIPS Code of Ethics
+  37. 10.
+    Broader impacts
+  38. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
+  39. Answer: [N/A]
+  40.
+Justification: The adoption of PARCO is not expected to have any negative societal impact.
+On the contrary, we envision machine learning approaches as PARCO may become more prevalent in solving CO problems and thus lead to positive societal benefits, including reduced resource consumption, better management of logistics such as disaster relief efforts, and democratization of optimization solvers.
+  41. 11.
+    Safeguards
+  42.
+Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
+  43. Answer: [N/A]
+  44. Justification: There is no high risk of misuse for our released assets.
+  45. 12.
+    Licenses for existing assets
+  46.
+Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
+  47. Answer: [Yes]
+  48. Justification: Creators of assets (e.g., code, data, models) used in the paper are properly credited in Table˜3.
+  49. 13.
+    New assets
+  50. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
+  51. Answer: [Yes]
+  52. Justification: All introduced assets are well documented. We provide clear instructions, including checkpoints release through HuggingFace.
+  53. 14.
+    Crowdsourcing and research with human subjects
+  54.
+Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
+  55. Answer: [N/A]
+  56. Justification: The paper does not involve crowdsourcing nor research with human subjects.
+  57. 15.
+    Institutional review board (IRB) approvals or equivalent for research with human subjects
+  58.
+Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
+  59. Answer: [N/A]
+  60. Justification: The paper does not involve crowdsourcing nor research with human subjects.
+  61. 16.
+    Declaration of LLM usage
+  62.
+Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research?
+Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
+  63. Answer: [N/A]
+  64. Justification: LLMs are used only for writing, editing, or formatting purposes and do not impact the core methodology, scientific rigor, or originality of the research.